### v1

# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

Abstract
We study the problem of differentially-private (DP) stochastic (convex-concave) saddle-points in the polyhedral setting. We propose $(\varepsilon,\delta)$-DP algorithms based on stochastic mirror descent that attain nearly dimension-independent convergence rates for the expected duality gap, a type of guarantee that was known before only for bilinear objectives. For convex-concave and first-order-smooth stochastic objectives, our algorithms attain a rate of $\sqrt{\log(d)/n}+(\log(d)^{3/2}/[n\varepsilon])^{1/3}$, where $d$ is the dimension of the problem and $n$ the dataset size. Under an additional second-order-smoothness assumption, we improve the rate on the expected gap to $\sqrt{\log(d)/n}+(\log(d)^{3/2}/[n\varepsilon])^{2/5}$. Under this additional assumption, we also show, by using bias-reduced gradient estimators, that the duality gap is bounded by $\log(d)/\sqrt{n}+\log(d)/[n\varepsilon]^{1/2}$ with constant success probability. This result provides evidence of the near-optimality of the approach. Finally, we show that combining our methods with acceleration techniques from online learning leads to the first algorithm for DP Stochastic Convex Optimization in the polyhedral setting that is not based on Frank-Wolfe methods. For convex and first-order-smooth stochastic objectives, our algorithms attain an excess risk of $\sqrt{\log(d)/n}+\log(d)^{7/10}/[n\varepsilon]^{2/5}$, and when additionally assuming second-order-smoothness, we improve the rate to $\sqrt{\log(d)/n}+\log(d)/\sqrt{n\varepsilon}$. Instrumental to all of these results are various extensions of the classical Maurey Sparsification Lemma, which may be of independent interest.
1 Introduction
Protecting privacy of users’ data has become a key desideratum in machine learning. Such a requirement is fundamental in situations involving e.g. medical or financial data. In this context, differential privacy (DP) has become the de facto standard for private data analysis, and it has been intensively studied for nearly 2 decades (Dwork et al., 2006b, 2014).
Recently, an emerging line of work has focused on solving stochastic saddle-point [^1] We will restrict ourselves to the convex-concave setting, and therefore omit this from the problem denomination (SSP) problems under DP (Boob and Guzmán, 2023; Zhang et al., 2022a; Yang et al., 2022; Kang et al., 2022; Bassily et al., 2023). While this problem has been primarily motivated by group fairnes in machine learning (Jagielski et al., 2019; Li et al., 2019; Lowy et al., 2023), its generality may be of use in problems as varied as learning under distributional robustness (Mohajerin Esfahani and Kuhn, 2018) and synthetic data generation (Gaboardi et al., 2014). In relation to the latter problem, a long list of works has explored the idea of using zero-sum games (corresponding to bilinear saddle-points with polyhedral feasible sets) for private query answering and private synthetic data generation (e.g.,  Hardt and Rothblum (2010); Hardt et al. (2012); Hsu et al. (2013); Gaboardi et al. (2014)).
Under privacy constraints, there is a major discrepancy in the literature between the bilinear and convex-concave settings. While bilinear problems have convergences rates –in terms of the expected duality gap– which scale only poly-logarithmically with the ambient dimensions (e.g., Hardt and Rothblum (2010)), all known rates for general convex-concave objectives scale polynomially with the dimension (e.g., Bassily et al. (2023)). This can be attributed to the fact that all existing results for convex-concave settings work under Euclidean geometry, which assumes $\ell_{2}$ bounded feasible sets and objectives that are Lipschitz and smooth with respect to $\|\cdot\|_{2}$. Therefore, known polynomial-in-the-dimension lower bounds for DP Stochastic Convex Optimization (SCO) in the Euclidean setting (Bassily et al., 2014) apply. Recently, (Asi et al., 2021b; Bassily et al., 2021a) showed that it is possible to avoid the lower bounds for DP-SCO under additional geometric structure on the problem, which we refer to as the polyhedral setting. It consists of privately minimizing an stochastic objective which is Lipschitz and smooth with respect to $\|\cdot\|_{1}$ over a polyhedral feasible set. Inspired by these recent developments, in this work we extend the scope of polyhedral settings to the class of convex-concave, Lipschitz and smooth [^2] We only consider smooth objectives because even for DP-SCO in the polyhedral setting, smoothness is a necessary condition to get nearly dimension-independent rates (Asi et al., 2021b). SSP problems (see Section 4 for a formal problem statement). Below we summarize our results and main contributions.
Summary of Contributions.
We prove the first nearly-dimension independent convegence rates for convex-concave DP-SSP in the polyhedral setting. Furthemore, for (nearly) quadratic objectives, our rates are optimal. In addition, our mirror-descent based algorithms achieve nearly-dimension independent convegence rates for DP-SCO in the polyhedral setting. Moreover, our work is the first that does not use a Frank-Wolfe based method to get this type of guarantee. Table 1 summarizes the obtained rates.
  • We propose two algorithms for DP-SSP, based on (entropic) Stochastic Mirror Descent (SMD). As opposed to previous results for bilinear settings where privatization by sampling from vertices leads to unbiased gradient estimators (Hsu et al., 2013), the main challenge we encounter is controlling the bias in the more general (stochastic) convex-concave setting. The first algorithm (Section 4.1) controls bias by simply using a large number of samples in every iteration. The second algorithm (Section 4.2) builds upon the first one by incorporating a technique from Blanchet and Glynn (2015) to reduce the bias of gradients evaluated at the private estimators of the iterates.
  • In our algorithms, the iterates are privately estimated by sampling vertices proportionally to their weights, a technique known as Maurey’s Sparsification Lemma (Pisier, 1980-1981) [^3] This method is also known as the Maurey Empirical Method.. For nonlinear objectives, evaluating the duality gap and gradients of the objective at the iterate estimators introduces bias. To control the bias and other estimation errors, we provide various extensions of this classical result that work in terms of functional approximations. We state these results in more generality than what is needed for our analyses, as they could be of independent interest. See Section 3 for more details.
  • We give the first mirror descent style algorithm for DP-SCO in the polyhedral setting with nearly dimension-independent excess risk guarantees by combining the ideas we used for DP-SSP with the Anytime Online-to-Batch Converstion from Cutkosky (2019). See Section 5. While these rates are not optimal, we expect they could inspire future work into more practical algorithms for this setting.
Setting
Objective
Previous best-known rate
Our rate
SSP
Bilinear
$\color[rgb]{1,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$ (Hardt and Rothblum, 2010)
$\color[rgb]{1,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$
(Thm. 4.1)
Quadratic
      None
$\color[rgb]{1,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$
(Thm. 4.1)
$L_{2}$-smooth
      None
$\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{2/5}$
(Thm. 4.1)
$L_{1}$-smooth
      None
$\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/3}$
(Thm. 4.2)
$L_{2}$-smooth
      None
$\color[rgb]{1,0,0}\frac{\ell}{\sqrt{n}}+\frac{\ell}{\sqrt{n\varepsilon}}{}^{(\ast)}$
(Thm. 4.3)
SCO
$L_{2}$-smooth
      None
$\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}}{\sqrt{n\varepsilon}}$
(Thm. 5.1)
$L_{1}$-smooth
$\color[rgb]{1,0,0}\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{2/3}}{(n\varepsilon)^{2/3}}$ (Asi et al., 2021b)
$\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{7/10}}{(n\varepsilon)^{2/5}}$
(Thm. 5.1)
Table 1: We omit polylog($n,1/\delta$) factors, function-class parameters and constants. We refer to Lipschitz and smooth functions as $L_{1}$-smooth. If the function has, in addition, smooth partial derivatives, then we refer to it as $L_{2}$-smooth. We denote $\ell=\log(d_{x})+\log(d_{y})$ and $\ell_{x}=\log(d_{x})$. Rates in red are optimal up to polylog factors. ${}^{(\ast)}$ This rate holds with constant probability while the rest hold in expectation.
1.1 Related Work
DP stochastic saddle-points.
Stochastic (non-private) saddle-point problems have connections to optimization, economics, machine learning, and many other fields. One such algorithm for solving these problems is the stochastic entropic mirror-descent algorithm (Nemirovski et al., 2009). Recently, a line of research has studied DP stochastic saddle-points (Boob and Guzmán, 2023; Zhang et al., 2022a; Yang et al., 2022; Kang et al., 2022; Bassily et al., 2023). These works focus exclusively on Euclidean setups, which necessarily leads to a polynomial dependence on the dimension; furthermore, techniques of these works are entirely unrelated to ours. In addition, with only limited exceptions (Boob and Guzmán, 2023; Bassily et al., 2023), most works provide guarantees on the weak duality gap. Bassily et al. (2023) justify that the (strong) duality gap is a more meaningful and challenging efficiency estimate. This is the convergence criterion that we use in this work.
Optimization in the polyhedral setting.
Closest to this work is DP stochastic convex optimization (SCO) in $\ell_{1}$ setup (Asi et al., 2021b; Bassily et al., 2021b; Han et al., 2022), as well as earlier work on private empirical risk minimization (ERM) (Jain and Thakurta, 2014; Talwar et al., 2014, 2015). These works attain nearly dimension-independent excess risk rates with private variants of the (stochastic) Frank-Wolfe method. Even though this method leads to optimal rates for DP-SCO, it is not known to yield any nontrivial rates for saddle-point problems. Even without DP constraints, it is not known how to solve SSP problems with Frank-Wolfe methods unless strong additional assumptions are made: namely, either simultaneously assuming smoothness and strong convexity/strong concavity [^4] These assumptions are indeed strong: If $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ is an $L_{1}$-smooth and $\mu$-strongly convex function w.r.t. $\|\cdot\|_{1}$, then $L/\mu=\Omega(d)$. This follows by conjugate duality and the argument in (d’Aspremont et al., 2018, Example 4.1), or strong convexity of the feasible sets (Gidel et al., 2017). For these reasons, those assumptions are of no use for the interests of this work.
Maurey Sparsification.
This is a classical technique in high-dimensional convex geometry (Pisier, 1980-1981). The idea is based on sampling vertices with the barycentric coordinates weights and taking the empirical average, as nicely illustrated in the introduction of the book (Vershynin, 2018). This approximation is established for norms with nontrivial concentration estimates for independent random sums, e.g. quantified by their Rademacher type (Pisier, 2016). This technique has been used systematically as a compression-type argument to obtain faster rates in statistical estimation and learning (e.g. (Barron, 1993; Zhang, 2002)). More recently, Acharya et al. (2019) prove a result similar to Lemma 3.1, a functional approximation for randomized sparsification, in the context of model compression for distributed learning of generalized linear models. They also proposed a sparsified mirror descent algorithm that works in the $\ell_{1}/\ell_{q}$ setting where $q<\infty$, resorting on smoothnes properties of the Bregman divergence. Furthemore, they only sparsify the output models of each machine. We do not follow this approach. Instead, we sparsify every iterate of mirror descent. Furthermore, we show that entropic SMD (using a nonsmooth Bregman divergence) with iterate sparsification works for convex and smooth losses, which to our knowledge is an entirely new result, and does not follow from direct ex-post model compression.
Debiasing.
This is a technique proposed in the simulation literature (Blanchet and Glynn, 2015), to leverage biased estimators whose accuracy improves with larger sample size. The main difference between the works on simulation and our approach is the finite sample size, which limits the amount of bias-reduction. Asi et al. (2021a) studied bias-reduction in regularization for SCO, but their results are unrelated to ours. In concurrent work (Ghazi et al., 2024), a gradient debiasing procedure is developed for Euclidean DP-SCO. Although our results do not depend on their techniques, we borrow some inspiration from them. Note futher that most of our results do not use bias-reduction.
Minimizing the worst-group loss.
In independent and concurrent work, Zhou and Bassily (2024) investigate empirical and population worst-group risk minimization under DP, with a focus on group fairness and robust learning. While the convex-linear loss functions of their work correspond to a particular case of ours (convex-concave), their use of Euclidean geometry makes their results incomparable to ours. Furthermore, their sequential regularization approach is largely independent from ours.
2 Preliminaries
Given $x\in\mathbb{R}^{d}$, we denote the $1$-norm and $\infty$-norm by $\|x\|_{1}=\sum_{j\in[d]}|x_{j}|$ and $\|x\|_{\infty}=\max_{j\in[d]}|x_{j}|$, respectively. A differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{0}$-Lipschitz w.r.t. $\|\cdot\|_{1}$ over a set $\mathcal{X}$ if $\|\nabla F(x)\|_{\infty}\leq L_{0}$ for all $x\in\mathcal{X}$. It is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$ if $\|\nabla F(x^{1})-\nabla F(x^{2})\|_{\infty}\leq L_{1}\|x^{1}-x^{2}\|_{1}$ for all $x^{1},x^{2}\in\mathcal{X}$. We say a function is first-order-smooth if it is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. We say a function is second-order-smooth if it is first-order-smooth and its partial derivates are $L_{2}$-smooth w.r.t $\|\cdot\|_{1}$. [^5] Note that this assumption is milder than the typical second-order smoothness assumption in optimization, and it is satisfied by the widely used quadratic and logistic losses. For $a,b:\mathbb{R}^{p}\mapsto\mathbb{R}$, we say $a\lesssim b$ if there exists an absolute constant $C>0$ such that for all $x\in\mathbb{R}^{p}$, $a(x)\leq Cb(x)$. When $a\lesssim b$ and $b\lesssim a$, we write $a\eqsim b$. We denote by $e_{i}$ the $i$-th canonical vector (the dimension being clear from context). The standard simplex in $\mathbb{R}^{d}$ is the set $\Delta_{d}=\{x\in\mathbb{R}^{d}:\|x\|_{1}=1,x_{j}\geq 0\text{ for all }j\in[d]\}$. A vector $x\in\Delta$ induces a probability distribution over the canonical vectors, which we denote by $P_{x}$. Hence, $\hat{x}\sim P_{x}$ means that $\hat{x}$ takes the value $e_{i}$ with probability $x_{i}$. Note that if $\hat{x}\sim P_{x}$, then $\mathbb{E}[\hat{x}]=x$. Finally, we define the notion of differential privacy.
Definition 2.1 (Differential Privacy Dwork et al. (2006a)).
A randomized algorithm $\mathcal{A}:\mathcal{Z}^{n}\mapsto\mathcal{X}$ is $(\varepsilon,\delta)$-differentially private (abbreviated $(\varepsilon,\delta)$-DP) if for any pair of datasets $S$ and $S^{\prime}$ differing in at most one data point (denoted by $S\sim S^{\prime}$) and any event $\mathcal{E}\subseteq\mathcal{X}$, $\mathbb{P}[\mathcal{A}(S)\in\mathcal{E}]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(S^{\prime})\in\mathcal{E}]+\delta$. When $\delta=0$, we denote the $(\varepsilon,0)$-DP property as $\varepsilon$-DP.
3 Maurey Sparsification with Functional Value Approximation
The classical Maurey Sparsification Lemma (Pisier, 1980-1981) is an application of the probabilistic method that leads to an approximate and (nearly) dimension-independent version of Carathéodory’s Theorem (Carathéodory, 1911). While this method is well-known to apply with smooth norms, it appears that for norms such as $\|\cdot\|_{1}$ such approximations are out of reach. [^6] In fact, $\|\cdot\|_{1}$ is as far as being smooth as possible, in the sense of having trivial Rademacher type (Pisier, 2016). Furthermore, polynomial in $d$ mean estimation lower bounds with respect to $\|\cdot\|_{1}$ are known (Waggoner, 2015). We circumvent this limitation by directly working with function value approximations. This introduces its own additional challenges which we address in Section 4. Proofs in this section can be found in Appendix D.
We start with a simple adaptation of Maurey’s method for function approximation in the general smooth case. A similar proof appears in (Acharya et al., 2019).
Lemma 3.1 (Function value approximation in expectation).
[…]
$\mathbb{E}\left[F\left(\sum_{t=1}^{T}\lambda_{t}a^{t}\right)-F\left(\sum_{t=1}^{T}\lambda_{t}x^{t}\right)\right]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right].$
Proof.
Let $\bar{x}^{T}=\sum_{t=1}^{T}\lambda_{t}x^{t}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\lambda_{t}a^{t}$. Noting that $F(\bar{a}^{T})=F(\bar{x}^{T}+\bar{a}^{T}-\bar{x}^{T})$, and using repeatedly the $L_{1}$-smoothness of $F$, we have
$F(\bar{a}^{T})$ $\leq F(\bar{x}^{T}+\bar{a}^{T-1}-\bar{x}^{T-1})+\langle\nabla F(\bar{x}^{T}+\bar{a}^{T-1}-\bar{x}^{T-1}),\lambda_{T}(a^{T}-x^{T})\rangle+\frac{L_{1}\lambda_{T}^{2}\left\|a^{T}-x^{T}\right\|^{2}}{2}$
$\ldots$ $\leq F\left(\bar{x}^{T}\right)+\sum_{t\in[T]}\underbrace{\Big{\langle}\nabla F\Big{(}\bar{x}^{T}+\bar{a}^{t-1}-\bar{x}^{t-1}\Big{)},\lambda_{t}(a^{t}-x^{t})\Big{\rangle}}_{\alpha_{t}}+\frac{L_{1}}{2}\sum_{t\in[T]}\lambda_{t}^{2}\|a^{t}-x^{t}\|^{2}$
where $\alpha_{t}=\Big{\langle}\nabla F\Big{(}\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\Big{)},\lambda_{t}(a^{t}-x^{t})\Big{\rangle}$ is a martingale difference sequence. Hence, $\mathbb{E}\big{[}\sum_{t\in[T]}\alpha_{t}\big{]}=0$ and $\mathbb{E}[F(\bar{a}^{T})-F(\bar{x}^{T})]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$. ∎
We now instantiate Lemma 3.1 for the space $(\mathbb{R}^{d},\|\cdot\|_{1})$ with $\lambda_{i}=\frac{1}{T}$, $i\in[T]$. Here, $x^{t}\in\Delta_{d}$ represents the $t$-th iterate of our algorithm and $a^{t}\sim P_{x^{t}}$ a private, unbiased estimator of $x^{t}$.
Corollary 3.1 (Function value approximation w.r.t. $\|\cdot\|_{1}$ over the simplex).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors in $\Delta_{d}$ and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then
$\left|\mathbb{E}\left[F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right|\leq\frac{2L_{1}}{T}.$
The above result bounds the bias incurred by evaluating gradients at sampled vertices. More concretely, $\left\|\mathbb{E}\left[\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right\|_{\infty}\lesssim\frac{L_{2}}{T}$ for second-order-smooth functions since the same bound is true for the bias of each partial derivative (see Corollary D.1). We can also bound the bias for first-order-smooth functions by $\frac{L_{1}}{\sqrt{T}}$, which we do by first approximating the partial derivatives using finite differences, and then using the above corollary directly for the smooth objective function (see Corollary D.2). These results are new, to the best of our knowledge.
We will also need to bound the second-moment of the error incurred by evaluating gradients at sampled vertices. Towards this, first we note that the in-expectation guarantees obtained in Lemma 3.1 can be turned into high-probability guarantees by controlling the increments of the martingale difference sequence $(\alpha_{t})_{t\in[T]}$ with the bounded differences inequality.
Lemma 3.2 (Function value approximation with high probability).
[…]
$\mathbb{P}\left[F\left(\sum_{t=1}^{T}\lambda_{t}a^{t}\right)-F\left(\sum_{t=1}^{T}\lambda_{t}x^{t}\right)\geq\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\beta\frac{L_{0}D}{\sqrt{2}}\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\,\right]\leq\exp(-\beta^{2}).$
We use this high-probability guarantee to give bounds on the expected second-moment error, with respect to $\|\cdot\|_{\infty}$. We have the following result, which we instantiate for two particular cases.
Lemma 3.3 (Second moment of maximum expected error).
Let $F_{1},...,F_{M}:\mathbb{R}^{d}\to\mathbb{R}$ be a collection of $L_{0}$-Lipschitz and $L_{1}$-smooth functions. Let $x^{1},\ldots,x^{T}\in\mathbb{R}^{d}$, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$ and $a^{1},...,a^{T}$ be a collection of random vectors such that $\mathbb{E}[a^{t}|a^{t-1},\ldots,a^{1}]=x^{t}$ and $\|a^{t}-x^{t}\|_{1}\leq D$ a.s. for all $t\in[T]$. Let $\bar{a}^{T}=\sum_{t\in[T]}\lambda_{t}a^{t}$ and $\bar{x}^{T}=\sum_{t\in[T]}\lambda_{t}x^{t}$. Then
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{x})-F_{j}(\bar{a}^{T})|^{2}\right]\leq\frac{5L_{1}^{2}D^{4}}{4}\left(\sum_{t=1}^{T}\lambda_{t}^{2}\right)^{2}+\frac{L_{0}^{2}D^{2}(6+\log(M))}{2}\sum_{t=1}^{T}\lambda_{t}^{2}.$
Note that when $F$ is second-order-smooth, applying Lemma 3.3 with $F_{j}=\nabla_{j}F$, $j\in[d]$, leads to $\mathbb{E}\|\nabla F\left(\frac{1}{T}\sum_{t\in[T]}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t\in[T]}x^{t}\right)\|^{2}_{\infty}\lesssim\frac{L_{2}^{2}}{T^{2}}+\frac{L_{1}^{2}(1+\log(d))}{T}$. We can also get a bound of $\frac{L_{1}^{2}}{T^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d)}}{\sqrt{T}}$ for first-order-smooth functions by an approximation with finite differences and Lemma 3.3 with $F_{j}(\cdot)=F(\cdot+re_{j})$ for $j\in[d]$. See Corollaries D.3 and  D.4 for details.
4 Differentially Private Stochastic Saddle Points
Problem set-up.
[…]
$\min_{x\in\Delta_{x}}\max_{y\in\Delta_{y}}F_{\mathcal{D}}(x,y),$
under the restriction of using $(\varepsilon,\delta)$-DP algorithms. The efficiency of a candidate solution $(x,y)\in\mathcal{X}\times\mathcal{Y}$, is measured by the duality gap, $\operatorname{Gap}(x,y)=\max_{v\in\mathcal{X},w\in\mathcal{Y}}(F_{\mathcal{D}}(x,w)-F_{\mathcal{D}}(v,y))$: this is a natural measure as it bounds the gap from Nash equilibria when $x,y$ are seen as players in a zero-sum game. Instrumental to the accuracy analysis of algorithms is the so-called saddle operator (a.k.a. the monotone operator associated to the saddle-point problem): given $f:{\cal X}\times{\cal Y}\mapsto\mathbb{R}$, we let $\Phi(x,y)=\big{(}\nabla_{x}f(x,y),-\nabla_{y}f(x,y)\big{)}$ (Rockafellar, 1970).
4.1 Private Stochastic Mirror Descent with Sampling from Vertices
Our first algorithm is a (entropic) stochastic mirror descent (SMD) method with sampling from vertices; see Algorithm 1. A known observation is that sampling from vertices with exponential weights (as induced by SMD) guarantees privacy as it is an instance of the exponential mechanism (e.g., Hsu et al. (2013)). In the bilinear case, this is enough to obtain guarantees for the duality gap as this sampling provides an unbiased first-order stochastic oracle. However, for nonlinear objectives (e.g., general convex-concave losses), it is unclear the accuracy of sampling from vertices.
A key challenge is that the provided extensions of Maurey’s Lemma lead to biased updates. We can reduce this bias to any desired accuracy by increasing the sample size; however, an inherent tradeoff arises: on the one hand, accuracy improves; on the other hand, the privacy degrades. In the following results we optimally tradeoff both considerations under different assumptions on the objective function. Theorem 4.1 considers second-order-smoothness, while Theorem 4.2 only assumes first-order smoothness. We defer the explicit choice of parameters and proof to Appendix E.
Algorithm 1 Private stochastic mirror descent with sequential Maurey sparsification
0:  Dataset $S=\{z^{1},...,z^{n}\}$, Number of steps $T$, Step-size $\tau$, Number of iterate samples $K$
1:  $x^{1}=(1/d_{x},...,1/d_{x})$, $y^{1}=(1/d_{y},...,1/d_{y})$
2:  for $t=1$ to $T$ do
3:     Let $B^{t}$ be a batch of $B=n/T$ fresh samples
[…]
5:     $g^{t}_{x}=\nabla_{x}F_{\mathcal{D}}(\hat{x}^{t},\hat{y}^{t};B^{t})$, $g^{t}_{y}=-\nabla_{y}F_{\mathcal{D}}(\hat{x}^{t},\hat{y}^{t};B^{t})$
6:     $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau g^{t}_{x,j}\right),\quad\forall j\in[d_{x}]$  and  $y^{t+1}_{i}\propto y^{t}_{i}\exp\left(\tau g^{t}_{y,i}\right),\quad\forall i\in[d_{y}]$
7:  end for
8:  return  $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$, where $\tilde{x}^{t}\sim P_{x^{t}}$, $\tilde{y}^{t}\sim P_{y^{t}}$
Theorem 4.1 (Guarantees for Algorithm 1 with second order smoothness.).
Suppose the convex-concave objective $f(\cdot,\cdot;z)$ is second-order-smooth. Then there exists $T,\tau$, and $K$ such that Algorithm 1 returns $(\tilde{x},\tilde{y})$ with
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1}+L_{2}){\small\bigg{[}\sqrt{\frac{\log(d_{x})+\log(d_{y})}{n}}+\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{2/5}\bigg{]}.}$
Moreover, if $L_{2}=0$, then there exists $T,\tau$ and $K$ such that Algorithm 1 returns $(\tilde{x},\tilde{y})$ with
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\bigg{[}\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{n}}+\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg{]}.$
In both cases, the output $(\tilde{x},\tilde{y})$ is $(\varepsilon,\delta)$-DP.
Remark 4.1.
The duality gap bound for $L_{2}=0$ (i.e. quadratic objective) in Theorem 4.1 is optimal up to constant factors, at least in the case where $\log(d_{x})\simeq\log(d_{y})$. This follows from a lower bound for the problem of DP Synthetic Data Generation (details in Appendix G). We also note that the proof of this rate only requires $L_{2}$ to be ‘small’ (Remark E.1 for details), rather than exactly 00, so the optimal rates hold for more general nearly quadratic objectives.
Theorem 4.2 (Guarantees for Algorithm 1 with first order smoothness.).
Suppose the convex-concave objective $f(\cdot,\cdot;z)$ is first-order-smooth. Then there exists $T,\tau$ and $K$ such that Algorithm 1 returns an $(\varepsilon,\delta)$-DP point, $(\tilde{x},\tilde{y})$ with
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\left[\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{n}}+\bigg{(}\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/3}\bigg{]}.$
To prove these theorems, we will need some technical intermediate results (Appendix E.1 for proofs). The first is a condition on the stepsize that ensures the output $(\tilde{x},\tilde{y})$ is private.
Proposition 4.1.
Algorithm 1 is $(\varepsilon,\delta)$-DP if $\tau\leq B\varepsilon/[8L_{0}\sqrt{2T(K+1)\log(1/\delta)}].$
For the accuracy analysis, note that Algorithm 1 produces both a private sequence $(\tilde{x}^{t},\tilde{y}^{t})_{t\in[T]}$ and a non-private one $(x^{t},y^{t})_{t\in[T]}$. We first show a bound for the average of the non-private iterates.
Lemma 4.1 (Implicit in (Nemirovski et al., 2009)).
Let $(x^{t},y^{t})$ be the (non-private) iterates of Algorithm 1 generated in line 6, and let $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t}$ and $\bar{y}=\sum_{t\in[T]}y^{t}$. Furthermore, let $g^{t}_{x}$ and $g^{t}_{y}$ be the estimators constructed in line 5 of the algorithm. Then, the following inequality holds.
$T\operatorname{Gap}(\bar{x},\bar{y})$ $\leq\frac{2(\log(d_{x})+\log(d_{y}))}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2})$
$\qquad+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
where $g^{t}=(g^{t}_{x},g^{t}_{y})$ and the sequences $(w^{t})_{t\in[T]}$ and $(v^{t})_{t\in[T]}$ are completely determined by the randomness of the algorithm up to iteration $t$.
For explicit definitions of $(w^{t},v^{t})$ and the proof of Lemma 4.1, see Appendix E.1. Now we quantify the discrepancy between the gap of the private (sampled) and non-private iterates. Since the duality gap is defined in terms of function values, we would like to use Maurey’s Sparsification Lemma. However, since the duality gap involves a maximum –and the maximum of smooth functions is not necessarily smooth– the sparsification lemma does not directly apply. We handle this issue with a smoothing argument, and get the following lemma (see Appendix E.1 for proof).
Lemma 4.2.
Let $(\tilde{x},\tilde{y})$ be the output of Algorithm 1 and define $(\bar{x},\bar{y})$ as in Lemma 4.1. Then,
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{4L_{1}}{T}+\frac{2L_{1}\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{T}}.$
With these results, we are ready to prove Theorem 4.1.
Proof of Theorem 4.1.
Denote $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$ the private output of Algorithm 1, and $(\bar{x},\bar{y})=\frac{1}{T}\sum_{t=1}^{T}(x^{t},y^{t})$ its ‘non-private’ average. Using Lemma 4.1 and the fact that $\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2}\leq 5L_{0}^{2}$ by $L_{0}$-Lipschitzness of the objective we obtain
$\operatorname{Gap}(\bar{x},\bar{y})\leq\frac{2(\log(d_{x})+\log(d_{y}))}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
Since $w^{t},v^{t}$ are deterministic when conditioning on the randomness up to iteration $t$, taking expectations on both sides of the previous inequality leads to
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]$ $\lesssim\frac{2(\log(d_{x})+\log(d_{y}))}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}.$ (1)
Since we are assuming second order smoothness, then $|\mathbb{E}[\nabla_{x,j}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x,j}\mid\mathcal{F}_{t}]|\leq\frac{2L_{2}}{K}$ by Corollary 3.1. Similarly, for any $i\in[d_{y}]$ we have $|\mathbb{E}[-\nabla_{y,i}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y,i}\mid\mathcal{F}_{t}]|\leq 2L_{2}/K$. We conclude that $\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}\leq 2L_{2}/K$. In addition, from Lemma 4.2 it follows that $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq 8L_{1}\sqrt{\log(d_{x})+\log(d_{y})}/\sqrt{T}$. These bounds lead to
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\lesssim L_{1}\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{T}}+\frac{(\log(d_{x})+\log(d_{y}))}{\tau T}+\tau L_{0}^{2}+\frac{L_{2}}{K}.$ (2)
We obtain the first convergence rate by optimizing this bound over $T,\tau,L$, taking into account the constraint over $\tau$ that Proposition 4.1 requires for the privacy. ∎
4.2 Private Stochastic Mirror Descent with Bias-Reduced Sampling from Vertices
The limitation that prevents Algorithm 1 from attaining optimal rates is the bias of the stochastic gradient oracles. This is a consequence of sampling from vertices, and can only be mitigated by increasing the sampling rate. However, this eats up the privacy budget. In this section, we propose circumventing this limitation by combining our gradient oracles with a bias-reduction method.
Bias-Reduced Gradient Oracle.
We say a random variable $N$ is distributed as a truncated geometric with parameters $(p,M)$, denoted by $\mbox{TG}(p,M)$, if $\mathbb{P}[N=k]=p^{k}\mathbbm{1}_{(k\in\{0,1,...,M\})}/C_{M}$, where the normalization constant is $p\leq C_{M}\leq 1$. To construct a bias-reduced estimator of the gradient at $(x,y)$, we start by sampling $N\sim\mbox{TG}(0.5,M)$. Then, we construct a batch $B$ of $\max\{1,2^{N^{t}}/\alpha\}$ fresh samples, where $\alpha$ is a parameter to be choosen later. This batch $B$ combined with the gradient estimates (see Algorithm 2 and note similar algorithms in Blanchet and Glynn (2015)) reduces the bias while preserving privacy and accuracy (see Thm. 4.3). To do so, we control the bias and second moment of the gradient estimator. The proof of the next result is deferred to Appendix E.2.
Algorithm 2 BiasReducedGradient($x,y,N,B$)
[…]
  $(\bar{x}_{+},\bar{y}_{+})=\frac{1}{2^{N+1}}\sum_{i\in[2^{N+1}]}(\hat{x}^{i},\hat{y}^{i})$ ; $(\bar{x}_{-},\bar{y}_{-})=\frac{1}{2^{N}}\sum_{i\in[2^{N}]}(\hat{x}^{i},\hat{y}^{i})$ ; $(x_{0},y_{0})=(\hat{x}^{1},\hat{y}^{1})$
  $g_{x}=C_{M}2^{N}(\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B))+\nabla_{x}f(x_{0},y_{0};B)$
  $g_{y}=C_{M}2^{N}(\nabla_{y}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{y}f(\bar{x}_{-},\bar{y}_{-};B))+\nabla_{y}f(x_{0},y_{0};B)$
  return  $g_{x},g_{y}$
Lemma 4.3 (Bias and Second Moment of Algorithm 2).
[…]
$\|\mathbb{E}[\Phi_{\mathcal{D}}(x,y)-(g_{x},-g_{y})]\|_{\infty}$ $\leq\frac{2L_{2}}{2^{M}},$
[…]
Consequently, the bias scales inversely proportional to $2^{M}$, where $M$ is largest possible realization of $N$. Hence, the expected bias is the same as if we were sampling approximately $2^{M}$ times while we actually sample only $2^{N}$ times. This idea leads to a more favorable privacy-utility tradeoff. Another consequence is the order of the second moment, $M(\log(d_{x})+\log(d_{y}))$. Here $M$ will be choosen as a logarithmic quantity. Hence, the method allows us to improve the bias at a only a logarithmic cost in the second moment bound. Note, however, a disadvantage of this gradient oracle is we can only control the second moment for second-order-smooth objective functions.
Bias-reduced Algorithm, Algorithm 3.
We propose an algorithm which is essentially Algorithm 1 combined with the bias-reduced gradient oracle. The main challenges encountered are the oracles becoming heavy-tailed (moments are exponentially larger than their expectation), and due to a randomization of batch sizes, it is difficult to control the privacy budget. To handle the heavy-tailed nature of our estimators, we resort to efficiency estimates from Markov’s inequality which only certify success with constant probability. On the other hand, the privacy budget analysis uses the Fully Adaptive Composition Theorem of DP (Whitehouse et al., 2023), which allows us in particular to set up the mechanisms and their privacy parameters in a random (but predictable) way.
Algorithm 3 Private stochastic mirror descent with bias-reduced sampling from vertices
0:  Dataset $S$, Step-size $\tau$, Stopping parameter $U\geq 1$, TG parameter $M$, Batch parameter $\alpha$.
  $x^{1}=(1/d_{x},...,1/d_{x})$, $y^{1}=(1/d_{y},...,1/d_{y}),t=1,N^{1}\sim\mbox{TG}(0.5,M)$
  while $\sum_{i\in[t-1]}2^{N^{i}}\leq U-2^{M}$ do
     Let $B^{t}$ be a fresh batch of $\max\{1,2^{N^{t}}/\alpha\}$ samples.
     $(g^{t}_{x},g^{t}_{y})=\operatorname{BiasReducedGradient}(x^{t},y^{t},N^{t},B^{t})$
     $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau g^{t}_{x,j}\right),\quad\forall j\in[d_{x}]$  and  $y^{t+1}_{i}\propto y^{t}_{i}\exp\left(\tau g^{t}_{y,i}\right),\quad\forall i\in[d_{y}]$
     $N^{t+1}\sim\mbox{TG}(0.5,M)$and set $t=t+1$
  end while
  return  $(\tilde{x},\tilde{y})=\frac{1}{t}\sum_{i=1}^{t}(\tilde{x}^{i},\tilde{y}^{i})$ where $\tilde{x}^{i}\sim P_{x^{i}}$, $\tilde{y}^{i}\sim P_{y^{i}}$
Theorem 4.3 (Accuracy and Privacy for Algorithm 3).
Suppose the objective $f(\cdot,\cdot;z)$ is convex-concave and second-order-smooth. Let $(\tilde{x},\tilde{y})$ be the output of Algorithm 3 with $M=\log(\sqrt{U})$, stopping parameter $U=\min\left\{\frac{n\varepsilon\sqrt{L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}}}{\sqrt{4\cdot 48\cdot 81(\log(d_{x})+\log(d_{y}))\log(1/\delta)}L_{0}},\frac{n}{2}\right\}$, $\alpha=\left(\frac{2\varepsilon^{2}}{48\cdot 81\log(1/\delta)(\tau L_{0})^{2}n}\right)^{1/3}$ and step-size $\tau=\sqrt{\frac{(\log(d_{x})+\log(d_{y}))}{(L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2})U}}$. If $U\geq 4$, then $(\tilde{x},\tilde{y})$ is $(\varepsilon,\delta)$-DP and
$\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\lesssim\sqrt{\frac{[(\log(d_{x})+\log(d_{y})][L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]M^{2}}{U}}\,\right]\geq 0.99.$
Many ideas from the analysis of Algorithm 1 are borrowed to prove Theorem 4.3. There are two main differences that make the analysis considerably more involved. First, our algorithm runs with random batch sizes and for a random number of steps, so the Adaptive Composition of DP is not useful anymore. Instead, we prove the privacy guarantees using the Fully Adaptive Composition Theorem of DP (Whitehouse et al., 2023). Secondly, our main tool to bound the duality gap is Lemma 4.1, which involves a sum of the second moments of the bias-reduced gradient estimators and of the bias terms. We control these errors in expectation w.r.t. the randomness in the truncated geometric random variables. But then, we sum these terms over a random number of iterations that depends on the realization of truncated geometric random variables. Hence, it is not clear that Lemma 4.3 will be directly applicable. To handle this, we borrow tools from the literature of stopping times and prove that, indeed, Lemma 4.3 is applicable.
Theorem 4.3 only provides guarantees with constant probability because the number of iterations depends on the realization of the truncated geometric random variables, which are, themselves, heavy tailed. In short, the number of steps will be in expectation close to $U$, but the heavy tailed nature of the (exponential of) truncated geometric random variables prevents in expectation or high probability guarantees. The detailed proof of Theorem 4.3 can be found in Appendix E.2.
Algorithm 4 Private stochastic mirror descent with sequential Maurey sparsification for DPSCO
0:  Dataset $S=\{z^{1},...,z^{n}\}$, Total number of steps $T$, Step-size $\tau$, Sampling parameter $K$, Number of steps for updating private average iterate $q$, Initial point $x^{1}=(1/d_{x},...,1/d_{x})$
1:  for $t=1$ to $T$ do
2:     $w^{t}=\frac{(t-1)w^{t-1}+x^{t}}{t}$
3:     if $t\leq q$ or $t=0$ mod $q$ then
[…]
5:     else
6:        $\hat{w}^{t}=\hat{w}^{t-1}$
7:     end if
8:     $g^{t}=\nabla f(\hat{w}^{t};B^{t})$, where $B^{t}$ is a batch of $B=n/T$ fresh samples.
9:     $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau g^{t}_{j}\right),\quad\forall j\in[d_{x}]$
10:  end for
[…]
5 Differentially Private Stochastic Convex Optimization
[…]
Lemma 5.1 (Anytime online-to-batch Cutkosky (2019)).
Let $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ be a convex set. For any sequence $\beta_{t}\in\mathbb{R}^{+},g^{t}\in\mathbb{R}^{d_{x}}$, suppose an online learner predicts $x^{t}$ and receives $t$-th loss $\ell_{t}(x)=\langle g^{t},x\rangle$. Define $w^{t}=\sum_{i\in[t]}\frac{\beta_{i}x^{i}}{\sum_{j\in[t]}\beta_{j}}$. Then, for any convex, differentiable $F$ over $\mathcal{X}$ and any $x\in\mathcal{X},$
$F(w^{T})-F(x)\leq\frac{1}{\sum_{j\in[T]}\beta_{j}}\bigg{(}\operatorname{Regret}_{T}(x)+\sum_{t\in[T]}\langle\beta_{t}\nabla F(w^{t})-g_{t},x_{t}-x\rangle\bigg{)}.$
An advantage of this conversion is that the weighted averages change slowly from one iteration to the next, which is very useful in the smooth case (Zhang et al., 2022b). Using Lemma 5.1 with weights $\beta_{j}=1$ to optimize over $\Delta_{x}$, the average iterates satisfy $\|w^{t}-w^{t-1}\|_{1}=\|x^{t}-w^{t-1}\|_{1}/t\leq 2/t$. Consequently, $P_{w^{t-1}}$ and $P_{w^{t}}$ are (eventually) similar distributions. We exploit this idea in Algorithm 4 by maintaining estimates of the private iterates for a large number of steps, aggressively reducing the number of samples and improving the privacy. Theorem 5.1 provides a faster rate compared to the one we would get with the plain online-to-batch conversion. Proofs in Appendix F.
Theorem 5.1 (Guarantees for Algorithm 4).
Let $f(\cdot;z)$ be convex and first-order-smooth. Then, there exists a parameter setting for $T,K,q,\tau$ such that Algorithm 4 run with these as inputs returns $\hat{w}^{T}$ such that $\hat{w}^{T}$ is $(\varepsilon,\delta)$-DP and
$\mathbb{E}[F_{\cal D}(\hat{w}^{T})-F_{\cal D}(x)]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\log(d_{x})}{n}}+L_{0}\frac{\log(d_{x})^{7/10}\log(1/\delta)^{1/5}}{(n\varepsilon)^{2/5}}\right]\\ +L_{1}\log(\sqrt{n}\log(d_{x}))\left[\frac{1}{\log(d_{x})\sqrt{n}}+\frac{\log(1/\delta)^{1/5}}{\log(d_{x})^{4/5}(n\varepsilon)^{2/5}}\right].$
If $f(\cdot;z)$ is convex and second-order-smooth , Algorithm 4 with a different choice for $T,K,q,\tau$ outputs $\hat{w}^{T}$ that is $(\varepsilon,\delta)$-DP and satisfies
$\mathbb{E}[F_{\cal D}(\hat{w}^{T})-F_{\cal D}(x)]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\log(d_{x})}{n}}+\frac{\log(d_{x})\log(1/\delta)^{1/4}}{\sqrt{n\varepsilon}}\right]+\frac{L_{2}}{\sqrt{n}\log(d_{x})^{1/4}}\\ +\frac{L_{2}\log(d_{x})^{1/4}\log(1/\delta)^{1/4}}{\sqrt{n\varepsilon}}+L_{1}\log(\sqrt{n\log(d_{x})})\Big{[}\frac{1}{\sqrt{n}}+\frac{\log(1/\delta)^{1/4}}{\sqrt{n\varepsilon}}\Big{]}.$
Conclusion.
We presented mirror descent based algorithms for DP-SSP in the polyhedral setting. Our algorithms are the first achieving polylogarithmic-in-the-dimension error beyond bilinear objectives. In fact, for nearly quadratic objectives our rates are optimal. For second-order-smooth objectives, we showed our algorithms, in combination with bias-reduction techniques, achieve (up to logarithmic factors) the optimal rates with constant success probability. An important question remains: Are the optimal rates achievable in expectation for first and second-order-smooth functions? We leave this question for future work. We also apply our ideas to DP-SCO in the polyhedral setting, where we also leave open the problem of achieving optimal rates with mirror descent based algorithms.
Acknowledgments.
T. González’s research was partially done as a Research Intern at Google Brain. C. Guzmán’s research was partially done as a Visiting Researcher at Google. C. Guzmán’s research was partially supported by INRIA Associate Teams project, ANID FONDECYT 1210362 grant, ANID Anillo ACT210005 grant, and National Center for Artificial Intelligence CENIA FB210017, Basal ANID. C. Paquette is a Canadian Institute for Advanced Research (CIFAR) AI chair, Quebec AI Institute (MILA) and C. Paquette was supported by a Discovery Grant from the Natural Science and Engineering Research Council (NSERC) of Canada, NSERC CREATE grant Interdisciplinary Math and Artificial Intelligence Program (INTER-MATH-AI)”, an external grant from Google, and Fonds de recherche du Québec – Nature et technologies (FRQNT) New University Researcher’s Start-Up Program.
Appendix A Auxiliary results for privacy and convergence analysis
For completeness, we state some important known results for privacy and convergence analysis.
Theorem A.1 (Advanced composition (Corollary 3.21 in Dwork et al. [2014])).
Given target privacy parameters $10<\varepsilon^{\prime}<$ 1 and $\delta^{\prime}>0$, to ensure $(\varepsilon^{\prime},T\delta+\delta^{\prime})$ cumulative privacy loss over the adaptive composition of $T$ mechanisms, it suffices that each mechanism is $(\varepsilon,\delta)$-DP, where $\varepsilon=\frac{\varepsilon^{\prime}}{2\sqrt{2T\log(1/\delta^{\prime})}}$.
Theorem A.2 (Fully Adaptive Composition Whitehouse et al. [2023]).
Suppose $(\mathcal{A}_{n})_{n\geq 1}$ is a sequence of algorithms such that, for any $n\geq 1$, $\mathcal{A}_{n}$ is $(\varepsilon_{n},\delta_{n})$-DP with respect to a dataset $S$ when conditioned on $\mathcal{A}_{1:n-1}$. Let $\varepsilon\geq 0$ and $\delta=\delta^{\prime}+\delta^{\prime\prime}$ be the target privacy parameters such that $\delta^{\prime},\delta^{\prime\prime}>0$. Consider the function $N:\mathbb{R}^{\infty}_{\geq 0}\times\mathbb{R}^{\infty}_{\geq 0}\to\mathbb{N}$ given by
$N((\varepsilon_{n})_{n\geq 1},(\delta_{n})_{n\geq 1})=\inf\left\{n:\varepsilon<\sqrt{2\log\left(\frac{1}{\delta^{\prime}}\right)\sum_{m\leq n+1}\varepsilon_{m}^{2}}+\frac{1}{2}\sum_{\sum_{m\leq n+1}}\varepsilon_{m}^{2}\text{ or }\delta^{\prime\prime}<\sum_{m\leq n+1}\delta_{m}\right\}.$
Then, the algorithm $\mathcal{A}_{1:N(S)}$ is $(\varepsilon,\delta)$-DP, where $N(S)=N((\varepsilon_{n}(S))_{n\geq 1},(\delta_{n}(S))_{n\geq 1})$.
Theorem A.3 (Exponential mechanism (Theorem 3.10 in Dwork et al. [2014])).
Let $J:=\{e_{1},...,e_{d}\}$. Consider a (score) function $s:{\cal Z}^{n}\times J\mapsto\mathbb{R}$. Define the sensitivity of the (score) function as $\Delta_{s}:=\max_{j\in J}\max_{S\sim S^{\prime}}|s(S,j)-s(S^{\prime},j)|$. Then, sampling an object $j$ from $J$ according to $\mathbb{P}(j=e_{i})\propto\exp\left(\frac{\varepsilon s(S,i)}{2\Delta_{s}}\right)$ is $\varepsilon$-DP. As a consequence, sampling according to $\mathbb{P}(j=e_{i})\propto\exp\left(s(S,i)\right)$ is $2\Delta_{s}-$DP.
Lemma A.1 (Regret guarantees of SMD, Implicit in [Nemirovski et al., 2009]).
For every $x\in\Delta_{d}$. After $T$ iterations of Online Multiplicative Weights, we obtain
$\operatorname{Regret}_{T}(x)\leq\frac{\log(d)}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2},$
where $g^{t}\in\partial f_{t}(x)$.
Lemma A.2 (Vector martingale concetration in $\ell_{\infty}$ norm Juditsky and Nemirovski [2008]).
Let $d\geq 3$ and $\{x^{t}\}$ be a sequence of random vectors such that $\mathbb{E}[\|x^{t}\|_{\infty}]<\infty$ and $\mathbb{E}[x_{t+1}|x_{1:t}]=0$ for all $t$. Then, for a constant $0<c\leq 2e$, we have
$\mathbb{E}\Big{[}\big{\|}\sum_{t=1}^{T}x^{t}\Big{\|}_{\infty}^{2}\Big{]}\leq c\log(d)\sum_{t=1}^{T}\mathbb{E}[\|x^{t}\|_{\infty}^{2}].$
Lemma A.3 (Corollary 2 in Juditsky et al. [2011]).
Let $\xi^{1},\xi^{2},...$ be a sequence in $\mathbb{R}^{d}$ and $\mathcal{X}$ be a compact, convex set. Define the sequence $w^{1}\in\mathbb{R}^{d}$, $w^{t+1}=\operatorname{argmin}_{x\in\mathcal{X}}D_{\psi}(w^{t},x)+\langle\xi^{t},x\rangle$. Then, we have that for all $w\in\mathcal{X}$ and all iterations $T$:
$\sum_{t\in[T]}\langle\xi^{t},w^{t}-w\rangle\leq D_{\psi}(w^{1},w)+\frac{1}{2}\sum_{t\in[T]}\|\xi^{t}\|^{2}_{*}$
Appendix B Motivating Examples
We provide two motivating examples for our setting.
Private Synthetic Data Generation
[…]
$\min_{x\in\Delta_{|\mathcal{Z}|}}\max_{y\in\Delta_{|\mathcal{Q}|}}\mathbb{E}_{z\sim\mathcal{D}}\Big{[}\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)\Big{]}.$ (3)
While the algorithm we propose in Section 4.1 can solve this problem, applying them directly would result in a slightly suboptimal dependence on the histogram size, $|{\cal Z}|$. We propose instead a variant of Algorithm 1 that leverages the specific structure of this problem. This algorithm attains the optimal rates for this problem [Bun et al., 2014]. We defer the details to Appendix G.
Minimizing the maximal loss
Consider a collection of convex losses $f_{i}:\mathcal{X}\times\mathcal{Z}\to[-B,B]$, $i\in[d_{y}]$. Suppose $\mathcal{X}$ is a polytope. We are interested in minimizing the maximal expected loss $\max_{i\in[d_{y}]}\{F_{i}(x):=\mathbb{E}_{z\sim\mathcal{D}}[f_{i}(x;z)]\}$. This can be formulated as follows:
$\min_{x\in\mathcal{X}}\{\max_{i\in[d_{y}]}F_{i}(x)\}=\min_{x\in\mathcal{X}}\max_{y\in\Delta_{y}}\mathbb{E}_{z\sim\mathcal{D}}\left[\sum_{i\in[d_{y}]}y_{i}f_{i}(x;z)\right].$
If the functions $(f_{i})_{i\in[d_{y}]}$ are $L_{0}$-Lipschitz and $L_{1}$-smooth on $x$ w.r.t $\|\cdot\|_{1}$, then the problem objective $f(x,y;z)=\sum_{i\in[d_{y}]}y_{i}f_{i}(x;z)$ is convex-concave, $\max\{L_{0},B\}$-Lipschitz, $\max\{L_{0},L_{1}\}$-smooth on $(x,y)$. If the functions $(f_{i})_{i\in[d_{y}]}$ are also $L_{2}$-second-order-smooth on $x$, then $f(x,y)$ is $\max\{L_{1},L_{2}\}$-second-order-smooth on $(x,y)$. Examples of convex losses that are Lipschitz, smooth and second-order-smooth are the quadratic and the logistic loss. Finally, the maximal loss formulation that we present encompasses applications to fairness [Jagielski et al., 2019, Li et al., 2019, Lowy et al., 2023] and robustness [Mohajerin Esfahani and Kuhn, 2018] under privacy constraints.
Appendix C On the Generality of the Simplex for Polyhedral Setups
In previous works [Asi et al., 2021b, Bassily et al., 2021a], polyhedral setups are described in more generality than the standard simplex as feasible set. Here we show that in fact our formulation directly extends to those cases, and therefore our assumptions are without loss of generality.
Consider a DP-SSP problem
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}F(x,y)$ (4)
where
  • ${\cal X}\subseteq\mathbb{R}^{d_{x}}$ and ${\cal Y}\subseteq\mathbb{R}^{d_{y}}$ are polytopes with $K_{x}$ and $K_{y}$ vertices respectively, that we denote by $\{x_{1},\ldots,x_{K_{x}}\}$ and $\{y_{1},\ldots,y_{K_{y}}\}$, respectively. We denote by $D$ the $\|\cdot\|_{1}$-diameter of $\mathcal{X}\times\mathcal{Y}$.
  • $F$ is $L_{0}^{F}$-Lipschitz, $L_{1}^{F}$-smooth w.r.t $\|\cdot\|_{1}$ and has $L_{2}^{F}$-smooth w.r.t $\|\cdot\|_{1}$ partial derivatives.
We consider the linear transformation
[…]
In other words, $\Lambda$ is a matrix whose $i$-th column is $(x_{i}-x_{1},\mathbf{0}_{d_{y}})$ for $i\in[K_{x}]$ and $(\mathbf{0}_{d_{x}},y_{i}-y_{1})$ for $i\in\{K_{x}+1,...,K_{x}+K_{y}\}$, where $\mathbf{0}_{d}$ is the vector of all zeros in $\mathbb{R}^{d}$. We reformulate problem (4) as
$\min_{\lambda\in\Delta_{K_{x}}}\max_{\mu\in\Delta_{K_{y}}}[G(\lambda,\mu):=F(\Lambda(\lambda,\mu)+(x_{1},y_{1}))].$
Naturally, stochastic oracles can be computed for $G$ based on those for $F$ and on $\Lambda$. Next, we show how the constants for $G$ can be bounded in terms those of $F$. First, note that
$|G(\lambda_{1},\mu_{1})-G(\lambda_{2},\mu_{2})|$ $=|F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))|$
$\leq L_{0}^{F}\|\Lambda(\lambda_{1},\mu_{1})-\Lambda(\lambda_{2},\mu_{2})\|_{1}$
$=L_{0}^{F}\left\|\left(\sum_{i=1}^{K_{x}}(\lambda_{1,i}-\lambda_{2,i})(x_{i}-x_{1}),\sum_{j=1}^{K_{y}}(\mu_{1,j}-\mu_{2,j})(y_{j}-y_{1})\right)\right\|_{1}$
$\leq L_{0}^{F}\sum_{i=1}^{K_{x}}|\lambda_{1,i}-\lambda_{2,i}|\|x_{i}-x_{1}\|_{1}+L_{0}\sum_{j=1}^{K_{y}}|\mu_{1,j}-\mu_{2,j}|\|y_{j}-y_{1}\|_{1}$
$\leq L_{0}^{F}D(\|\lambda_{1}-\lambda_{2}\|_{1}+\|\mu_{1}-\mu_{2}\|_{1})$
$=L_{0}^{F}D\|(\lambda_{1},\mu_{1})-(\lambda_{2},\mu_{2})\|_{1}.$
Next,
$\|\nabla G(\lambda_{1},\mu_{1})-\nabla G(\lambda_{2},\mu_{2})\|_{\infty}$
$=\big{\|}\Lambda^{T}\big{[}\nabla F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\big{]}\big{\|}_{\infty}$
$=\max_{i\in[K_{x}+K_{y}]}|\langle\Lambda^{T}_{i},\nabla F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\rangle|$
$\leq\max_{i\in[K_{x}+K_{y}]}\|\Lambda^{T}_{i}\|_{1}\|\nabla F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\|_{\infty}$
$\leq\max_{i\in[K_{x}+K_{y}]}\|\Lambda^{T}_{i}\|_{1}L_{1}^{F}\|\Lambda(\lambda_{1},\mu_{1})-\Lambda(\lambda_{2},\mu_{2})\|_{1}$
$\leq L_{1}^{F}D^{2}\|(\lambda_{1},\mu_{1})-(\lambda_{2},\mu_{2})\|_{1}.$
where in the last line we used that the rows of $\Lambda^{T}$ are the columns of $\Lambda$, and they have the form $x_{i}-x_{1}$ or $y_{i}-y_{1}$, so their $\ell_{1}$ norm is at most $D$. We also used that $\|\Lambda(\lambda_{1},\mu_{1})-\Lambda(\lambda_{2},\mu_{2})\|_{1}\leq D\|(\lambda_{1},\mu_{1})-(\lambda_{2},\mu_{2})\|_{1}$, which was implictly proved when we showed that $G$ is $L_{0}D$-Lipschitz. Finally, notice that
$\|\nabla\nabla_{j}G(\lambda_{1},\mu_{1})-\nabla\nabla_{j}G(\lambda_{2},\mu_{2})\|_{\infty}=\max_{k\in[K_{x}+K_{y}]}\big{|}\nabla_{k,j}G(\lambda_{1},\mu_{1})-\nabla_{k,j}G(\lambda_{2},\mu_{2})\big{|}$
$=\max_{k\in[K_{x}+K_{y}]}\big{|}\nabla_{k}\langle\Lambda^{T}_{j},\nabla F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\rangle\big{|}$
$=\max_{k\in[K_{x}+K_{y}]}\left|\sum_{i\in[K_{x}+K_{y}]}\Lambda^{T}_{j,i}\Big{[}\nabla_{k,i}F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla_{k,i}F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\Big{]}\right|$
$=\max_{k\in[K_{x}+K_{y}]}\left|\sum_{i\in[K_{x}+K_{y}]}\Lambda^{T}_{j,i}\langle\Lambda^{T}_{k},\nabla\nabla_{i}F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla\nabla_{i}F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\right|$
$\leq\max_{k\in[K_{x}+K_{y}]}\sum_{i\in[K_{x}+K_{y}]}|\Lambda_{j,i}|\|\Lambda^{T}_{k}\|_{1}\|\nabla\nabla_{i}F(\Lambda(\lambda_{1},\mu_{1})+(x_{1},y_{1}))-\nabla\nabla_{i}F(\Lambda(\lambda_{2},\mu_{2})+(x_{1},y_{1}))\|_{\infty}$
$\leq\max_{k\in[K_{x}+K_{y}]}\|\Lambda^{T}_{k}\|_{1}\|\Lambda^{T}_{j}\|_{1}L_{2}^{F}\|\Lambda(\lambda_{1},\mu_{1})-\Lambda(\lambda_{2},\mu_{2})\|_{1}$
$\leq L_{2}^{F}D^{3}\|(\lambda_{1},\mu_{1})-(\lambda_{2},\mu_{2})\|_{1}.$
where we use the same arguments we used for Lipschitzness and smoothness. We conclude that $G$ is $L_{0}^{G}$-Lipschitz, $L_{1}^{G}$-smooth with $L_{2}^{G}$-smooth partial derivatives for some $L_{0}^{G}\leq L_{0}^{F}D$, $L_{1}^{G}\leq L_{1}^{F}D^{2}$ and $L_{2}^{G}\leq L_{2}^{F}D^{3}$.
Appendix D Missing results from Section 3
We state explicitly the bounds for the bias of gradients.
Corollary D.1 (Gradient bias with second-order-smoothness).
Suppose $F:\mathbb{R}^{d}\to\mathbb{R}$ has partial derivates that are $L_{2}$-smooth w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors in $\Delta_{d}$ and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for every $t\in[T]$. Then
$\left\|\mathbb{E}\left[\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right\|_{\infty}\leq\frac{2L_{2}}{T}.$
Proof.
This follows directly from applying Corollary 3.1 to each partial derivate: for all $j\in[d]$, we have
$\left|\mathbb{E}\left[\nabla_{j}F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-\nabla_{j}F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right|\leq\frac{2L_{2}}{T}.$
Corollary D.2 (Gradient bias with first-order-smoothness).
Suppose $F:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors in $\Delta_{d}$ and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for every $t\in[T]$. Then
$\left\|\mathbb{E}\left[\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right\|_{\infty}\leq\frac{4L_{1}}{\sqrt{T}}.$
Proof.
Denote $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t},\bar{a}=\frac{1}{T}\sum_{t\in[T]}a^{t}$. First, note that by $L_{1}$-smoothness of $F$ we have for all $j\in[d]$:
$F(\bar{x}+re_{j})\leq F(\bar{x})+\langle\nabla F(\bar{x}),re_{j}\rangle+\frac{L_{1}r^{2}}{2}\|e_{j}\|_{1}^{2},$
for all $r\in\mathbb{R}^{+}$. Re-arranging this expression
$\langle\nabla F(\bar{x}),e_{j}\rangle\leq\frac{F(\bar{x})-F(\bar{x}+re_{j})}{r}+\frac{L_{1}r}{2},$
and similarly
$\langle-\nabla F(\bar{a}),e_{j}\rangle\leq\frac{F(\bar{a}+re_{j})-F(\bar{a})}{r}+\frac{L_{1}r}{2}.$
Adding these inequalities and taking expectations we get
$\mathbb{E}[\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle]\leq\frac{\mathbb{E}\left[F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})\right]}{r}+L_{1}r.$
By applying Corollary 3.1 to $F(\cdot)$ and $F(\cdot+re_{j})$ we obtain $\mathbb{E}\left[F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})\right]\leq 4L_{1}/T$. Hence, for all $j\in[d]$
$\mathbb{E}[\nabla_{j}F(\bar{x})-\nabla_{j}F(\bar{a})]=\mathbb{E}[\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle]\leq\frac{4L_{1}}{rT}+L_{1}r.$
Choosing $r=2/\sqrt{T}$ gives the result.
Now, we prove Lemma 3.2 and Lemma 3.3.
Lemma 3.2.
Proceeding as in the proof of Lemma 3.1, we get
$F(\bar{a}^{T})-F\left(\bar{x}^{T}\right)\leq\sum_{t=1}^{T}\alpha_{t}+\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2},$
where $\alpha_{t}=\lambda_{t}\left\langle\nabla F\left(\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\right),a^{t}-x^{t}\right\rangle$. Hence
$\mathbb{P}[F(\bar{a}^{T})-F\left(\bar{x}^{T}\right)\geq\alpha]\leq\mathbb{P}\left[\sum_{t=1}^{T}\alpha_{t}+\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\geq\alpha\right]=\mathbb{P}\left[\sum_{t\in[T]}\alpha_{t}\geq\alpha^{\prime}\right]$
for $\alpha=\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\alpha^{\prime}$. Next, we bound the probability on the right-hand side by using the bounded differences inequality. Indeed, note that $\{\alpha_{i}\}_{i\in[T]}$ is a martingale difference sequence wrt to $\{a^{i}\}_{i\in[T]}$ (we argued in the proof of Lemma 3.1 that $\alpha_{t}$ conditioned on $a^{1},...,a^{t-1}$ is centered), and furthermore $|\alpha_{i}|\leq L_{0}D\lambda_{i}$. Thus
$\mathbb{P}\left[\sum_{t=1}^{T}\alpha_{t}\geq\beta\frac{L_{0}D}{\sqrt{2}}\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\right]\leq\exp(-\beta^{2}),$
and we conclude that
$\mathbb{P}\left[F(\bar{a}^{T})-F(\bar{x}^{T})\geq\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\beta\frac{L_{0}D}{\sqrt{2}}\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\right]\leq\exp(-\beta^{2}).$
Lemma 3.3.
Define the constants $A=\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}$ and $C=\frac{L_{0}D}{\sqrt{2}}\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}$. We will introduce a quantity, $B$, in the following computation, and later optimize over the choice of this $B$. Hence,
[…]
$=\frac{(A+B)^{2}}{2}+2MC\left[C\exp\left(-(B/C)^{2}\right)+A\mathbb{P}[X\geq B/C]\right]$
where the first inequality comes from the union bound, the second from Lemma 3.2, and in the last line, we use density of a normal random variable, $X\sim{\cal N}(0,1/2)$. From standard gaussian tail bounds, we have $\mathbb{P}[X\geq B/C]\leq\exp\left(-(B/C)^{2})\right)$. Hence,
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{x})-F_{j}(\bar{a}^{T})|^{2}\right]\leq\frac{(A+B)^{2}}{2}+2MC(C+A)\exp\left(-(B/C)^{2}\right).$
Setting $B=C\sqrt{\log(M)}$ gives
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{x})-F_{j}(\bar{a}^{T})|^{2}\right]$ $\leq A^{2}+C^{2}\log(M)+2C^{2}+2AC$
$\leq 5A^{2}+(6+\log(M))C^{2}$
$=\frac{5L_{1}^{2}D^{4}}{4}\left(\sum_{t=1}^{T}\lambda_{t}^{2}\right)^{2}+\frac{L_{0}^{2}D^{2}(6+\log(M))}{2}\sum_{t=1}^{T}\lambda_{t}^{2}.$
Finally, we present two results regarding gradient approximation error that follow from Lemma 3.3.
Corollary D.3 (Gradient approximation error with second-order-smoothness).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$ with $L_{2}$-smooth partial derivatives. Let $x^{1},...,x^{T}\in\Delta_{d}$ be a collection of deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for every $t\in[T]$. Then
$\mathbb{E}\left\|\nabla F\left(\frac{1}{T}\sum_{t\in[T]}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t\in[T]}x^{t}\right)\right\|^{2}_{\infty}\leq\frac{20L_{2}^{2}}{T^{2}}+\frac{8L_{1}^{2}(6+\log(d))}{T}.$
Proof.
Applying Lemma 3.3 with $F_{j}=\nabla_{j}F$, $j\in[d]$, leads to directly to the result. ∎
Corollary D.4 (Gradient approximation error with first-order-smoothness).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a function that is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}\in\Delta_{d}$ be a collection of deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for every $t\in[T]$. Then
$\mathbb{E}\left\|\nabla F\left(\frac{1}{T}\sum_{t\in[T]}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t\in[T]}x^{t}\right)\right\|^{2}_{\infty}\leq\frac{43L_{1}^{2}}{\sqrt{12+\log(d)}T^{3/2}}+\frac{17(L_{0}^{2}+L_{1}^{2})\sqrt{(12+\log(d))}}{\sqrt{T}}.$
Proof.
Denote $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t},\bar{a}=\frac{1}{T}\sum_{t\in[T]}a^{t}$. Similarly to the proof of Corollary D.2, using $L_{1}$-smoothness of $F$,
$\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle$ $\leq\frac{F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})}{r}+L_{1}r$
$\leq\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})|+|F(\bar{x})-F(\bar{a})|}{r}+L_{1}r$
Since $-F$ is also $L_{1}$-smooth, then
$\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle\leq\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})|+|F(\bar{x})-F(\bar{a})|}{r}+L_{1}r.$
We conclude that
$\left\|\nabla F\left(\bar{a}\right)-\nabla F\left(\bar{x}\right)\right\|^{2}_{\infty}$ $=\max_{j\in[d]}\left|\nabla_{j}F\left(\bar{a}\right)-\nabla_{j}F\left(\bar{x}\right)\right|^{2}=\max_{j\in[d]}\left|\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle\right|^{2}$
$\leq 3\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})|^{2}+|F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}+L_{1}^{2}r^{2}$
$=3\left(\frac{|F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}+L_{1}^{2}r^{2}+\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})|^{2}}{r^{2}}\right)$
By Lemma 3.3 with $F_{j}(\cdot)=F(\cdot+re_{j})$ and $M=d$, we have
$\mathbb{E}\left[\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})|^{2}}{r^{2}}\right]\leq\frac{20L_{1}^{2}}{T^{2}r^{2}}+\frac{8L_{0}^{2}(6+\log(d))}{Tr^{2}}$
Similarly, by Lemma 3.3 with $F_{1}(\cdot)=F(\cdot)$ and $M=1$, we have
$\mathbb{E}\left[\frac{|F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}\right]\leq\frac{20L_{1}^{2}}{T^{2}r^{2}}+\frac{48L_{0}^{2}}{Tr^{2}}.$
Putting everything together and taking expectations
$\mathbb{E}[\left\|\nabla F\left(\bar{a}\right)-\nabla F\left(\bar{x}\right)\right\|^{2}_{\infty}]\leq\frac{120L_{1}^{2}}{T^{2}r^{2}}+\frac{24L_{0}^{2}(12+\log(d))}{Tr^{2}}+3L_{1}^{2}r^{2}.$
Choosing $r^{2}=\sqrt{\frac{8(12+\log(d))}{T}}$
$\mathbb{E}[\left\|\nabla F\left(\bar{a}\right)-\nabla F\left(\bar{x}\right)\right\|^{2}_{\infty}]\leq\frac{43L_{1}^{2}}{\sqrt{12+\log(d)}T^{3/2}}+\frac{17(L_{0}^{2}+L_{1}^{2})\sqrt{(12+\log(d))}}{\sqrt{T}}.$
Appendix E Missing results from Section 4
E.1 Missing results from Section 4.1
Here we provide the missing results of the analysis presented in Section 4. We start with a remark, which is easily observed from the proof of Theorem 4.1.
Remark E.1.
In Theorem 4.1, we can replace the condition $L_{2}=0$ and get the same convergence rate for the gap if $L_{2}\lesssim(L_{0}+L_{1})\left\{\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{n}}+\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\right\}$. In particular, this indicates that the optimal rates are achievable for a broader class than only quadratic functions.
We continue by giving the explicit parameter setting to get the rates in Theorem 4.1: If
$L_{2}\lesssim(L_{0}+L_{1})\left\{\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{n}}+\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\right\},$
we set the following parameters
$T\eqsim\min\left\{n,\frac{n\varepsilon}{\sqrt{(\log(d_{x})+\log(d_{y}))\log(1/\delta)}}\right\},$
$\tau\eqsim\min\left\{\frac{1}{L_{0}}\sqrt{\frac{\log(d_{x})+\log(d_{y})}{T}},\frac{n\varepsilon}{L_{0}T\sqrt{TK\log(1/\delta)}}\right\},K=1.$
Otherwise, we choose the following
$T\eqsim\min\left\{n,\left[\frac{n\varepsilon}{(\log(d_{x})+\log(d_{y}))^{1/4}\sqrt{\log(1/\delta)}}\right]^{4/5}\right\},$
$\tau\eqsim\min\left\{\frac{1}{L_{0}}\sqrt{\frac{\log(d_{x})+\log(d_{y})}{T}},\frac{n\varepsilon}{L_{0}T\sqrt{TK\log(1/\delta)}}\right\},K\eqsim\sqrt{\frac{T}{\log(d_{x})+\log(d_{y})}}.$
Next, we present the parameter setting to get the rates in Theorem 4.2:
$T\eqsim\min\left\{n,\frac{(n\varepsilon)^{2/3}}{\log(1/\delta)^{1/3}}\right\},$
$\tau\eqsim\min\left\{\frac{1}{L_{0}}\sqrt{\frac{\log(d_{x})+\log(d_{y})}{T}},\frac{n\varepsilon}{L_{0}T\sqrt{TK\log(1/\delta)}}\right\},K\eqsim\frac{T}{\log(d_{x})+\log(d_{y})}.$
Below, we prove Proposition 4.1, which gives a condition on the stepsize that guarantees the privacy of the algorithm.
Proposition 4.1.
We start proving that for all $t\in[T],k\in[K]$, $\tilde{x}^{t+1},\tilde{y}^{t+1}$ and $\hat{x}^{t+1,k},\hat{y}^{t+1,k}$ are private given the previous private iterates $\hat{x}^{1},\tilde{x}^{1},\hat{y}^{1},\tilde{y}^{1},...,\hat{x}^{t},\tilde{x}^{t},\hat{y}^{t},\tilde{y}^{t}$. First, notice that $\tilde{x}^{1},\tilde{y}^{1}$ and $\hat{x}^{1,k},\hat{y}^{1,k}$ are trivially private because they don’t depend on the dataset $S$. Now we consider $\tilde{x}^{t+1},\hat{x}^{t+1,k}$. Unrolling the $x$ update, we have
$x^{t+1}_{j}\propto\exp\left\{-\tau\left(\sum_{i=1}^{t}g^{i}_{x,j}\right)\right\},$
where $g^{t}_{x,j}$ denotes the $j$-th coordinate of $g^{t}_{x}$ for every $t\in[T]$. Hence, $\tilde{x}^{t+1}$ and $\hat{x}^{t+1,k}$ are vertices of the simplex $\Delta_{x}$, and each of them is choosen with the exponential mechanism with score function
$s^{t}_{x}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{x,j}\right).$
The privacy guarantee of the exponential mechanism in Theorem A.3 implies that $\tilde{x}^{t+1}$ and $\hat{x}^{t+1,k}$ are $2\Delta(s^{t}_{x})$-DP, where $\Delta(s^{t}_{x})$ is the sensitivity of $s^{t}_{x}$. Using the same idea, we argue that $\tilde{y}^{t+1}$ and $\hat{y}^{t+1,k}$ are all $2\Delta(s^{t}_{y})$-DP, where $\Delta(s^{t}_{y})$ is the sensitivity of $s^{t}_{y}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{y,j}\right)$.
By $L_{0}$-Lipschitzness of $f$ we have that $\max_{j\in[d]}\left|s^{t}_{x}(S,j)-s^{t}_{x}(S^{\prime},j)\right|\leq\frac{2L_{0}\tau}{B}$ when the past iterates $\hat{x}^{1},\tilde{x}^{1},\hat{y}^{1},\tilde{y}^{1},...,\hat{x}^{t},\tilde{x}^{t},\hat{y}^{t},\tilde{y}^{t}$ are fixed (advanced composition of DP allows to assume this). Hence, every sampled vertex $\tilde{x}^{t+1},\hat{x}^{t+1,k}$ is $\frac{4L_{0}\tau}{B}$-DP conditioned on the previous private vertices. With an analogous proof, $\tilde{y}^{t+1}$ and $\hat{y}^{t+1,k}$ are $\frac{4L_{0}\tau}{B}$-DP given $\hat{x}^{1},\tilde{x}^{1},\hat{y}^{1},\tilde{y}^{1},...,\hat{x}^{t},\tilde{x}^{t},\hat{y}^{t},\tilde{y}^{t}$. By the Advanced Composition (Theorem A.1), it suffices that each sampled vertex is $\frac{\varepsilon}{2\sqrt{2T(K+1)\log(1/\delta)}}$-DP in order to make the entire Algorithm 1 $(\varepsilon,\delta)$-DP. Then, for the algorithm to be $(\varepsilon,\delta)$-DP, it suffices that $\frac{4L_{0}\tau}{B}\leq\frac{\varepsilon}{2\sqrt{2T(K+1)\log(1/\delta)}}$. ∎
Next, we provide an explicit definition of the vectors $w^{t},v^{t}$ in Lemma 4.1:
$w^{1}=(1/d_{x},...,1/d_{x})\in\mathbb{R}^{d_{x}},w^{t+1}:=\operatorname{argmin}_{x\in\Delta_{x}}\left(\tau\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g_{x}^{t},x\rangle+\sum_{i\in[d_{x}]}w^{t}_{i}\log(w^{t}_{i}/x_{i})\right)$
$v^{1}=(1/d_{y},...,1/d_{y})\in\mathbb{R}^{d_{y}},v^{t+1}:=\operatorname{argmin}_{y\in\Delta_{y}}\left(\tau\langle-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g_{y}^{t},y\rangle+\sum_{i\in[d_{y}]}v^{t}_{i}\log(v^{t}_{i}/y_{i})\right).$
Note that $(w{t+1},v^{t+1})$ is measurable w.r.t. the algorithmic randomness up to iteration $t$ (i.e. it is a predictable process). We now prove the lemma.
Lemma 4.1.
Using the convex-concave properties of the loss:
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]$ $\leq\sum_{t\in[T]}[F_{\mathcal{D}}(x^{t},v)-F_{\mathcal{D}}(w,y^{t})]\leq\sum_{t\in[T]}\langle\Phi_{\mathcal{D}}(x^{t},y^{t}),(x^{t},y^{t})-(w,v)\rangle$
$\leq\operatorname{Regret}_{T}^{x}(w)+\operatorname{Regret}_{T}^{y}(v)+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle.$
We can bound the regret terms using the guarantees of SMD (see Lemma A.1):
$\operatorname{Regret}_{T}^{x}(w)\leq\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}_{x}\|_{\infty}^{2}\text{ and }\operatorname{Regret}_{T}^{y}(v)\leq\frac{\log(d_{y})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}_{y}\|_{\infty}^{2}$
To bound the other term, we use Lemma A.3. With our definition of $(w^{t})_{t\in[T]}$ we have
$\sum_{t\in[T]}\langle\tau(\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}),w^{t}-w\rangle\leq\log(d_{x})+\frac{1}{2}\sum_{t\in[T]}\tau^{2}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}.$
Using this result we can conclude that
$\sum_{t\in[T]}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w\rangle\\ \leq\sum_{t\in[T]}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w^{t}\rangle+\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}.$
Analogously, for the $y$ iterate
$\sum_{t\in[T]}\langle-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y},y^{t}-v\rangle\\ \leq\sum_{t\in[T]}\langle-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y},y^{t}-v^{t}\rangle+\frac{\log(d_{y})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}^{2}.$
Both bounds above imply that for any $w\in\Delta_{x},v\in\Delta_{y}$,
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]$ $\leq\frac{2(\log(d_{x})+\log(d_{y}))}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2})$
$\qquad+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
Since the right hand side does not depend on $w,v$, we can maximize the left hand sinde over $w\in\Delta_{x},v\in\Delta_{y}$ and this way we get the desired result. ∎
We now give a proof for Lemma 4.2, which allows us to bound the discrepancy on the duality gap evaluated at private and non-private iterates.
Lemma 4.2.
Define $F_{\lambda}(x):=\max_{y\in\Delta_{y}}[F_{\mathcal{D}}(x,y)+\lambda\phi(y)]$ where $\phi(y)=\sum_{j\in[d_{y}]}y_{j}\log(1/y_{j})$. Since $F_{\mathcal{D}}(x,y)$ is $L_{1}$-smooth in $(x,y)$ and concave in $y$, and $\lambda\phi(y)$ is $\lambda$-strongly convex in $y$, we have by Proposition 1 in Blondel et al. [2022] we have that $F_{\lambda}$ is $\big{(}L_{1}+\frac{L_{1}^{2}}{\lambda}\big{)}$-smooth wrt $\|\cdot\|_{1}$. It follows from Corollary 3.1 applied to $F_{\lambda}$ that
$\mathbb{E}[F_{\lambda}(\tilde{x})-F_{\lambda}(\bar{x})]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}.$
In addition, it is easy to see that $\max_{v\in\Delta_{y}}F_{\mathcal{D}}(x,v)\leq F_{\lambda}(x)\leq\lambda\log(d_{y})+\max_{v\in\Delta_{y}}F_{\mathcal{D}}(x,v)$ for all $x\in\Delta_{x}$, since $0\leq\phi(y)\leq\log(d_{y})$. This implies that
$\mathbb{E}\big{[}\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)\big{]}$
$=\mathbb{E}\big{[}\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-F_{\lambda}(\tilde{x})\big{]}+\mathbb{E}[F_{\lambda}(\tilde{x})-F_{\lambda}(\bar{x})]+\mathbb{E}\big{[}F_{\lambda}(\bar{x})-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)\big{]}$
$\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{y}).$
Analogously,
$\mathbb{E}\big{[}\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\bar{y})-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\tilde{y})\big{]}\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{x}).$
Using both inequalities, we get
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]$ $=\mathbb{E}[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\tilde{y})]$
$=\mathbb{E}[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)]+\mathbb{E}[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\bar{y})]$
$+\mathbb{E}[\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\bar{y})-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\tilde{y})]$
$\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{y})+\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]+\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{x})$
$=\frac{4L_{1}}{T}+\frac{4L_{1}^{2}}{\lambda T}+\lambda(\log(d_{x})+\log(d_{y}))+\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})].$
Choosing $\lambda=L_{1}\frac{2}{\sqrt{T(\log(d_{x})+\log(d_{y}))}}$ gives the bound. ∎
We now proceed by providing the proof of Theorem 4.2.
Theorem 4.2.
We follow the same proof as that of Theorem 4.1 until the inequality (1):
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{2(\log(d_{x})+\log(d_{y}))}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}.$
From Corollary D.2 we have $\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}\leq\frac{4L_{1}}{\sqrt{K}}$. In addition, from Lemma 4.2 it follows that $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq 8L_{1}\sqrt{\log(d_{x})+\log(d_{y})}/\sqrt{T}$. These bounds lead to
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\leq\frac{8L_{1}\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{T}}+\frac{2(\log(d_{x})+\log(d_{y}))}{\tau T}+5\tau L_{0}^{2}+\frac{2\sqrt{2}L_{1}}{\sqrt{K}}.$
Setting $T\eqsim\min\left\{n,\frac{(n\varepsilon)^{2/3}}{\log(1/\delta)^{1/3}}\right\}$, $\tau\eqsim\min\left\{\frac{1}{L_{0}}\sqrt{\frac{\log(d_{x})+\log(d_{y})}{T}},\frac{n\varepsilon}{L_{0}T\sqrt{TK\log(1/\delta)}}\right\}$ and $K\eqsim\frac{T}{\log(d_{x})+\log(d_{y})}$ gives
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\log(d_{x})+\log(d_{y})}{n}}+\left(\frac{\sqrt{\log(1/\delta)}(\log(d_{x})+\log(d_{y}))^{3/2}}{n\varepsilon}\right)^{1/3}\right].$
E.2 Missing proofs from Section 4.2
We prove Lemma 4.3, which provides guarantees on the bias-reduced gradient estimators.
Lemma 4.3.
We start with the bias bound. First,
$\|\mathbb{E}[(g_{x},g_{y})]-\nabla F_{\mathcal{D}}(x,y)\|_{\infty}$ $=\max\{\|\mathbb{E}[g_{x}]-\nabla_{x}F_{\mathcal{D}}(x,y)\|_{\infty},\|\mathbb{E}[g_{y}]-\nabla_{y}F_{\mathcal{D}}(x,y)\|_{\infty}\}.$
Note that, by the conditional independence of $(\bar{x}_{+,k},\bar{y}_{+,k})$, $(\bar{x}_{-,k},\bar{y}_{-,k})$ and $B_{k}$
$\quad\mathbb{E}[g_{x}]=\mathbb{E}\left[\sum_{k=0}^{M}(\nabla_{x}f(\bar{x}_{+,k},\bar{y}_{+,k};B_{k})-\nabla_{x}f(\bar{x}_{-,k},\bar{y}_{-,k};B_{k}))+\frac{2^{-k}}{C_{M}}\nabla_{x}f(x_{0},y_{0};B_{k})\right]$
$=\mathbb{E}\left[\sum_{k=0}^{M}\big{(}\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+,k},\bar{y}_{+,k})-\nabla_{x}F_{\mathcal{D}}(\bar{x}_{-,k},\bar{y}_{-,k})\big{)}+\nabla_{x}F_{\mathcal{D}}(x_{0},y_{0})\right]$
$=\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+,M},\bar{y}_{+,M})-\nabla_{x}F_{\mathcal{D}}(x_{-,0},y_{-,0})+\nabla_{x}F_{\mathcal{D}}(x_{0},y_{0})\right]$
$=\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+,M},\bar{y}_{+,M})\right],$
where in the last step we used that $(x_{-,0},y_{-,0})$ and $(x_{0},y_{0})$ are identically distributed. Finally, by Corollary 3.1 we get $\|\mathbb{E}[g_{x}]-\nabla_{x}F_{\mathcal{D}}(x,y)\|_{\infty}=\max_{j\in[d]}|\mathbb{E}[g_{x,j}]-\nabla_{x,j}F_{\mathcal{D}}(x,y)|\leq\frac{2L_{2}}{2^{M}}$. Clearly, an analogous bound holds for the partial gradient with respect to $y$.
We now proceed with the second moment bound.
$\mathbb{E}\|g_{x}\|_{\infty}^{2}$ $=\sum_{k=0}^{M}\mathbb{E}\Big{\|}C_{M}2^{k}[\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B_{k})-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B_{k})]+\nabla_{x}f(x_{0},y_{0};B_{k})\Big{\|}_{\infty}^{2}\cdot\frac{2^{-k}}{C_{M}}$
$\leq 2\sum_{k=0}^{M}C_{M}2^{k}\mathbb{E}\Big{\|}\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B_{k})-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B_{k})\Big{\|}_{\infty}^{2}+\frac{2^{-k}}{C_{M}}\mathbb{E}\|\nabla_{x}f(x_{0},y_{0};B_{k})\|_{\infty}^{2}$
$\leq 2L_{0}^{2}+4C_{M}\sum_{k=0}^{M}2^{k}\mathbb{E}\Big{\|}\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B_{k})-\nabla_{x}f(x,y;B_{k})\Big{\|}_{\infty}^{2}$
$\quad+4C_{M}\sum_{k=0}^{M}\mathbb{E}2^{k}\Big{\|}\nabla_{x}f(x,y;B_{k})-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B_{k})\Big{\|}_{\infty}^{2}.$
Now, note that by Corollary D.3
$\mathbb{E}\Big{\|}\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B_{k})-\nabla_{x}f(x,y;B_{k})\Big{\|}_{\infty}^{2}\leq\frac{20L_{2}^{2}}{2^{2(k+1)}}+\frac{(12+2\log(d_{x}))L_{1}^{2}}{2^{k+1}},$
and
$\mathbb{E}\Big{\|}\nabla_{x}f(x,y;B_{k})-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B_{k})\Big{\|}_{\infty}^{2}\leq\frac{20L_{2}^{2}}{2^{2k}}+\frac{(12+2\log(d_{x}))L_{1}^{2}}{2^{k}}.$
Hence, $\mathbb{E}\|g_{x}\|_{\infty}^{2}\lesssim L_{0}^{2}+L_{2}^{2}+M\log(d_{x})L_{1}^{2}$. Analogously, $\mathbb{E}\|g_{y}\|_{\infty}^{2}\lesssim L_{0}^{2}+L_{2}^{2}+M\log(d_{y})L_{1}^{2}$. ∎
Now we will prove Theorem 4.3. Define
$\mathcal{T}=\sup\left\{T\in\mathbb{N}:\sum_{t=1}^{T}2^{N_{t}}\leq U-2^{M}\right\}.$
Then, Algorithm 3 performs $\mathcal{T}+1$ optimization steps (see condition in Line 2). We will use later that, by definition of $\mathcal{T}$ we have $\sum_{t=1}^{\mathcal{T}}2^{N_{t}}\leq U-2^{M}$, and since $2^{N_{t}}\leq 2^{M}$ we have $\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq U$. We will first prove the privacy guarantees. Since this algorithm runs with random batch sizes and for a random number of steps, we analyze its privacy by the Fully Adaptive Composition Theorem [Whitehouse et al., 2023].
Proposition E.1.
Let $10<\delta<$ 1, $0<\varepsilon<8\log(1/\delta)$ and $U\leq\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}}$. Then, Algorithm 3 is $(\varepsilon,\delta)$-DP.
Proof.
By unraveling the recursion that defines the $x$ iterate, it is easy to see that its $j$-th coordinate has the form
$x^{t+1}_{j}\propto\exp\left\{-\tau\sum_{r=1}^{t}\Big{[}2^{N_{r}}\Big{(}\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big{)}+\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{r})\Big{]}\right\},$
hence sampling a vertex from $\Delta_{x}$ where the $j$-vertex is choosen with probability $x^{t}_{j}$ corresponds to an application of the exponential mechanism with score function given by
$s^{t}_{x}=-\tau\sum_{r=1}^{t}\Big{[}2^{N_{r}}\Big{(}\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big{)}+\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{r})\Big{]}.$
Next, note that in iteration $t$ of Algorithm 3, we sample a vertex $\tilde{x}^{t}\sim P_{x^{t}}$ in line 8 and we also sample $2^{N_{t}+1}$ times from $P_{x^{t}}$ to construct the bias-reduced gradient (with respect to $x$) estimate in line 1 of Algorithm 2. Analogously, sampling vertices from $\Delta_{y}$ corresponds to the exponential mechanism with score function
$s^{t}_{y}=-\tau\sum_{r=1}^{t}\Big{[}2^{N_{r}}\Big{(}\nabla_{y}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{y}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big{)}+\nabla_{y}f(x_{0}^{r},y_{0}^{r};B^{r})\Big{]}.$
Denote by $\Delta(s_{x}^{t})$ and $\Delta(s_{y}^{t})$ the sensitivities of the score functions $s_{x}^{t}$ and $s_{y}^{t}$, respectively. Recall that exponential mechanism with an score function that has sensitivity $\Delta$ is $2\Delta$-DP. Hence, based on the fully adaptive composition of DP (Theorem A.2), the algorithm is $(\varepsilon,\delta)$-DP if
$\sqrt{2\log(1/\delta)\sum_{t=1}^{\mathcal{T}+1}(2^{N_{t}+1}+1)(4\Delta(s_{x}^{t})^{2}+4\Delta(s_{y}^{t})^{2})}+\frac{1}{2}\sum_{t=1}^{\mathcal{T}+1}(2^{N_{t}+1}+1)(4\Delta(s_{x}^{t})^{2}+4\Delta(s_{y}^{t})^{2})\leq\varepsilon.$
Next, we will calculate the sensitivity of $s_{x}^{t}$ (it can be done analogously for $s_{y}^{t}$). Fix a pair of neighboring datasets $S$ and $S^{\prime}$. Denote by $s_{x}^{t}$ the score function with $S$ and by $s_{x}^{\prime t}$ with $S^{\prime}$. When conditioning on the randomness up to iteration $t$ (recall the batch-size is a random quantity), either $B^{r}=B^{\prime r}$ for all $r\in[t]$, in which case $s_{x}^{t}=s_{x}^{\prime t}$, or there is exactly one index $r$ such that $B^{r}$ differ from $B^{\prime r}$ in exactly one datapoint (say $z^{*}$ and $z^{\prime*}$). In this case, $|s_{x}^{t}-s_{x}^{\prime t}|$ can be calculated as
$|s_{x}^{t}-s_{x}^{\prime t}|$
$=\tau\Big{|}2^{N_{r}}\Big{(}\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big{)}+\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{r})$
$\qquad-2^{N_{r}}\Big{(}\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{\prime r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{\prime r})\Big{)}-\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{\prime r})\Big{|}$
$\leq\tau\Big{|}\frac{2^{N_{r}}\alpha}{2^{N_{r}}}\Big{[}\Big{(}\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};z^{*})-\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};z^{\prime*})\Big{)}+\Big{(}\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};z^{*})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};z^{\prime*})\Big{)}\Big{]}$
$\qquad+\frac{\alpha}{2^{N_{r}}}\Big{(}\nabla_{x}f(x_{0}^{r},y_{0}^{r};z^{*})-\nabla_{x}f(x_{0}^{r},y_{0}^{r};z^{\prime*})\Big{)}\Big{|}$
$\leq\tau\alpha L_{0}(4+1/2^{N_{r}})\leq 4.5\tau\alpha L_{0}.$
Analogously, $\Delta(s_{y}^{t})\leq 4.5\tau\alpha L_{0}$. Hence, each time we use the exponential mechanism to sample a vertex (regardless of whether we are sampling vertices from $\Delta_{x}$ or of $\Delta_{y}$), the mechanism is $9\tau\alpha L_{0}$-DP. Plugging in the condition for privacy mentioned above, we now need
$\sqrt{2\log(1/\delta)\sum_{t=1}^{\mathcal{T}+1}(4\cdot 2^{N_{t}}+2)(9\tau\alpha L_{0})^{2}}+\frac{1}{2}\sum_{t=1}^{\mathcal{T}+1}(4\cdot 2^{N_{t}}+2)(9\tau\alpha L_{0})^{2}\leq\varepsilon.$
We further restrict this stopping time by upper bounding each of the above terms by $\varepsilon/2$. Further noticing that such a bound for the first term implies the same bound for the second term so long as $\varepsilon\leq 8\log(1/\delta)$, we conclude that it suffices
$4\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}+2(\mathcal{T}+1)\leq 6\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq\frac{\varepsilon^{2}}{8\log(1/\delta)(9\tau\alpha L_{0})^{2}}.$
Noting that
$4\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}+2(\mathcal{T}+1)$ $\leq 6\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq 6\sum_{t=1}^{\mathcal{T}}2^{N_{t}}+6\cdot 2^{M}\leq 6(U-2^{M})+6\cdot 2^{M}=6U,$
the condition is satisfied when $U\leq\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}}$. ∎
Now we start the accuracy analysis of Algorithm 3. First, we note that the resulting guarantee on the gap that we prove in Theorem 4.3 holds with constant probability is of the order
$\log(n)\left(L_{0}+L_{2}+L_{1}\sqrt{(\log(d_{x})+\log(d_{y}))}\right)\sqrt{\frac{\log(d_{x})+\log(d_{y})}{n}}\\ +\log(n)\sqrt{L_{0}\sqrt{L_{0}^{2}+L_{2}^{2}+L_{1}^{2}\log(n)(\log(d_{x})+\log(d_{y}))}}\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}.$
Up to poly-logarithmic factors in $n,d_{x},d_{y}$, such rate is indeed optimal, as it will be mentioned Appendix G. Now, we proceed to prove the theorem.
Theorem 4.3.
First, recall from the privacy analysis that we need $U\leq\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}}$. At the same time, we need $\sum_{t\in[\mathcal{T}+1]}\max\{1,2^{N_{t}}/\alpha\}\leq n$, to ensure Algorithm 3 does not exceed the sample size. For this condition to hold, it suffices that $\mathcal{T}+1\leq n/2$ and $\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq n\alpha/2$. This way, we only need $\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq\min\{n,n\alpha/2\}$. Since $\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq U$, then for this last condition to hold, it is sufficient that $U\leq\min\{n,n\alpha/2\}$. Hence, defining $U\leq\min\left\{\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}},n\alpha/2,n/2\right\}$ ensures that we don’t exceed our privacy budget and that we respect the given sample size constraint. We choose $\alpha=\left(\frac{2\varepsilon^{2}}{48\cdot 81\log(1/\delta)(\tau L_{0})^{2}n}\right)^{1/3}$ so $U=\min\left\{\frac{(n\varepsilon)^{2/3}}{(4\cdot 48\cdot 81\log(1/\delta)^{1/3}(\tau L_{0})^{2/3}},\frac{n}{2}\right\}$.
Now, let $(\tilde{x},\tilde{y})$ be the output of the algorithm. We will prove that $\operatorname{Gap}(\tilde{x},\tilde{y})\leq A/B$ with probability at least $0.99$, for some $A,B$ to be defined later. In other words, we will prove that $\mathbb{P}[\operatorname{Gap}(\tilde{x},\tilde{y})\geq A/B]\leq 0.01$. Notice that
$\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\geq\frac{A}{B}\right]$ $=\mathbb{P}\left[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\cdot\frac{1}{\mathcal{T}+1}\geq\frac{A}{B}\right]$
$\leq\mathbb{P}\left[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\geq A\right]+\mathbb{P}\left[\frac{1}{(\mathcal{T}+1)}\geq\frac{1}{B}\right]$
Note the second probability can be bounded as follows:
$\mathbb{P}\left[\frac{1}{\mathcal{T}+1}\geq\frac{1}{B}\right]=\mathbb{P}\left[\mathcal{T}\leq B-1\right]\leq\mathbb{P}\left[\sum_{t\in[B]}2^{N_{t}}>U-2^{M}\right]\leq\frac{B\mathbb{E}[2^{N_{1}}]}{U-2^{M}},$
where the first inequality follows from the fact that $\mathcal{T}\leq B-1$ implies that $\sum_{t\in[B]}2^{N_{i}}>U$, and the second follows from Markov’s inequality. By using Markov’s inequality again we can see that
$\mathbb{P}\left[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\geq A\right]\leq\frac{\mathbb{E}\left[(\mathcal{T}+1)(\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y}))\right]+\mathbb{E}\left[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})\right]}{A}.$
Let us bound $\mathbb{E}\left[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})\right]$. From Lemma 4.1, we have
$(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})$ $\lesssim\frac{\log(d_{x})+\log(d_{y})}{\tau}+\frac{\tau}{2}\sum_{t=1}^{\mathcal{T}+1}(L_{0}^{2}+\|g^{t}_{x}\|_{\infty}^{2}+\|g^{t}_{y}\|_{\infty}^{2})$
$\qquad+\sum_{t=1}^{\mathcal{T}+1}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-(g_{x}^{t},g_{y}^{t}),(x^{t},y^{t})-(w^{t},v^{t})\rangle$
Since $\mathcal{T}+1\leq\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq U$, then
$\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\|g^{t}_{x}\|_{\infty}^{2}\right]$ $=\mathbb{E}\left[\sum_{t=1}^{U}\|g^{t}_{x}\|_{\infty}^{2}\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]$
$=\mathbb{E}\left[\sum_{t=1}^{U}\mathbb{E}[\|g^{t}_{x}\|_{\infty}^{2}\mathbbm{1}_{(\mathcal{T}\geq t-1)}\mid\mathcal{F}_{t}]\right]$
$=\mathbb{E}\left[\sum_{t=1}^{U}\mathbb{E}[\|g^{t}_{x}\|_{\infty}^{2}]\mathbbm{1}_{(\mathcal{T}\geq t-1)}\right]$
$\lesssim\left(L_{0}^{2}+L_{2}^{2}+\log(d_{x})ML_{1}^{2}\right)\mathbb{E}\left[\sum_{t=1}^{U}\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]$
$\lesssim\left(L_{0}^{2}+L_{2}^{2}+\log(d_{x})ML_{1}^{2}\right)\mathbb{E}[\mathcal{T}+1],$
where $\mathcal{F}_{t}$ is the sigma algebra generated by all the randomness in the algorithm that generated $x^{t}$. In the second and third equalities, we used that $\{\mathcal{T}\geq t-1\}=\Big{\{}\sum_{k\in[t-1]}2^{N_{k}}\leq U\Big{\}}$ is $\mathcal{F}_{t}$-measurable (since $2^{N_{1}},...,2^{N_{t-1}}$ are used to generate $x^{t}$), and hence it can be taken out of the conditional expectation. Next, for the first inequality we used that $2^{N_{t}}$ is independent of $\mathcal{F}_{t}$, so we can use Lemma 4.3 to bound the conditional expectation of the second moment. In the last line we used the fact that the expectation can be written as the sum of the tail probabilities. Analogously, we can obtain $\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\|g^{t}_{y}\|_{\infty}^{2}\right]\lesssim L_{0}^{2}+L_{2}^{2}+\log(d_{y})ML_{1}^{2}\mathbb{E}[\mathcal{T}+1]$. In a similar way, we can use the same technique of introducing the indicator of $\mathcal{T}+1\geq t$ and then Lemma 4.3 to bound the bias term
$\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-(g_{x}^{t},g_{y}^{t}),(x^{t},y^{t})-(w^{t},v^{t})\rangle\right]\lesssim\frac{L_{2}\mathbb{E}[\mathcal{T}+1]}{2^{M}}.$
Now, we take expectations on both sides of the expression given by Lemma 4.1 and then use the bounds we calculated above for the second moment and bias of the bias-reduced gradient estimators:
$\mathbb{E}[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})]$
$\lesssim\frac{\log(d_{x})+\log(d_{y})}{\tau}+\tau[L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))ML_{1}^{2}]\mathbb{E}[\mathcal{T}+1]+\frac{L_{2}\mathbb{E}[\mathcal{T}+1]}{2^{M}}.$
To bound $\mathbb{E}[\mathcal{T}+1]$ we note that $\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq U$, so using Wald’s identity [Wald, 1944] yields
$U$ $\geq\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\right]=\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\mathbb{E}[2^{N_{t}}\mathbbm{1}_{(\mathcal{T}+1\geq t)}|N_{t-1},...,N_{1}]\right]$
$=\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\mathbb{E}[2^{N_{t}}]\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]$
$=\mathbb{E}[2^{N_{1}}]\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]$
$=\mathbb{E}[2^{N_{1}}]\mathbb{E}\left[\mathcal{T}+1\right]\eqsim M\mathbb{E}\left[\mathcal{T}+1\right].$
Hence, $\mathbb{E}\left[\mathcal{T}+1\right]\lesssim U/\mathbb{E}[2^{N_{1}}]\eqsim U/M$. It follows that
$\mathbb{E}[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})]$ $\lesssim\frac{\log(d_{x})+\log(d_{y})}{\tau}+\tau[L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))ML_{1}^{2}]\frac{U}{M}+\frac{L_{2}U}{M2^{M}}$
$\lesssim\frac{\log(d_{x})+\log(d_{y})}{\tau}+\tau[L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]U+\frac{L_{2}U}{2^{M}}.$
Choosing $M=\log_{2}(\sqrt{U})$ and $\tau=\sqrt{\frac{(\log(d_{x})+\log(d_{y}))}{(L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2})U}}$
$\mathbb{E}[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})]\lesssim\sqrt{[(\log(d_{x})+\log(d_{y})][L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]U}+L_{2}\sqrt{U}.$
Now, let’s bound $\mathbb{E}\left[(\mathcal{T}+1)(\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y}))\right]$. We can assume that $\tilde{x}$’s and $\tilde{y}$’s are sampled after the While Loop is over in the Algorithm, thus by Lemma 4.2
$\mathbb{E}\left[(\mathcal{T}+1)(\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y}))\right]$ $=\mathbb{E}\left[(\mathcal{T}+1)\mathbb{E}_{\tilde{x},\tilde{y}}\left[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})\mid\mathcal{T}\right]\right]$
$\leq\mathbb{E}\left[(\mathcal{T}+1)\left(\frac{4L_{1}}{\mathcal{T}+1}+\frac{2L_{1}\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{\mathcal{T}+1}}\right)\right]$
$=4L_{1}+2L_{1}\sqrt{\log(d_{x})+\log(d_{y})}\mathbb{E}\left[\sqrt{\mathcal{T}+1}\right]$
$\leq 4L_{1}+2L_{1}\sqrt{[\log(d_{x})+\log(d_{y})]\mathbb{E}[(\mathcal{T}+1)]}$
$\lesssim\sqrt{[(\log(d_{x})+\log(d_{y})]L_{1}^{2}\frac{U}{M}}$
where the second step follows by Jensen’s inequality. In summary, we have
$\mathbb{P}\left[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\geq A\right]\lesssim\frac{\sqrt{[(\log(d_{x})+\log(d_{y})][L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]U}}{A}.$
Notice that
$\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\geq\frac{A}{B}\right]$ $\lesssim\frac{\sqrt{[(\log(d_{x})+\log(d_{y})][L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]U}}{A}$
$+\frac{BM}{U-\sqrt{U}}.$
Assuming $U\geq 4$ we have that $U-\sqrt{U}\geq U/2$. Hence, choosing
$A\eqsim\sqrt{[(\log(d_{x})+\log(d_{y})][L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]U}\quad,\quad B\eqsim U/M$
we get
$\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\lesssim\sqrt{\frac{[(\log(d_{x})+\log(d_{y})][L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}]\log(U)^{2}}{U}}\right]\geq 0.99.$
Finally, note that with our choice of $\tau$, we obtain $U=\min\left\{\frac{n\varepsilon\sqrt{L_{0}^{2}+L_{2}^{2}+(\log(d_{x})+\log(d_{y}))L_{1}^{2}}}{\sqrt{4\cdot 48\cdot 81(\log(d_{x})+\log(d_{y}))\log(1/\delta)}L_{0}},\frac{n}{2}\right\}$. ∎
Appendix F Missing details from Section 5
We give the explicit parameter setting to get the convergence rate for Lipschitz and smooth functions claimed in Theorem 5.1:
$T=\min\left\{n,\frac{(n\varepsilon)^{4/5}}{(\log(d_{x})\log(1/\delta))^{2/5}}\right\},q=\sqrt{T}/\log(d_{x}),K=T/\log(d_{x})$
$\tau\eqsim\min\left\{\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+(L_{0}^{2}+L_{1}^{2})q\sqrt{\log(d_{x})/K}+L_{1}^{2}q/[\sqrt{\log(d_{x})}K^{3/2}])T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\}.$
We also give the explicit parameter setting to get the convergence rate for Lipschitz and smooth functions with smooth partial derivatives claimed in Theorem 5.1
$T=\min\Big{\{}n,\frac{n\varepsilon}{\log(d_{x})\sqrt{\log(1/\delta)}}\Big{\}},K=\sqrt{T\log(d_{x})},q=\sqrt{T/\log(d_{x})},$
$\tau\eqsim\min\left\{\left(\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})T}}\right),\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\}.$
In the following result, we make explicit the advantage of using the anytime online-to-batch approach in terms of privacy. Namely, due to the substantial sampling reduction, our privacy budget degrades much more nicely.
Proposition F.1.
Algorithm 4 is $(\varepsilon,\delta)$-DP when $\tau\leq\frac{B\varepsilon}{8L_{0}\sqrt{2(TK/q+qK)\log(1/\delta)}}$.
Proof.
Note that $w^{t}=\frac{1}{t}\sum_{i\in[t]}x^{i}$. Hence, for any $j\in[d_{x}]$
$w^{t}_{j}=\frac{1}{t}\sum_{i\in[t]}x^{i}_{j}\propto\sum_{i\in[t]}\exp\left(-\tau\sum_{k\in[i-1]}g^{k}_{j}\right)=\exp\left(\log\left(\sum_{i\in[t]}\exp\left(-\tau\sum_{k\in[i]}g^{k}_{j}\right)\right)\right).$
Denote by $G_{j}(S)=\left(\sum_{k\in[i-1]}g^{k}_{j}\right)_{i\in[t]}$ the vector of the cumulative sum of the gradients that our algorithm produces when it has been fed with dataset $S$. Also, define the log-sum-exp function as $\operatorname{LSE}((x_{1},...,x_{k}))=\log(\exp(x_{1})+...+\exp(x_{k}))$. With this new notation
$w^{t}_{j}\propto\exp(\operatorname{LSE}(\tau G_{j}(S))).$
Two well know properties of the $\operatorname{LSE}(x)$ function are that it is convex on its argument and its gradient is the softmax function of $x$. As a consequence of the second property mentioned, it is clear that $\|\nabla\operatorname{LSE}(x)\|_{1}=1$ for every $x$. By convexity of $\operatorname{LSE}$, we have
$\operatorname{LSE}(-\tau G_{j}(S))-\operatorname{LSE}(-\tau G_{j}(S^{\prime}))$ $\leq\langle\nabla\operatorname{LSE}(-\tau G_{j}(S)),\tau G_{j}(S^{\prime})-\tau G_{j}(S)\rangle$
$\leq\tau\|\nabla\operatorname{LSE}(-\tau G(S))\|_{1}\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}$
$=\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}.$
Next, note that by $L_{0}$-Lipschitzness of $f$
$\max_{j\in[d_{x}]}\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}=\tau\max_{j\in[d_{x}],i\in[t]}\left|\sum_{k\in[i-1]}(g^{k}_{j}-g^{\prime k}_{j})\right|\leq\frac{2L_{0}}{B}.$
Hence, sampling vertices from $\Delta_{x}$ according to the probabilities induced by $w^{t}$ is equivalent to sampling with the exponential mechanism with score function $s(S,j)=\operatorname{LSE}(\tau G_{j}(S))$, whose sensitivity is $\frac{2\tau L_{0}}{B}$. Since we sample $K$ times during the first $q$ iterations, then every $q$ iterations and then $K$ times to sparsify the last iterate, the overall number of samplings will be $TK/q+qK+K$. Since each of them is $\frac{4\tau L_{0}}{B}$-DP, then by Advanced Composition of DP (Theorem A.1) we only need that
$\frac{4\tau L_{0}}{B}\leq\frac{\varepsilon}{2\sqrt{2(TK/q+(q+1)K)\log(1/\delta)}}.$
Proof of Theorem 5.1.
Theorem 5.1.
Let $x\in\Delta_{x}$ be a minimizer of $F_{\mathcal{D}}(\cdot)$. From Lemma 5.1 we have
$F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)\leq\frac{1}{T}\left[\operatorname{Regret}_{T}(x)+\sum_{t\in[T]}\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right],$
where $\operatorname{Regret}_{T}(x)\leq\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2}$ by Lemma A.1. Using that $\|g^{t}\|_{\infty}\leq L_{0}$ and taking expectations, we obtain
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]\leq\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right].$
Let us bound the term $\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$. For $t\leq q$, we have $g^{t}=\nabla f(\hat{w}^{t};B^{t})$, so
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]$ $=\mathbb{E}[\langle\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]$
$=\mathbb{E}_{x^{t}}[\langle\mathbb{E}_{\hat{w}^{t}}[\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t})\mid x^{t}],x^{t}-x\rangle]$
$\leq 2\|\mathbb{E}_{\hat{w}^{t}}[\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t})]\|_{\infty}$
$\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
where in the first equality we used that $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ and for the second inequality we used Corollaries D.2 and D.1. Now we consider the case where $t>q$. First, denote by $t_{0}=q\lfloor t/q\rfloor$. That is, $t_{0}$ was the last time that $\hat{w}$ was updated by the algorithm before the $t$-th iteration. Hence, $g^{t}=\nabla f(\hat{w}^{t_{0}};B^{t})$. Now, we decompose the error into three terms and bound each of them individually:
$\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle$ $=\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle$
$+\langle\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t}),x^{t}-x\rangle$
$+\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle.$
Clearly $\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle]=0$, since $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ conditioned on the the randomness up to iteration $t$. For the second term, note that
$\langle\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t}),x^{t}-x\rangle$ $\leq\|\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t})\|_{\infty}\|x^{t}-x\|_{1}$
$\leq 2L_{1}\|w^{t}-w^{t_{0}}\|_{1}\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\|w^{k+1}-w^{k}\|_{1}$
$\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\frac{2}{k+1}\leq\frac{4L_{1}q}{q\lfloor t/q\rfloor+1},$
where the first inequality follows from Hölder, the second from the $L_{1}$-smoothness of $f$, the third from the triangle inequality for the $\ell_{1}$ norm and the fourth from the stability of the average iterates. Now, let us bound the third term: $\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle$. We will further decompose this term into
$\langle\tilde{\Delta}^{t_{0}},x^{t}-x\rangle=\underbrace{\langle\tilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle}_{=:A_{1}}+\underbrace{\langle\tilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle}_{=:A_{2}},$
where $\tilde{\Delta}^{t_{0}}$ denotes $\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})$ and $\tilde{x}^{t}$ is a ficticious iterate obtained by starting at $x^{t_{0}}$ and performing updates $\tilde{x}_{j}^{t+1}\propto\tilde{x}_{j}^{t}\exp\big{(}-\tau\nabla_{j}f(w^{t_{0}};B^{k})\big{)}$. We start by bounding $A_{1}$:
$\mathbb{E}_{\hat{w}^{t_{0}}}[\langle\tilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle]$ $=\langle\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})],\tilde{x}^{t}-x\rangle$
$\leq 2\|\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})]\|_{\infty}$
$\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
where the equality holds since $\tilde{x}^{t}$, as opposed to $x^{t}$, does not depend on $\hat{w}^{t_{0}}$, and the last inequality follows from Corollaries D.2 and D.1. By the law of iterated expectation,
$\mathbb{E}[A_{1}]=\mathbb{E}[\langle\tilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle]\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
We bound now $A_{2}$. Note that $\langle\tilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle\leq\|\tilde{\Delta}^{t_{0}}\|_{\infty}\|x^{t}-\tilde{x}^{t}\|_{1}$. Note that for $j\in[d_{x}]$
$(x^{t}-\tilde{x}^{t})_{j}=\frac{x^{t_{0}}_{j}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(\hat{w}^{t_{0}};B^{k})\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{i}f(\hat{w}^{t_{0}};B^{k})\right)}-\frac{x^{t_{0}}_{j}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(w^{t_{0}};B^{k})\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(-\tau\sum_{k=t_{0}}^{t-1}\nabla_{i}f(w^{t_{0}};B^{k})\right)}.$
Denote $\hat{G}_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(\hat{w}^{t_{0}};B^{k}),G_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(w^{t_{0}};B^{k})$. We rewrite the equality above in terms of $\hat{G}_{j}$ and $G_{j}$ and bound $(x^{t}-\tilde{x}^{t})_{j}$ as follows
$(x^{t}-\tilde{x}^{t})_{j}$ $=\frac{x^{t_{0}}_{j}\exp\left(\hat{G}_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(\hat{G}_{i}\right)}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$=\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)\exp\left(\hat{G}_{j}-G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)\exp\left(\hat{G}_{i}-G_{i}\right)}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[\exp\left(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right)-1\right],$
and similarly
$(\tilde{x}^{t}-x^{t})_{j}$ $=\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)\exp\left(\hat{G}_{j}-G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)\exp\left(\hat{G}_{i}-G_{i}\right)}$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[1-\exp\left(-2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right)\right]$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right].$
Combining both inequalities, we obtain
$|(x^{t}-\tilde{x}^{t})_{j}|\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\underbrace{\max\left\{2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|,\exp\left(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right)-1\right\}}_{:=\psi},$
and hence
$\|x^{t}-\tilde{x}^{t}\|_{1}=\sum_{j\in[d_{x}]}|(x^{t}-\tilde{x}^{t})_{j}|\leq\sum_{j\in[d_{x}]}\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\psi=\psi.$
Next, we use that $e^{x}-1\leq 2x$ for $x\leq 1$. Assume for now $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$, so $\psi=4\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|$. Using the definition of $\hat{G}_{j},G_{j}$, it follows that
$\|x^{t}-\tilde{x}^{t}\|_{1}\leq\psi$ $=4\max_{j\in[d_{x}]}\tau\left|\sum_{k=t_{0}}^{t-1}[\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})]\right|$
$\leq 4\tau\max_{j\in[d_{x}]}\sum_{k=t_{0}}^{t-1}\left|\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})\right|$
$\leq 4\tau\sum_{k=t_{0}}^{t-1}\left\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\right\|_{\infty}.$
This has two implications. The first is that $\psi\leq 8\tau qL_{0}$. This further implies that in order to ensure that our assumption $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$ holds, it suffices that $8\tau qL_{0}\leq 2$, or equivalently $\tau\leq 1/(4L_{0}q)$. We impose this condition on $\tau$ later when optimizing the convergence rate. The second is the following bound on $A_{2}$
$A_{2}$ $=\langle\tilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle$
$=\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-\tilde{x}^{t}\rangle$
$\leq\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}\|x^{t}-\tilde{x}^{t}\|_{1}$
$\leq 4\tau\sum_{k=t_{0}}^{t-1}\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}$
$\leq 2\tau\sum_{k=t_{0}}^{t-1}\Big{(}\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\Big{)}$
$\leq 2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\Big{(}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\Big{)}$
$\lesssim\begin{cases}q\tau\big{[}\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big{]}&\text{without second order smoothnes}\\ q\tau\big{[}\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big{]}&\text{with second order smoothness}\end{cases},$
where we bounded the second moment of the estimation errors with the results in Corollaries D.3 and D.4. We conclude that for all $t>q$
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle]$
$=\mathbb{E}\big{[}\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle\big{]}+\mathbb{E}\big{[}\langle\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t}),x^{t}-x\rangle\big{]}$
$\quad+\mathbb{E}\big{[}\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle\big{]}$
$\leq 0+\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+2\|\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})]\|_{\infty}$
$\quad+2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\mathbb{E}\left[\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\right]$
$\lesssim\begin{cases}\frac{L_{1}}{\sqrt{K}}+q\tau\big{[}\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big{]}&\text{without second order smoothnes}\\ \frac{L_{2}}{K}+q\tau\big{[}\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big{]}&\text{with second order smoothness}\end{cases}$
And recall that for $t\leq q$
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{without second order smoothnes}\\ \frac{4L_{2}}{K}&\text{with second order smoothness}\end{cases}.$
Finally, we plug this bounds into our excess risk bound. We start considering the case of second order smooth functions:
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$
$\leq\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\left(\frac{L_{2}}{K}+\mathbbm{1}_{(t>q)}\left[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+q\frac{L_{2}^{2}\tau}{K^{2}}+q\frac{L_{1}^{2}\log(d_{x})\tau}{K}\right]\right)$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big{]}+\frac{L_{2}}{K}+\frac{1}{T}\sum_{t=q+1}^{T}\left[\frac{L_{1}q}{q\lfloor t/q\rfloor+1}\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big{]}+\frac{L_{2}}{K}+\frac{L_{1}q\log(T/q)}{T},$ (5)
where the last line follows from $\sum_{t=q+1}^{T}\left[\frac{1}{q\lfloor t/q\rfloor+1}\right]\leq q\left(\frac{1}{q}+\frac{1}{2q}+...+\frac{1}{(T/q)q}\right)\lesssim\log\Big{(}\frac{T}{q}\Big{)}$. Now, by Theorem F.1, and recalling that the batch size is $B=n/T$, we have a constraint on $\tau>0$ in order to satisfy privacy. We also have the constraint of $\psi\leq 1$, which holds true only when $\tau\leq 1/(8qL_{0})$. Optimizing (5) under these constraints, we have that we should choose
$\tau\eqsim\min\left\{\left(\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})T}}\right),\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\},$
we obtain
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$ $\lesssim\sqrt{\frac{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})\log(d_{x})}{T}}+\frac{L_{0}q\log(d_{x})}{T}$
$\qquad+\frac{L_{0}\sqrt{(TK/q+qK)\log(1/\delta)}\log(d_{x})}{n\varepsilon}+\frac{L_{2}}{K}+\frac{L_{1}q\log(T/q)}{T}.$
Note that the sparsification error is, by Corollary 3.1, $\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(w^{T})]\lesssim\frac{L_{1}}{K}$. Combining this with the previous inequality and setting $q=\sqrt{T/\log(d_{x})},K=T/q=\sqrt{T\log(d_{x})}$, leads to
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]$ $\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}+\frac{L_{2}}{T^{3/4}\log(d_{x})^{1/4}}+\frac{L_{1}+L_{2}}{\sqrt{T\log(d_{x})}}$
$+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}.$
$\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}+\frac{L_{2}}{\sqrt{T}\log(d_{x})^{1/4}}$
$+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}.$
Hence letting $T=\min\left\{n,\frac{n\varepsilon}{\log(d_{x})\sqrt{\log(1/\delta)}}\right\}$ gives the claimed rate for second-order-smooth functions. Now, we look at the case where the partial derivates are not necessarily smooth. Now, we have
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$
$\leq\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\left(\frac{L_{1}}{\sqrt{K}}+\mathbbm{1}_{(t>q)}\left[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+\frac{L_{1}^{2}q\tau}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q\tau}{\sqrt{K}}\right]\right)$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big{]}+\frac{L_{1}}{\sqrt{K}}+\frac{1}{T}\sum_{t=q+1}^{T}\left[\frac{L_{1}q}{q\lfloor t/q\rfloor+1}\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big{]}+\frac{L_{1}}{\sqrt{K}}+\frac{L_{1}q\log(T/q)}{T}.$
By Corollary 3.1 the sparsification error is $\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(w^{T})]\lesssim\frac{L_{1}}{K}$, which implies
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big{[}L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big{]}+\frac{L_{1}}{\sqrt{K}}+\frac{L_{1}q\log(T/q)}{T}.$
Setting $T=\min\left\{n,\frac{(n\varepsilon)^{4/5}}{(\log(d_{x})\log(1/\delta))^{2/5}}\right\}$, $q=\sqrt{T}/\log(d_{x})$, $K=T/\log(d_{x})$ and
$\tau\eqsim\min\left\{\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+(L_{0}^{2}+L_{1}^{2})q\sqrt{\log(d_{x})/K}+L_{1}^{2}q/[\sqrt{\log(d_{x})}K^{3/2}])T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\}$ leads to the convergence rate claimed for first-order-smooth functions. ∎
Appendix G Algorithm for Statistical Synthetic Data Generation
We introduced the Statistical Synthetic Data Generation problem in Appendix B. Recall its SSP formulation (3)
$\min_{x\in\Delta_{|\mathcal{Z}|}}\max_{y\in\Delta_{|\mathcal{Q}|}}\mathbb{E}_{z\sim\mathcal{D}}\Big{[}\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)\Big{]}.$
We denote the objective function by $f(x,y;z)=\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)$. This objective is convex-concave, since it is linear on $x$ and on $y$. It is easy to check that the it is also Lipschitz and smooth w.r.t $\|\cdot\|_{1}$. Furthemore, it is quadratic, so we could use Algorithm 1 directly. However, the algorithm below leverages the structure of the objective that we are optimizing in this specific formulation. The algorithm is the following
Algorithm 5 DP synthetic data generation with DP stochastic mirror descent and Maurey sparsification
0:  Dataset $S$, Number of steps $T$, step-sizes $\tau_{x},\tau_{y}$.
1:  $x^{1}=(1/|\mathcal{Z}|,...,1/|\mathcal{Z}|),y^{1}=(1/|\mathcal{Q}|,...,1/|\mathcal{Q}|)$
2:  for $t=1$ to $T$ do
3:     $\hat{y}^{t}\sim P_{y^{t}}$
4:     $g^{t}_{x}=\nabla_{x}f(x^{t},\hat{y}^{t};S)$
5:     $g^{t}_{y}=\nabla_{y}f(x^{t},\hat{y}^{t};S)$
6:     $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau_{x}g^{t}_{x,j}\right),\quad\forall j\in[|\mathcal{Z}|]$
7:     $y^{t+1}_{j}\propto y^{t}_{j}\exp\left(\tau_{y}g^{t}_{y,j}\right),\quad\forall j\in[|\mathcal{Q}|]$
8:  end for
9:  $\bar{x}=\frac{1}{T}\sum_{t=1}^{T}x^{t}$
[…]
11:  return  $\tilde{x}$
We obtain the following theorem.
Theorem G.1.
Let $10<\varepsilon,\delta<$ 1. If $|\mathcal{Q}|\leq|\mathcal{Z}|^{C}$, Algorithm 5 with $T=\frac{6n\varepsilon}{16\sqrt{2\log(1/\delta)}\log(|{\cal Q}|)},\tau_{x}=\sqrt{\frac{\log(|\mathcal{Z}|)}{9T}},\tau_{y}=\frac{\log(|{\cal Q}|)}{6\sqrt{\log(|\mathcal{Z}|)T}}$ constructs an $(\varepsilon,\delta)$-DP synthetic dataset $\tilde{S}$ of $n$ samples, such that
$\mathbb{E}_{\mathcal{A},S}\Big{[}\max_{q\in\mathcal{Q}}|\mathbb{E}_{z\sim\mathcal{D}}[q(z)]-q(\tilde{S})|\Big{]}\leq 20\sqrt{\frac{\log(|{\cal Q}|)}{n}}+2(12+C/3)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$
Furthermore, this rate is optimal.
The proof of Theorem G.1 follows the same ideas we used for analyzing Algorithm 1.
Proof.
We prove the optimality, privacy and convergence rate claimed in the theorem seperately.
Optimality: The optimality of the rate is a folk result. The private lower bound follows from [Bun et al., 2014] combined with a resampling argument from [Bassily et al., 2019]. The non-private term comes from an application of Fano’s inequality [Cover and Thomas, 2006].
Privacy Analysis: Since $x^{1},y^{1}$ are trivially private, then $\tilde{y}^{1}$ is also private, as it is sampled according to the (private) probability vector $y^{1}$. Next, we have that $x^{2}$ is immediately private, as it only depends on $x_{1}$ and $\hat{y}_{1}$, and not on the data. However, $g_{y}^{1}$ depends on the data, so in order to privatize $y^{2}$ we create a private version of it using the the exponential mech as we did in Algorithm 1. Since we are using a full batch to estimate $g^{t}_{y}$, the sensitivity of the score function $s^{t}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{y,j}\right)$ is $\frac{2TL_{0}\tau}{n}$. Now, we have that $x^{2}$ and $\hat{y}^{2}$ are both $\frac{4TL_{0}\tau}{n}$-DP. We inductively continue this argument, and using the advanced composition (Theorem A.1) to conclude that the average $\bar{x}$ is $(\varepsilon,\delta)$-DP so long as
$\frac{4TL_{0}\tau_{y}}{n}\leq\frac{\varepsilon}{2\sqrt{2T\log(1/\delta)}}.$
This condition holds under the parameter setting of Theorem G.1. The privacy of the output $\tilde{x}$ follows from post-processing of DP.
Accuracy Analysis: We start by noticing that
$\mathbb{E}\left[\max_{q\in{\cal Q}}|\mathbb{E}_{z}[q(z)]-q(\tilde{S})|\right]$ $=\mathbb{E}\left[\max_{y\in\Delta_{|{\cal Q}|}}\sum_{j\in[|{\cal Q}|]}y_{j}(\mathbb{E}_{z}[q(z)]-q(\tilde{S}))\right]$
$=\mathbb{E}\left[\max_{y\in\Delta_{|{\cal Q}|}}F_{\mathcal{D}}(\tilde{x},y)\right]$
$\leq\mathbb{E}\left[\max_{y\in\Delta_{|{\cal Q}|},x\in\Delta_{|\mathcal{Z}|}}(F_{\mathcal{D}}(\tilde{x},y)-F_{\mathcal{D}}(x,\bar{y})).\right]$
where the inequality comes from the fact that $F_{\mathcal{D}}(x_{\mathcal{D}},\bar{y})=0$ (where $x_{\mathcal{D}}$ is the probability vector associated with the unknown distribution $\mathcal{D}$ supported on $\mathcal{Z}$), so $\min_{x\in\Delta_{|\mathcal{Z}|}}F_{\mathcal{D}}(x,\bar{y})\leq 0$. Hence, it suffices that we bound $\mathbb{E}\left[\operatorname{Gap}(\tilde{x},\bar{y})\right]$. To this end, first we bound $\mathbb{E}\left[\operatorname{Gap}(\bar{x},\bar{y})\right]$. Analogously to the proof of Lemma 4.1, we note that
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]\leq\operatorname{Regret}_{T}^{x}(w)+\operatorname{Regret}_{T}^{y}(v)+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle$
$\leq\frac{\log(|\mathcal{Z}|)}{\tau_{x}}+\frac{\tau_{x}}{2}\sum_{t\in[T]}\|g^{t}_{x}\|_{\infty}^{2}+\frac{\log(|{\cal Q}|)}{\tau_{y}}+\frac{\tau_{y}}{2}\sum_{t\in[T]}\|g^{t}_{y}\|_{\infty}^{2}+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle.$
Next, we have
$\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w,v)\rangle=\sum_{t=1}^{T}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w\rangle+\langle-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y},y^{t}-v\rangle$
$\leq\sum_{t=1}^{T}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w\rangle+2\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}$
$\leq\sum_{t=1}^{T}\langle(\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w^{t}\rangle)+\frac{\log(|\mathcal{Z}|)}{\tau_{x}}+\sum_{t=1}^{T}\tau_{x}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}+2\|\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty},$
where the first inequality follows from Holder’s inequality and the second from Lemma A.3, where $(v^{t})_{t\in[T]}$ are defined analogously to Lemma 4.1. Using these bounds above together with the fact that $L_{0}=2$, and maximizing over $w\in\Delta_{|\mathcal{Z}|},v\in\Delta_{|{\cal Q}|}$, we arrive to
$\operatorname{Gap}(\bar{x},\bar{y})$
$\leq\frac{2\log(|\mathcal{Z}|)}{\tau_{x}T}+\frac{\log(|{\cal Q}|)}{\tau_{y}T}+18\tau_{x}+2\tau_{y}+\frac{1}{T}\sum_{t=1}^{T}\langle\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x},x^{t}-w^{t}\rangle+2\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}.$
Taking expectations, we obtain
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{2\log(|\mathcal{Z}|)}{\tau_{x}T}+\frac{\log(|{\cal Q}|)}{\tau_{y}T}+18\tau_{x}+2\tau_{y}+\frac{2}{T}\sum_{t=1}^{T}\mathbb{E}[\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}].$
We bound the standard deviation of $g^{t}_{y}$ using vector martingale concentration in $\ell_{\infty}$ norm (see Lemma A.2). More concretely, denoting $Q(z)=(q_{1}(z),...,q_{|{\cal Q}|}(z))$ and $Q(S)=\frac{1}{n}\sum_{j\in[n]}Q(z_{i_{j}})$,
$\mathbb{E}[\|-\nabla_{y}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y}\|_{\infty}]$ $=\mathbb{E}[\|Q(S)-\mathbb{E}_{z}[Q(z)]\|_{\infty}]$
$\leq\sqrt{\mathbb{E}[\|Q(S)-\mathbb{E}_{z}[Q(z)]\|_{\infty}^{2}]}$
$=\sqrt{\mathbb{E}\left[\left\|\sum_{j\in[n]}\frac{Q(z_{i_{j}})-\mathbb{E}_{z}[Q(z)]}{n}\right\|_{\infty}^{2}\right]}$
$\leq\sqrt{2e\log(|{\cal Q}|)\sum_{j\in[n]}\left\|\frac{Q(z_{i_{j}})-\mathbb{E}_{z}[Q(z)]}{n}\right\|_{\infty}^{2}}$
$\leq 5\sqrt{\frac{\log(|{\cal Q}|)}{n}},$
where the first inequality follows from Jensen’s inequality and the second from Lemma A.2. We have proved the following bound for the expected gap
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{2\log(|\mathcal{Z}|)}{\tau_{x}T}+\frac{\log(|{\cal Q}|)}{\tau_{y}T}+18\tau_{x}+2\tau_{y}+10\sqrt{\frac{\log(|{\cal Q}|)}{n}}.$
Setting $T=\frac{6n\varepsilon}{16\sqrt{2\log(1/\delta)}\log(|{\cal Q}|)},\tau_{x}=\sqrt{\frac{\log(|\mathcal{Z}|)}{9T}},\tau_{y}=\frac{\log(|{\cal Q}|)}{6\sqrt{\log(|\mathcal{Z}|)T}}$ and using $\log(|{\cal Q}|)\leq C\log(|\mathcal{Z}|)$, we obtain
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\leq 10\sqrt{\frac{\log(|{\cal Q}|)}{n}}+2(12+C/3)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$
Also, following the proof of Lemma 4.2 together with the fact that $L_{1}=2$, we can show that
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\bar{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{4}{n}+\frac{8}{\lambda n}+\lambda\log(|{\cal Q}|).$
Using $\lambda=\sqrt{8}/\sqrt{\log(|{\cal Q}|)T}$ gives $\mathbb{E}[\operatorname{Gap}(\tilde{x},\bar{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{4}{n}+6\sqrt{\frac{\log(|{\cal Q}|)}{n}}$. Finally, we conclude that
$\mathbb{E}\left[\max_{q\in{\cal Q}}|\mathbb{E}_{z}[q(z)]-q(\tilde{S})|\right]\leq\mathbb{E}[\operatorname{Gap}(\tilde{x},\bar{y})]\leq 20\sqrt{\frac{\log(|{\cal Q}|)}{n}}+2(12+C/3)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$