### v2

# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially- Private Stochastic Saddle-Point Problems

Abstract
We study the problem of differentially-private (DP) stochastic (convex-concave) saddle-points in the $\ell_{1}$ setting. We propose $(\varepsilon,\delta)$-DP algorithms based on stochastic mirror descent that attain nearly dimension-independent convergence rates for the expected duality gap, a type of guarantee that was known before only for bilinear objectives. For convex-concave and first-order-smooth stochastic objectives, our algorithms attain a rate of $\sqrt{\log(d)/n}+(\log(d)^{3/2}/[n\varepsilon])^{1/3}$, where $d$ is the dimension of the problem and $n$ the dataset size. Under an additional second-order-smoothness assumption, we show that the duality gap is bounded by $\sqrt{\log(d)/n}+\log(d)/\sqrt{n\varepsilon}$ with high probability, by using bias-reduced gradient estimators. This rate provides evidence of the near-optimality of our approach, since a lower bound of $\sqrt{\log(d)/n}+\log(d)^{3/4}/\sqrt{n\varepsilon}$ exists. Finally, we show that combining our methods with acceleration techniques from online learning leads to the first algorithm for DP Stochastic Convex Optimization in the $\ell_{1}$ setting that is not based on Frank-Wolfe methods. For convex and first-order-smooth stochastic objectives, our algorithms attain an excess risk of $\sqrt{\log(d)/n}+\log(d)^{7/10}/[n\varepsilon]^{2/5}$, and when additionally assuming second-order-smoothness, we improve the rate to $\sqrt{\log(d)/n}+\log(d)/\sqrt{n\varepsilon}$. Instrumental to all of these results are various extensions of the classical Maurey Sparsification Lemma [Pisier:1980], which may be of independent interest.
keywords:
Differential Privacy, Stochastic Saddle Point Problem, Mirror Descent, Sparse Approximation
{MSCcodes}
90C47, 90C15, 62D99, 68P27
1 Introduction
In this work, we study the stochastic saddle-point (SSP) problem, that is, we aim to find a pair $(x^{*},y^{*})$ that solves
(1) $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\{F_{\mathcal{D}}(x,y):=\mathbb{E}_{z\sim\mathcal{D}}[f(x,y;z)]\},$
where the distribution $\mathcal{D}$ over the sample space $\mathcal{Z}$ is unknown, and only iid data from the target distribution, $S\sim\mathcal{D}^{n}$, are available. We assume that, for each $z\in\mathcal{Z}$, the function $(x,y)\to f(x,y;z)$ is convex-concave, and the function is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t $\|\cdot\|_{1}$ (see Section 1.3 for details). We also assume that the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ are the standard simplices in $\mathbb{R}^{d_{x}}$ and $\mathbb{R}^{d_{y}}$, denoted by $\Delta_{x}$ and $\Delta_{y}$ from now on [^1] Note that we can reduce SSP problems over polytope feasible sets to this setting. In this case, the dimensions $d_{x}$ and $d_{y}$ will correspond to the number of vertices of $\mathcal{X}$ and $\mathcal{Y}$, respectively.. We will refer to SSP problems satisfying these assumptions as the $\ell_{1}$ setting’ [JNLS:2009, NesterovNemirovski:2013]. See Section 1.4 for important problems where these assumptions hold.
In modern machine learning applications, the observed dataset $S$ usually contains sensitive user information. In this context, Differential Privacy (DP) has become the de facto standard for privacy-preserving algorithms. A randomized algorithm $\mathcal{A}(S)$ is $(\varepsilon,\delta)$-differentially private if for any pair of datasets $S$ and $S^{\prime}$ differing in at most one data point and any event $\mathcal{E}$ in the output space,
$\mathbb{P}[\mathcal{A}(S)\in\mathcal{E}]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(S^{\prime})\in\mathcal{E}]+\delta.$
We study the Differentially Private Stochastic Saddle Point (DP-SSP) problem in the $\ell_{1}$ setting. More concretely, we provide differentially private algorithms that receive $S$ as input and return a pair $(x_{\mathcal{A}(S)},y_{\mathcal{A}(S)})$ that approximates an SSP. Given a pair $(x,y)$, we measure its efficiency by the duality gap, defined as
(2) $\operatorname{Gap}(x,y)=\max_{v\in\mathcal{X},w\in\mathcal{Y}}(F_{\mathcal{D}}(x,w)-F_{\mathcal{D}}(v,y)).$
There is a major discrepancy in the literature between the achievable rates of convergence for the duality gap between some structured problems and more general convex-concave settings . While some bilinear problems have convergences rates which scale only poly-logarithmically with the ambient dimensions
(e.g.,  [Hsu:2013]), all known rates for general convex-concave objectives scale polynomially with the dimension (e.g., [pmlr-v195-bassily23a]). This can be attributed to the fact that all existing results for convex-concave settings work under Euclidean geometry, which assumes general convex and compact feasible sets and objectives that are Lipschitz and smooth with respect to $\|\cdot\|_{2}$, as opposed to the $\ell_{1}$ setting considered in this work. Therefore, known polynomial dimension lower bounds for DP Stochastic Convex Optimization (SCO) in the Euclidean setting [Bassily:2014] apply. Recently, [asi2021private, bassily2021differentially] showed that it is possible to obtain excess risk bounds for DP-SCO under $\ell_{1}$ setting which are nearly independent of the dimension. Inspired by these recent developments, in this work we provide DP algorithms with nearly dimension-independent rates for convex-concave SSP problems in the $\ell_{1}$ setting. Our results extend considerably the class of SSP problems where it is possible to bypass polynomial in the dimension lower bounds, which so far only contain problems with bilinear objectives and simplex feasible sets.
1.1 Summary of Contributions
We prove the first nearly-dimension independent rates for convex-concave DP-SSP in the $\ell_{1}$ setting. These rates are optimal (up to logarithmic factors) for second-order-smooth objectives, and have the form $r_{NDP}+r_{DP}$ where $r_{NDP}$ is the optimal non-private SSP rate and $r_{DP}$ is an error rate due to differential privacy. Furthemore, the algorithms achieving these rates have an $O(n)$ gradient complexity. We also use our techniques to show nearly-dimension independent convergence rates for DP-SCO under the same assumptions considered in [asi2021private, bassily2021differentially]. Previous work used private versions of the variance-reduced stochastic Frank-Wolfe algorithm. By contrast, we are the first to achieve this using variants of mirror-descent algorithms. Table 1 summarizes the obtained rates.
  • We propose two algorithms for DP-SSP, based on (entropic) Stochastic Mirror Descent (SMD). As opposed to previous results for bilinear settings where privatization by sampling from vertices leads to unbiased gradient estimators [Hsu:2013], the main challenge we encounter is controlling the bias in the more general (stochastic) convex-concave setting. The first algorithm (Section 3.1) controls bias by simply using a large number of samples in every iteration. The second algorithm (Section 3.2) builds upon the first one by incorporating a stochastic simulation technique from [Blanchet:2015] to reduce the bias of gradients evaluated at the private estimators of the iterates.
  • In our algorithms, the iterates are privately estimated by sampling vertices proportionally to their weights, a technique known as Maurey’s Sparsification Lemma [Pisier:1980] [^2] This method is also known as the Maurey Empirical Method.. We provide various extensions of this classsical result, most notably for the duality gap and gradient sampling from vertices, that are new and we believe are of independent interest. These extensions are challenging, primarily due to the presence of bias in the underlying estimation task. See Section 2 and Subsection 3.1.2 for more details.
  • We give the first mirror descent style algorithm for DP-SCO in the polyhedral setting with nearly dimension-independent excess risk guarantees by combining the ideas we used for DP-SSP with the Anytime Online-to-Batch Conversion from [anytime-online-to-batch]. See Section 4. While these rates are not optimal, we expect they could inspire future work into more practical algorithms.
Setting Objective Previous best-known rate Our rate
SSP Bilinear $\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}\text{\tiny{(*)}}$ [Hardt:2010] $\color[rgb]{0,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$(*) (Thm. 3.1)
Quadratic     None $\color[rgb]{0,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$(*) (Thm. 3.1)
$L_{2}$-smooth     None $\color[rgb]{0,0,0}\frac{\ell}{\sqrt{n}}+\left(\frac{\ell^{2}}{n\varepsilon}\right)^{1/2}$(**) (Thm. 3.14)
$L_{1}$-smooth     None $\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/3}$ (Thm. 3.1)
SCO $L_{2}$-smooth     None $\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}}{\sqrt{n\varepsilon}}$ (Thm. 4.2)
$L_{1}$-smooth $\color[rgb]{0,0,0}\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{2/3}}{(n\varepsilon)^{2/3}}$(*) [asi2021private] $\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{7/10}}{(n\varepsilon)^{2/5}}$ (Thm. 4.2)
Table 1: Rates derived from our results. We omit polylog($n,1/\delta$) factors, function-class parameters and constants. We refer to first-order-smooth and second-order-smooth functions as $L_{1}$-smooth and $L_{2}$-smooth, respectively. We denote $\ell=\log(d_{x})+\log(d_{y})$ and $\ell_{x}=\log(d_{x})$. (*): these rates are optimal up to polylog factors. (**): this rate holds with high probability while the rest hold in expectation.
1.2 Related Work
DP stochastic saddle-points
Stochastic (non-private) saddle-point problems have connections to optimization, economics, machine learning, and many other fields. One such algorithm for solving these problems is the stochastic entropic mirror-descent algorithm [JNLS:2009], which is the basis for much of this work. Recently, a line of research has studied the differentially private (DP) version of stochastic saddle-point problems [boob2024optimal, DPSMO, DPSGDA, Kang:2022, pmlr-v195-bassily23a]. These works focus exclusively on Euclidean setups, which necessarily leads to algorithms that have a polynomial dependence on the dimension. Moreover, due to the Euclidean setup, the techniques used in these works do not necessarily apply to the setting considered here. In addition, with only limited exceptions [boob2024optimal, pmlr-v195-bassily23a], most previous work provide guarantees on the weak duality gap. The (strong) duality gap (2), on the other hand, is more meaningful (see [pmlr-v195-bassily23a] for justification), but also more difficult to bound, most notably in the private setting. This is the convergence criterion that we use in this work.
Optimization in the $\ell_{1}$ setting
Closest to this work is DP stochastic convex optimization (SCO) in the $\ell_{1}$ setting [asi2021private, bassily2021non, Han:2022], as well as earlier work on private empirical risk minimization (ERM) [Jain:2014, Talwar:2014, Talwar:2015]. These works attain nearly dimension-independent excess risk rates with private variants of the (stochastic) Frank-Wolfe method. Even though this method leads to optimal rates for DP-SCO, it is not known to yield any nontrivial rates for saddle-point problems. Even without DP constraints, it is not known how to solve SSP problems with Frank-Wolfe methods unless strong additional assumptions are made: namely, either simultaneously assuming smoothness and strong convexity/strong concavity [^3] These assumptions are indeed strong: If $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ is an $L_{1}$-smooth and $\mu$-strongly convex function w.r.t. $\|\cdot\|_{1}$, then $L_{1}/\mu=\Omega(d)$. This follows by conjugate duality and the argument in [dAspremont:2018, Example 4.1], or strong convexity of the feasible sets [Gidel:2017]. Clearly, these assumptions are not suitable for this work.
Maurey Sparsification
This is a classical technique in high-dimensional convex geometry [Pisier:1980]. The idea is based on sampling vertices with the barycentric coordinates weights and taking the empirical average, as nicely illustrated in the introduction of the book [Vershynin:2018]. This approximation is established for norms with concentration estimates for independent random sums, e.g., quantified by their Rademacher type [Pisier:2016]. This technique has been used systematically as a compression-type argument to obtain faster rates in statistical estimation and learning (e.g.,  [Barron:1993, Zhang:2002]). More recently, [Acharya:2019] proved a result similar to Lemma 2.1, a functional approximation for randomized sparsification, in the context of model compression for distributed learning of generalized linear models. They also proposed a sparsified mirror descent algorithm that works in the $\ell_{1}/\ell_{q}$ setting where $q<\infty$, resorting on smoothness properties of the Bregman divergence. Furthemore, they only sparsify the outputs of each machine. We do not follow this approach. Instead, we sparsify every iterate of mirror descent. Furthermore, we show that entropic SMD (using a nonsmooth Bregman divergence) with iterate sparsification works for convex and smooth losses, which to our knowledge is an entirely new result, and does not follow from direct ex-post model compression.
Debiasing
This is a technique proposed in the stochastic simulation literature [Blanchet:2015], to leverage biased estimators whose accuracy improves with larger sample size. The main difference between the works on simulation and our approach is the finite sample size, which limits the amount of bias-reduction. The work [Asi:2021bias] studied bias-reduction in regularization for SCO. In concurrent work [Ghazi:2024], a gradient debiasing procedure is developed for Euclidean DP-SCO. Although our results do not depend on their techniques, we borrow some inspiration from them. Note futher that most of our results do not use bias-reduction.
Concurrent and follow up work
In independent and concurrent work, the paper [Zhou:2024] investigates empirical and population worst-group risk minimization under DP, with a focus on group fairness and robust learning. While the convex-linear loss functions of their work correspond to a particular case of ours (convex-concave), their use of Euclidean geometry makes their results incomparable to ours. Furthermore, their sequential regularization approach is independent from ours. The follow up work [BGM24noneuclideanDP-SSP] considers DP-SSPs in $\ell_{p}/\ell_{q}$ setup with $p,q\in[1,2]$. In contrast to our work, their results focus exclusively in the nonsmooth setting, leading to unavoidable polynomial-in-the-dimension convergence rates for the strong duality gap.
1.3 Preliminaries
Given $x\in\mathbb{R}^{d}$, we denote the $1$-norm and $\infty$-norm by $\|x\|_{1}=\sum_{j\in[d]}|x_{j}|$ and $\|x\|_{\infty}=\max_{j\in[d]}|x_{j}|$, respectively. A differentiable function $F:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{0}$-Lipschitz w.r.t. $\|\cdot\|_{1}$ over a set $\mathcal{X}$ if $\|\nabla F(x)\|_{\infty}\leq L_{0}$ for all $x\in\mathcal{X}$. It is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$ over $\mathcal{X}$ if $\|\nabla F(x^{1})-\nabla F(x^{2})\|_{\infty}\leq L_{1}\|x^{1}-x^{2}\|_{1}$ for all $x^{1},x^{2}\in\mathcal{X}$. We say a function is first-order-smooth if it is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. We say a function is second-order-smooth if it is first-order-smooth and its partial derivatives are $L_{2}$-smooth w.r.t $\|\cdot\|_{1}$. [^4] Note that this assumption is milder than the typical second-order-smoothness assumption in optimization, and it is satisfied by the widely used quadratic and logistic losses. For $a,b:\mathbb{R}^{p}\mapsto\mathbb{R}$, we say $a\lesssim b$ if there exists an absolute constant $C>0$ such that for all $x\in\mathbb{R}^{p}$, $a(x)\leq Cb(x)$. When $a\lesssim b$ and $b\lesssim a$, we write $a\eqsim b$. We denote by $e_{i}$ the $i$-th canonical vector (the dimension being clear from context). The standard simplex in $\mathbb{R}^{d}$ is the set $\Delta_{d}=\{x\in\mathbb{R}^{d}:\|x\|_{1}=1,x_{j}\geq 0\text{ for all }j\in[d]\}$. A vector $x\in\Delta_{d}$ induces a probability distribution over the canonical vectors, which we denote by $P_{x}$. Hence, $\hat{x}\sim P_{x}$ means that $\hat{x}$ takes the value $e_{i}$ with probability $x_{i}$. Note that if $\hat{x}\sim P_{x}$, then $\mathbb{E}[\hat{x}]=x$. Given $f:{\cal X}\times{\cal Y}\mapsto\mathbb{R}$, we let $\Phi(x,y)=\big(\nabla_{x}f(x,y),-\nabla_{y}f(x,y)\big)$ be saddle operator (a.k.a. the monotone operator associated to the saddle-point problem) [Rockafellar:1970]. We denote $\ell=\log(d_{x})+\log(d_{y})$, where $d_{x},d_{y}$ are the ambient dimensions of the optimization variables in the SSP problem.
1.4 Motivating Examples
As motivating examples for the DP-SSP problem (1), we present two problems of interest in privacy-preserving machine learning and data analysis that fit into our problem setup.
DP Synthetic Data Generation
Consider a finite sample space $\mathcal{Z}=\{z_{1},...,z_{|\mathcal{Z}|}\}$ and an unknown distribution $\mathcal{D}$ supported on $\mathcal{Z}$. Suppose we have a dataset $S=\{z_{i_{1}},...,z_{i_{n}}\}\overset{iid}{\sim}\mathcal{D}$ and a set of queries $\mathcal{Q}=\{q:\mathcal{Z}\to[-1,1]\}$. Defining $q(S)=\frac{1}{|S|}\sum_{z\in S}q(z)$ and $q(\mathcal{D})=\mathbb{E}_{z\sim\mathcal{D}}[q(z)]$, our goal is to construct a synthetic dataset $\tilde{S}$ such that $\tilde{S}$ is private with respect to $S$ and $\max_{q\in{\cal Q}}|q(\tilde{S})-q(\mathcal{D})|$ is as small (in expectation) as possible. Using a linear relaxation for the discrete maximum over $\cal Q$} and viewing each query $q$ as vector $(q(z))_{z\in\mathcal{Z}}$, we obtain a bilinear SSP formulation as follows:
(3) $\min_{x\in\Delta_{|\mathcal{Z}|}}\max_{y\in\Delta_{|\mathcal{Q}|}}\mathbb{E}_{z\sim\mathcal{D}}\Big[\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)\Big].$
While our Algorithm 1, proposed in Section 3.1, can solve this problem, applying it directly would result in a slightly suboptimal dependence on the histogram size, $|{\cal Z}|$. However, a variant of it that leverages the specific structure of this problem attains the optimal rates [lower_bound_DP_query_answering]. More concretely, the following theorem holds.
Theorem 1.1.
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. If $|\mathcal{Q}|\leq|\mathcal{Z}|^{C}$, then a variant of Algorithm 1 constructs an $(\varepsilon,\delta)$-DP synthetic dataset $\tilde{S}$ of $n$ samples, such that
$\mathbb{E}\Big[\max_{q\in\mathcal{Q}}|\mathbb{E}_{z\sim\mathcal{D}}[q(z)]-q(\tilde{S})|\Big]\lesssim\sqrt{\frac{\log(|{\cal Q}|)}{n}}+(1+C)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$
Furthermore, this rate is optimal up to constant factors.
Minimizing the maximal loss under DP
Consider a collection of convex losses $f_{i}:\mathcal{X}\times\mathcal{Z}\to[-B,B]$, $i\in[d_{y}]$. Suppose $\mathcal{X}$ is a polytope. We are interested in minimizing $\max_{i\in[d_{y}]}\{F_{i}(x):=\mathbb{E}_{z\sim\mathcal{D}}[f_{i}(x;z)]\}$. This can be formulated as follows:
$\min_{x\in\mathcal{X}}\{\max_{i\in[d_{y}]}F_{i}(x)\}=\min_{x\in\mathcal{X}}\max_{y\in\Delta_{y}}\mathbb{E}_{z\sim\mathcal{D}}\left[\sum_{i\in[d_{y}]}y_{i}f_{i}(x;z)\right].$
The maximal loss formulation encompasses applications to fairness [Jagielski:2019, Fair_FL, Lowy:2023] and robustness [Mohajerin:2018] under privacy constraints. As a remark, if the functions $(f_{i})_{i\in[d_{y}]}$ are $L_{0}$-Lipschitz and $L_{1}$-smooth on $x$ w.r.t $\|\cdot\|_{1}$, then the problem objective $f(x,y;z)=\sum_{i\in[d_{y}]}y_{i}f_{i}(x;z)$ is convex-concave, $\max\{L_{0},B\}$-Lipschitz, $\max\{L_{0},L_{1}\}$-smooth on $(x,y)$. If the functions $(f_{i})_{i\in[d_{y}]}$ are also $L_{2}$-second-order-smooth on $x$, then $f(x,y)$ is $\max\{L_{1},L_{2}\}$-second-order-smooth on $(x,y)$. Examples of convex losses that are Lipschitz, smooth and second-order-smooth over compact sets are the quadratic and the logistic loss.
2 Maurey Sparsification with Functional Value Approximation
The classical Maurey Sparsification Lemma [Pisier:1980] is an application of the probabilistic method that leads to an approximate and (nearly) dimension-independent version of Carathéodory’s Theorem [Carathodory:1911]. While this method is well-known to apply with smooth norms, it appears that for norms such as $\|\cdot\|_{1}$ such approximations are out of reach. [^5] In fact, $\|\cdot\|_{1}$ is as far as being smooth as possible, in the sense of having trivial Rademacher type [Pisier:2016]. Furthermore, polynomial in $d$ mean estimation lower bounds with respect to $\|\cdot\|_{1}$ are known [Waggoner:2015]. We circumvent this limitation by directly working with function value approximations. These results will be crucial for the analysis of our algorithms in the forthcoming sections.
We start with a simple adaptation of Maurey’s method for function approximation in the general smooth case. A similar proof appears in [Acharya:2019]. Throughout this section we denote $\bar{x}^{T}=\sum_{t=1}^{T}\frac{x^{t}}{T}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\frac{a^{t}}{T}$ unless explicitly said otherwise.
Lemma 2.1 (Function value approximation in expectation).
Let $(\mathbf{E},\|\cdot\|)$ be a normed space, and let $F$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$, and $a^{1},...,a^{T}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$. Let $\bar{x}^{T}=\sum_{t=1}^{T}\lambda_{t}x^{t}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\lambda_{t}a^{t}$. Then $\big|\mathbb{E}\left[F\left(\bar{a}^{T}\right)-F\left(\bar{x}^{T}\right)\right]\big|\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$.
Proof 2.2.
Let $\alpha_{t}=\Big\langle\nabla F\Big(\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\Big),\lambda_{t}(a^{t}-x^{t})\Big\rangle$. Noting that $F(\bar{a}^{T})=F(\bar{x}^{T}+\bar{a}^{T}-\bar{x}^{T})$, and using repeatedly the $L_{1}$-smoothness of $F$, we have
$F(\bar{a}^{T})$ $\leq F(\bar{x}^{T}+\bar{a}^{T-1}-\bar{x}^{T-1})+\alpha_{T}+\frac{L_{1}\lambda_{T}^{2}\left\|a^{T}-x^{T}\right\|^{2}}{2}$
$\ldots$ $\leq F\left(\bar{x}^{T}\right)+\sum_{t\in[T]}\alpha_{t}+\frac{L_{1}}{2}\sum_{t\in[T]}\lambda_{t}^{2}\|a^{t}-x^{t}\|^{2}.$
Clearly $(\alpha_{t})_{t\in[T]}$ is a martingale difference sequence. Hence, $\mathbb{E}\big[\sum_{t\in[T]}\alpha_{t}\big]=0$ and consequently $\mathbb{E}[F(\bar{a}^{T})-F(\bar{x}^{T})]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$. Since F-F is also $L_{1}$-smooth, the same idea yields $\mathbb{E}[F(\bar{x}^{T})-F(\bar{a}^{T})]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$.
We now instantiate Lemma 2.1 for the space $(\mathbb{R}^{d},\|\cdot\|_{1})$ with $\lambda_{i}=\frac{1}{T}$, $i\in[T]$. In later sections we will see that $x^{t}\in\Delta_{d}$ represents the $t$-th iterate of our algorithm and $a^{t}\sim P_{x^{t}}$ a private, unbiased estimator of $x^{t}$.
Corollary 2.3 (Function value approximation w.r.t. $\|\cdot\|_{1}$ over the simplex).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}\in\Delta_{d}$ be deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then $\left|\mathbb{E}\left[F\left(\bar{a}^{T}\right)-F\left(\bar{x}^{T}\right)\right]\right|\leq\frac{2L_{1}}{T}$.
The above result bounds the bias incurred by evaluating gradients at sampled vertices. More concretely, the following corollaries hold.
Corollary 2.4 (Gradient bias with second-order-smoothness).
Suppose $F:\mathbb{R}^{d}\to\mathbb{R}$ has partial derivates that are $L_{2}$-smooth w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}\in\Delta_{d}$ be deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then $\left\|\mathbb{E}\left[\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right]\right\|_{\infty}\leq\frac{2L_{2}}{T}$.
Proof 2.5.
This follows directly from applying Corollary 2.3 to each partial derivate: for all $j\in[d]$, we have $\left|\mathbb{E}\left[\nabla_{j}F\left(\bar{a}^{T}\right)-\nabla_{j}F\left(\bar{x}^{T}\right)\right]\right|\leq\frac{2L_{2}}{T}$.
Corollary 2.6 (Gradient bias with first-order-smoothness).
Suppose $F:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$.Let $x^{1},...,x^{T}$ be a collection of deterministic vectors in $\Delta_{d}$ and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then $\left\|\mathbb{E}\left[\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right]\right\|_{\infty}\leq\frac{4L_{1}}{\sqrt{T}}$.
Proof 2.7.
Denote $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t},\bar{a}=\frac{1}{T}\sum_{t\in[T]}a^{t}$. By $L_{1}$-smoothness of $F$, we have for all $j\in[d],r\in\mathbb{R}^{+}$: $F(\bar{x}+re_{j})\leq F(\bar{x})+\langle\nabla F(\bar{x}),re_{j}\rangle+\frac{L_{1}r^{2}}{2}\|e_{j}\|_{1}^{2}$. Re-arranging this expression gives $\langle\nabla F(\bar{x}),e_{j}\rangle\leq\frac{F(\bar{x})-F(\bar{x}+re_{j})}{r}+\frac{L_{1}r}{2}$. By $L_{1}$-smoothness of F-F, we can obtain $\langle\nabla F(\bar{a}),e_{j}\rangle\leq\frac{F(\bar{a}+re_{j})-F(\bar{a})}{r}+\frac{L_{1}r}{2}$. Adding these inequalities and taking expectations we get
$\mathbb{E}[\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle]\leq\frac{\mathbb{E}\left[F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})\right]}{r}+L_{1}r.$
By applying Corollary 2.3 to $F(\cdot)$ and $F(\cdot+re_{j})$ we obtain
$\mathbb{E}\left[F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})\right]\leq 4L_{1}/T.$
Hence, for all $j\in[d]$, $\mathbb{E}[\nabla_{j}F(\bar{a})-\nabla_{j}F(\bar{x})]=\mathbb{E}[\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle]\leq\frac{4L_{1}}{rT}+L_{1}r.$ Analogously we can prove the same bound for $\mathbb{E}[\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle]$. Finally, choosing $r=2/\sqrt{T}$ gives the result.
We will also need to bound the second-moment of the error incurred by evaluating gradients at sampled vertices. Towards this, we first observe that the in-expectation guarantees obtained in Lemma 2.1 can be turned into high-probability guarantees by controlling the increments of the martingale difference sequence $(\alpha_{t})_{t\in[T]}$ with the bounded differences inequality.
Lemma 2.8 (Function value approximation with high probability).
Let $(\mathbf{E},\|\cdot\|)$ be a normed space and $F$ be an $L_{0}$-Lipschitz, $L_{1}$-smooth function w.r.t. $\|\cdot\|$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$ and $a^{1},...,a^{T}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$. Suppose $\|x^{t}-a^{t}\|\leq D$ a.s. for all $t\in[T]$. Define $\bar{x}^{T}=\sum_{t=1}^{T}\lambda_{t}x^{t}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\lambda_{t}a^{t}$. Then, for any $\beta>0$
$\mathbb{P}\left[F\left(\bar{a}^{T}\right)-F\left(\bar{x}^{T}\right)\geq\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\beta\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\,\right]\leq\exp(-\beta^{2}).$
Proof 2.9.
Let $\alpha_{t}=\lambda_{t}\left\langle\nabla F\left(\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\right),a^{t}-x^{t}\right\rangle$ as in the proof of Lemma 2.1. From that proof we have that $F(\bar{a}^{T})-F\left(\bar{x}^{T}\right)\leq\sum_{t=1}^{T}\alpha_{t}+\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}$. Hence
(4) $\mathbb{P}[F(\bar{a}^{T})-F\left(\bar{x}^{T}\right)\geq\alpha]\leq\mathbb{P}\left[\sum_{t=1}^{T}\alpha_{t}+\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\geq\alpha\right]=\mathbb{P}\left[\sum_{t\in[T]}\alpha_{t}\geq\alpha^{\prime}\right],$
where $\alpha=\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\alpha^{\prime}$. Next, we bound the probability on the right-hand side. Note that $(\alpha_{t})_{t\in[T]}$ is a martingale difference sequence and $|\alpha_{i}|\leq L_{0}D\lambda_{i}$, so the bounded differences inequality gives $\mathbb{P}\left[\sum_{t=1}^{T}\alpha_{t}\geq\beta\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\right]\leq\exp(-\beta^{2})$, for any $\beta>0$. Making $\alpha^{\prime}=\beta\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}$ and replacing back into Equation 4, we conclude the claimed result.
We now use Lemma 2.8 to give bounds on the expected second-moment error, with respect to $\|\cdot\|_{\infty}$.
Lemma 2.10 (Second moment of maximum expected error).
Let $(\mathbf{E},\|\cdot\|)$ be a normed space and $F_{1},...,F_{M}$ a collection of $L_{0}$-Lipschitz and $L_{1}$-smooth functions w.r.t $\|\cdot\|$. Let $x^{1},\ldots,x^{T}\in\mathbb{R}^{d}$, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$ and $a^{1},...,a^{T}$ be a collection of random vectors such that $\mathbb{E}[a^{t}|a^{t-1},\ldots,a^{1}]=x^{t}$ and $\|a^{t}-x^{t}\|\leq D$ a.s. for all $t\in[T]$. Let $\bar{a}^{T}=\sum_{t\in[T]}\lambda_{t}a^{t}$ and $\bar{x}^{T}=\sum_{t\in[T]}\lambda_{t}x^{t}$. Then
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}\right]\leq\frac{L_{1}^{2}D^{4}}{2}\left(\sum_{t=1}^{T}\lambda_{t}^{2}\right)^{2}+2L_{0}^{2}D^{2}(4+\log(M))\sum_{t=1}^{T}\lambda_{t}^{2}.$
Proof 2.11.
Define the constants $A=\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}$ and $C=\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}$. We will introduce a quantity, $B$, in the following computation, and later optimize over the choice of this $B$. We proceed as follows:
$\mathbb{E}\bigg[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}$ $\bigg]=\int_{0}^{\infty}\mathbb{P}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|\geq\beta\right]\beta d\beta$
$\hskip-65.44142pt\leq\int_{0}^{A+B}\beta d\beta+\sum_{j\in[M]}\int_{A+B}^{\infty}\mathbb{P}\left[|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|\geq\beta\right]\beta d\beta$
$\hskip-65.44142pt=\frac{(A+B)^{2}}{2}+\sum_{j\in[M]}\int_{B/C}^{\infty}\mathbb{P}\left[|F_{j}(\bar{x})-F_{j}(\bar{a}^{T})|\geq A+C\beta\right](A+C\beta)Cd\beta$
$\hskip-65.44142pt\leq\frac{(A+B)^{2}}{2}+2MC\left[C\int_{B/C}^{\infty}\exp\left(-\beta^{2}\right)\beta d\beta+A\int_{B/C}^{\infty}\exp\left(-\beta^{2}\right)d\beta\right]$
$\hskip-65.44142pt=\frac{(A+B)^{2}}{2}+2MC\left[C\exp\left(-(B/C)^{2}\right)+A\sqrt{\pi}\mathbb{P}[X\geq B/C]\right],$
where the first inequality comes from the union bound, the second from Lemma 2.8, and in the last line, we use density of a normal random variable, $X\sim{\cal N}(0,1/2)$. From Mill’s inequality, we have $\mathbb{P}[X\geq B/C]\leq\frac{C}{B\sqrt{\pi}}\exp\left(-(B/C)^{2})\right)$. Hence,
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}\right]\leq\frac{(A+B)^{2}}{2}+2MC(C+A)\exp\left(-(B/C)^{2}\right).$
Finally, setting $B=C\sqrt{\log(M)}$ if $M>1$ and $B=C$ if $M=1$ gives
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}\right]\leq 2A^{2}+(4+\log(M))C^{2}.$
We conclude this section by presenting two corollaries of Lemma 2.10.
Corollary 2.12 (Gradient approximation error with second-order-smoothness).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$ with $L_{2}$-smooth partial derivatives. Let $x^{1},...,x^{T}\in\Delta_{d}$ be a collection of deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then
$\mathbb{E}\left\|\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right\|^{2}_{\infty}\leq\frac{8L_{2}^{2}}{T^{2}}+\frac{8L_{1}^{2}(4+\log(d))}{T}.$
Proof 2.13.
Apply Lemma 2.10 with $F_{j}=\nabla_{j}F$, $j\in[d]$.
Corollary 2.14 (Gradient approximation error with first-order-smoothness).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a function that is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}\in\Delta_{d}$ be a collection of deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for every $t\in[T]$. Then
$\mathbb{E}\left\|\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right\|^{2}_{\infty}\leq\frac{8\sqrt{2}L_{1}^{2}}{T^{3/2}\sqrt{4+\log(d)}}+\frac{8\sqrt{2}(L_{0}^{2}+L_{1}^{2})\sqrt{(4+\log(d))}}{\sqrt{T}}.$
Proof 2.15.
Similarly to the proof of Corollary 2.6, using $L_{1}$-smoothness of $F$,
$\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle$ $\leq\frac{F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})}{r}+L_{1}r$
$\leq\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|}{r}+L_{1}r.$
Since F-F is also $L_{1}$-smooth, then
$\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle\leq\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|}{r}+L_{1}r.$
We conclude that
$\left\|\nabla F\left(\bar{a}\right)-\nabla F\left(\bar{x}\right)\right\|^{2}_{\infty}$ $=\max_{j\in[d]}\left|\nabla_{j}F\left(\bar{a}\right)-\nabla_{j}F\left(\bar{x}\right)\right|^{2}=\max_{j\in[d]}\left|\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle\right|^{2}$
$\hskip-28.45274pt\leq 2\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}+L_{1}^{2}r^{2}.$
By Lemma 2.10 with $F_{j}(\cdot)=F(\cdot+re_{j})-F(\cdot)$ and $M=d$, we have
$\mathbb{E}\left[\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}\right]\leq\frac{32L_{1}^{2}}{T^{2}r^{2}}+\frac{32L_{0}^{2}(4+\log(d))}{Tr^{2}}.$
Putting everything together and taking expectations
$\mathbb{E}[\left\|\nabla F\left(\bar{a}\right)-\nabla F\left(\bar{x}\right)\right\|^{2}_{\infty}]\leq\frac{64L_{1}^{2}}{T^{2}r^{2}}+\frac{64L_{0}^{2}(4+\log(d))}{Tr^{2}}+2L_{1}^{2}r^{2}.$
Choosing $r^{2}=\sqrt{\frac{32(4+\log(d))}{T}}$ gives the bound.
3 Differentially Private Stochastic Saddle Points
In this section we state and prove our main results regarding the DP-SSP problem in $\ell_{1}$ setting. Given a batch $B$ of iid samples from $\mathcal{D}$, we will denote $\nabla F_{\mathcal{D}}(x,y;B):=\frac{1}{|B|}\sum_{z\in B}\nabla f(x,y;z)$, and similarly for partial derivatives.
3.1 Private Stochastic Mirror Descent with Sampling from Vertices
Our first algorithm is a (entropic) stochastic mirror descent (SMD) method with sampling from vertices; see Algorithm 1. The following theorem holds.
Algorithm 1 Private stochastic mirror descent with vertex sampling
0: Dataset $S$, Number of steps $T$, Step-size $\tau$, Number of iterate samples $K$
[…]
2:for $t=1$ to $T$ do
3:  $(\hat{x}^{t},\hat{y}^{t})=\frac{1}{K}\sum_{k\in[K]}(\hat{x}^{t,k},\hat{y}^{t,k})$, where $(\hat{x}^{t,k},\hat{y}^{t,k})_{k\in[K]}\overset{\text{iid}}{\sim}P_{x^{t}}\times P_{y^{t}}$
4:  Let $B^{t}$ be a batch of $B=n/T$ fresh samples
5:  $g^{t}_{x}=\nabla_{x}F_{\mathcal{D}}(\hat{x}^{t},\hat{y}^{t};B^{t})$, $g^{t}_{y}=-\nabla_{y}F_{\mathcal{D}}(\hat{x}^{t},\hat{y}^{t};B^{t})$
6:  $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau g^{t}_{x,j}\right),\quad\forall j\in[d_{x}]$  and  $y^{t+1}_{i}\propto y^{t}_{i}\exp\left(\tau g^{t}_{y,i}\right),\quad\forall i\in[d_{y}]$
7:end for
8:return $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$, where $(\tilde{x}^{t},\tilde{y}^{t})_{t\in[T]}\overset{\text{iid}}{\sim}P_{x^{t}}\times P_{y^{t}}$
Theorem 3.1 (Guarantees for Algorithm 1).
[…]
(5) $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\ell}{n}}+\bigg(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/3}\bigg].$
If in addition $f(\cdot,\cdot;z)$ has $L_{2}$-smooth partial derivatives, then there exists a different setting for $T,\tau$ and $K$ such that Algorithm 1 is $(\varepsilon,\delta)$-DP and its output satisfies
(6) $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1}+L_{2})\bigg[\sqrt{\frac{\ell}{n}}+\left(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{2/5}\bigg].$
[…]
(7) $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\bigg[\sqrt{\frac{\ell}{n}}+\left(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg].$
Remark 3.2.
The duality gap bound for $L_{2}=0$ (i.e. quadratic objective) in Theorem 3.1 is optimal up to constant factors in the case where $\log(d_{x})\simeq\log(d_{y})$. This follows from the lower bound in [lower_bound_DP_query_answering] with a resampling argument from [Bassily:2019]. Moreover, the proof of this rate only requires $L_{2}\lesssim(L_{0}+L_{1})\Big\{\sqrt{\ell/n}+\sqrt{\ell^{3/2}\sqrt{\log(1/\delta)}/(n\varepsilon)}\Big\}$, so the optimal rates hold for more general nearly quadratic objectives.
3.1.1 Privacy Analysis of Algorithm 1
We prove that the algorithm is DP. More concretely, we provide a condition on the stepsize that suffices to get privacy.
Proposition 3.3.
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. Then, Algorithm 1 is $(\varepsilon,\delta)$-DP if $\tau\leq\frac{B\varepsilon}{16L_{0}\sqrt{T(K+1)\log(1/\delta)}}$.
The proof relies on the privacy guarantees of the exponential mechanism and on the advanced composition property of DP. For completeness, we include them below.
Theorem 3.4 (Exponential mechanism (Theorem 3.10 in [dwork2014DPfoundations])).
Consider a (score) function $s:{\cal Z}^{n}\times J\mapsto\mathbb{R}$, where $J=\{e_{i}\}_{i\in[d]}$. Define the sensitivity of $s$ as $\Delta_{s}:=\max_{e\in J}\max_{S,S^{\prime}\in\mathcal{Z}^{n}\text{ differing in }\leq 1\text{ element}}|s(S,e)-s(S^{\prime},e)|$. Then, sampling an object $e$ from $J$ according to $\mathbb{P}(e=e_{i})\propto\exp\left(\frac{\varepsilon s(S,e_{i})}{2\Delta_{s}}\right)$ is $\varepsilon$-DP. As a consequence, sampling $e$ according to $\mathbb{P}(e=e_{i})\propto\exp\left(s(S,e_{i})\right)$ is $2\Delta_{s}$-DP.
Theorem 3.5 (Fully Adaptive Composition [Whitehouse:2023]).
Suppose $(\mathcal{A}_{n})_{n\geq 1}$ is a sequence of algorithms such that, for any $n\geq 1$, $\mathcal{A}_{n}$ is $(\varepsilon_{n},\delta_{n})$-DP with respect to a dataset $S$ when conditioned on $\mathcal{A}_{1:n-1}$. Let $\varepsilon\geq 0$ and $\delta=\delta^{\prime}+\delta^{\prime\prime}$ be the target privacy parameters such that $\delta^{\prime},\delta^{\prime\prime}>0$. Consider the function $N:\mathbb{R}^{\infty}_{\geq 0}\times\mathbb{R}^{\infty}_{\geq 0}\to\mathbb{N}$ such that $N((\varepsilon_{n})_{n\geq 1},(\delta_{n})_{n\geq 1})$ is defined as
$\inf\left\{n:\varepsilon<\sqrt{2\log\left(\frac{1}{\delta^{\prime}}\right)\sum_{m\leq n+1}\varepsilon_{m}^{2}}+\frac{1}{2}\sum_{\sum_{m\leq n+1}}\varepsilon_{m}^{2}\text{ or }\delta^{\prime\prime}<\sum_{m\leq n+1}\delta_{m}\right\}.$
Then, the algorithm $\mathcal{A}_{1:N(S)}$ is $(\varepsilon,\delta)$-DP, where $N(S)=N((\varepsilon_{n}(S))_{n\geq 1},(\delta_{n}(S))_{n\geq 1})$.
The following corollary follows from the theorem above, by noting that $\varepsilon^{\prime}\leq\frac{\varepsilon}{2\sqrt{2T\log(1/\delta)}}$ and $0\leq\varepsilon\leq 8\log(1/\delta)$ imply $\varepsilon\geq\varepsilon^{\prime 2}T/2+\varepsilon^{\prime}\sqrt{2T\log(1/\delta)}$.
Corollary 3.6 (Advanced composition of pure DP mechanisms).
For $\varepsilon^{\prime}>0$. the $T$-fold adaptive composition of $\varepsilon^{\prime}$-DP mechanisms satisfies $(\epsilon,\delta)$-DP for any $\delta\in(0,1)$ as long as $\varepsilon^{\prime}\leq\frac{\varepsilon}{2\sqrt{2T\log(1/\delta)}}$ and $0\leq\varepsilon\leq 8\log(1/\delta)$.
Proof 3.7 (Proof of Proposition 3.3).
We start proving that for all $t\in[T],k\in[K]$, $\tilde{x}^{t+1},\tilde{y}^{t+1}$ and $\hat{x}^{t+1,k},\hat{y}^{t+1,k}$ are private given the previous private iterates. First, notice that $\tilde{x}^{1},\tilde{y}^{1}$ and $\hat{x}^{1,k},\hat{y}^{1,k}$ are trivially private because they don’t depend on the dataset $S$. Now we consider $\tilde{x}^{t+1},\hat{x}^{t+1,k}$. Unrolling the $x$ update, $x^{t+1}_{j}\propto\exp\left\{-\tau\left(\sum_{i=1}^{t}g^{i}_{x,j}\right)\right\}$, where $g^{t}_{x,j}$ denotes the $j$-th coordinate of $g^{t}_{x}$ for every $t\in[T]$. Hence, $\tilde{x}^{t+1}$ and $\hat{x}^{t+1,k}$ are vertices of the simplex $\Delta_{x}$, and each of them is choosen with the exponential mechanism with score function $s^{t}_{x}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{x,j}\right)$. The privacy guarantee of the exponential mechanism in Theorem 3.4 implies that $\tilde{x}^{t+1}$ and $\hat{x}^{t+1,k}$ are $2\Delta(s^{t}_{x})$-DP, where $\Delta(s^{t}_{x})$ is the sensitivity of $s^{t}_{x}$. Using the same idea, we argue that $\tilde{y}^{t+1}$ and $\hat{y}^{t+1,k}$ are all $2\Delta(s^{t}_{y})$-DP, where $\Delta(s^{t}_{y})$ is the sensitivity of $s^{t}_{y}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{y,j}\right)$.
By $L_{0}$-Lipschitzness of $f$ we have that $\max_{j\in[d]}\left|s^{t}_{x}(S,j)-s^{t}_{x}(S^{\prime},j)\right|\leq\frac{2L_{0}\tau}{B}$. Hence, every sampled vertex $\tilde{x}^{t+1},\hat{x}^{t+1,k}$ is $\frac{4L_{0}\tau}{B}$-DP, conditionally on the previously sampled vertices. Analogously, $\tilde{y}^{t+1}$ and $\hat{y}^{t+1,k}$ are $\frac{4L_{0}\tau}{B}$-DP, conditionally the on previously sampled vertices. By Advanced Composition of DP (Corollary 3.6), it suffices that each sampled vertex is $\frac{\varepsilon}{2\sqrt{4T(K+1)\log(1/\delta)}}$-DP in order to make the entire Algorithm 1 $(\varepsilon,\delta)$-DP. Then, for the algorithm to be $(\varepsilon,\delta)$-DP, it suffices that $\frac{4L_{0}\tau}{B}\leq\frac{\varepsilon}{2\sqrt{4T(K+1)\log(1/\delta)}}$.
3.1.2 Convergence Analysis of Algorithm 1
For the accuracy analysis, note that Algorithm 1 produces both a private sequence $(\tilde{x}^{t},\tilde{y}^{t})_{t\in[T]}$ and a non-private one $(x^{t},y^{t})_{t\in[T]}$. We will bound the expected duality gap evaluated at the private iterates’ average. To do this, we first show a gap bound on the non-private iterates’ average, and then we quantify the discrepancy between the duality gap evaluated at the private and non-private iterates’ average.
Lemma 3.8 (Duality Gap at non-private iterates’ average (Implicit in [JNLS:2009])).
Let $(x^{t},y^{t})$ be the (non-private) iterates of Algorithm 1 generated in line 6, and let $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t}$ and $\bar{y}=\frac{1}{T}\sum_{t\in[T]}y^{t}$. Furthermore, let $g^{t}_{x}$ and $g^{t}_{y}$ be the estimators constructed in line 5 of the algorithm and $g^{t}=(g^{t}_{x},g^{t}_{y})$. Denote $\widetilde{\Delta}_{t}=\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}$. Then, for some sequences $(w^{t})_{t\in[T]}\subset\Delta_{x}$ and $(v^{t})_{t\in[T]}\subset\Delta_{y}$ that are predictable w.r.t. the natural filtration, the following inequality holds
$T\operatorname{Gap}(\bar{x},\bar{y})$ $\leq\frac{2\ell}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\widetilde{\Delta}_{t}\|_{\infty}^{2})+\sum_{t=1}^{T}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
For the proof of this lemma, we will use the following auxiliary result. This result is well known in the SSP literature, and follows directly from regret bounds for the (online) mirror-descent method. Similarly to other works, we use this result in order to circumvent uniform convergence to bound the gap function.
Lemma 3.9 (Corollary 2 in [juditsky_VI]).
Let $\xi^{1},\xi^{2},$... be a sequence in $\mathbb{R}^{d}$ and $\mathcal{X}$ be a compact, convex set. Let $\psi:\mathbb{R}^{d}\to\mathbb{R}$ be a $1$-strongly convex w.r.t $\|\cdot\|$ and $D_{\psi}(x,y)=\psi(x)-\psi(y)-\langle\nabla\psi(y),x-y\rangle$. Define $w^{1}\in\mathcal{X}$, $w^{t+1}=\operatorname{argmin}_{x\in\mathcal{X}}D_{\psi}(w^{t},x)+\langle\xi^{t},x\rangle$. Then, for all $w\in\mathcal{X}$ and $T\in\mathbb{N}$: $\sum_{t\in[T]}\langle\xi^{t},w^{t}-w\rangle\leq D_{\psi}(w^{1},w)+\frac{1}{2}\sum_{t\in[T]}\|\xi^{t}\|^{2}_{*}$, where $\|\cdot\|_{*}$ is the dual norm of $\|\cdot\|$.
Proof 3.10 (Proof of Lemma 3.8).
By convexity-concavity of the loss and the convergence rate of Stochastic Mirror Descent in the simplex setup [JNLS:2009, juditsky_VI]:
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]\leq\sum_{t\in[T]}F_{\mathcal{D}}(x^{t},v)-F_{\mathcal{D}}(w,y^{t})$
$\leq\sum_{t\in[T]}\langle\Phi_{\mathcal{D}}(x^{t},y^{t}),(x^{t},y^{t})-(w,v)\rangle\leq\frac{\ell}{\tau}+\tau\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2}+\sum_{t=1}^{T}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w,v)\rangle,$
for all $w\in\Delta_{x},v\in\Delta_{y}$. Lemma 3.9 with $\xi^{t}=\tau(\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x})$, $\psi$ the entropy function, $\|\cdot\|=\|\cdot\|_{1}$ and $w^{1}=(1/d_{x},...,1/d_{x})$ gives
$\sum_{t\in[T]}\langle\tau(\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}),w^{t}-w\rangle\leq\log(d_{x})+\frac{1}{2}\sum_{t\in[T]}\tau^{2}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}.$
Using this result and denoting $\widetilde{\Delta}_{t,x}=\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}$ we can conclude that
$\sum_{t\in[T]}\langle\widetilde{\Delta}_{t,x},x^{t}-w\rangle\leq\sum_{t\in[T]}\langle\widetilde{\Delta}_{t,x},x^{t}-w^{t}\rangle+\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}.$
An analogous bound for the $y$ iterate holds. Hence, for any $w\in\Delta_{x},v\in\Delta_{y}$,
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]$ $\leq\frac{2\ell}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\widetilde{\Delta}_{t}\|_{\infty}^{2})+\sum_{t=1}^{T}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
Maximizing over $w\in\Delta_{x},v\in\Delta_{y}$ finishes the proof.
Now we quantify the discrepancy between the gap of the private (sampled) and non-private iterates. The duality gap is defined in terms of function values, but since it involves a maximum (and the maximum of smooth functions is not necessarily smooth) our sparsification results do not directly apply. We handle this issue with a smoothing argument that leads to the following lemma.
Lemma 3.11.
[…]
Proof 3.12 (Proof of Lemma 3.11).
Let $F_{\lambda}(x):=\max_{y\in\Delta_{y}}[F_{\mathcal{D}}(x,y)+\lambda\phi(y)]$ where $\phi$ is the negative entropy. Since $F_{\mathcal{D}}(x,y)$ is $L_{1}$-smooth in $(x,y)$ and concave in $y$, and $\lambda\phi(y)$ is $\lambda$-strongly convex in $y$, by Proposition 1 in [generalized_fenchel-young_losses], $F_{\lambda}$ is $\big(L_{1}+\frac{L_{1}^{2}}{\lambda}\big)$-smooth w.r.t $\|\cdot\|_{1}$. Hence by Corollary 2.3 applied to $F_{\lambda}$, we have $\mathbb{E}[F_{\lambda}(\tilde{x})-F_{\lambda}(\bar{x})]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}$. In addition, it is easy to see that $\max_{v\in\Delta_{y}}F_{\mathcal{D}}(x,v)\leq F_{\lambda}(x)\leq\lambda\log(d_{y})+\max_{v\in\Delta_{y}}F_{\mathcal{D}}(x,v)$ for all $x\in\Delta_{x}$, since $0\leq\phi(y)\leq\log(d_{y})$. Therefore,
$\mathbb{E}\big[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)\big]=\mathbb{E}\big[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-F_{\lambda}(\tilde{x})\big]+\mathbb{E}[F_{\lambda}(\tilde{x})-F_{\lambda}(\bar{x})]\\ +\mathbb{E}\big[F_{\lambda}(\bar{x})-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)\big]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{y}).$
Analogously, $\mathbb{E}\big[\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\bar{y})-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\tilde{y})\big]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{x})$. Using both inequalities, we get
[…]
$=\mathbb{E}[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)]+\mathbb{E}[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\bar{y})]$
[…]
$\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{y})+\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]+\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{x})$
$=\frac{4L_{1}}{T}+\frac{4L_{1}^{2}}{\lambda T}+\lambda\ell+\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})].$
Choosing $\lambda=L_{1}\frac{2}{\sqrt{T\ell}}$ gives the bound.
With these results, we are ready to prove the convergence guarantees.
Proof 3.13 (Convergence of Algorithm 1).
Following the notation in Algorithm 1, denote $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$ and $(\bar{x},\bar{y})=\frac{1}{T}\sum_{t=1}^{T}(x^{t},y^{t})$. Using Lemma 3.8 and the fact that $\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2}\leq 5L_{0}^{2}$ by $L_{0}$-Lipschitzness of the objective, we obtain $\operatorname{Gap}(\bar{x},\bar{y})\leq\frac{2\ell}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle$. Since $w^{t},v^{t}$ are deterministic when conditioning on the randomness up to iteration $t$, taking expectations on both sides of the previous inequality leads to
(8) $\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]$ $\lesssim\frac{2\ell}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}.$
[…]
Now, we prove Equation (6). Note that $|\mathbb{E}[\nabla_{x,j}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x,j}\mid\mathcal{F}_{t}]|\leq\frac{2L_{2}}{K}$ by Corollary 2.4 and the second-order-smoothness assumption. Similarly, we have $|\mathbb{E}[-\nabla_{y,i}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y,i}\mid\mathcal{F}_{t}]|\leq\frac{2L_{2}}{K}$ for all $i\in[d_{y}]$. We conclude that $\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}\leq\frac{2L_{2}}{K}$. Moreover, by Lemma 3.11, $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq 8L_{1}\sqrt{\ell}/\sqrt{T}$. Hence $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim L_{1}\frac{\sqrt{\ell}}{\sqrt{T}}+\frac{\ell}{\tau T}+\tau L_{0}^{2}+\frac{L_{2}}{K}$. Setting $T\eqsim\min\Big\{n,\frac{(n\varepsilon)^{4/5}}{\ell^{1/5}\log(1/\delta)^{2/5}}\Big\}$, $\tau\eqsim\min\Big\{\frac{1}{L_{0}}\sqrt{\frac{\ell}{T}},\frac{n\varepsilon}{L_{0}T\sqrt{TK\log(1/\delta)}}\Big\}$ and $K\eqsim\sqrt{\frac{T}{\ell}}$ proves (6).
Finally, we prove Equation (7). Since $L_{2}=0$, we can safely set $K=1$. Optimizing over $T,\tau$ gives the bound.
3.2 A constant probability duality gap bound via Private Stochastic Mirror Descent with Bias-Reduced Sampling from Vertices
The limitation that prevents Algorithm 1 from attaining optimal rates is the bias of the stochastic gradient oracles. This is a consequence of sampling from vertices, and can only be mitigated by increasing the sampling rate. However, this eats up the privacy budget. In this section, we propose circumventing this limitation by combining our gradient oracles with a bias-reduction method; see Algorithm 3. We first show how that this procedure leads to improved rates, but only with constant probability. In Section 3.3 we show how to obtain a high probability bound of roughly the same order.
Theorem 3.14 (Guarantees for Algorithm 3).
[…]
$\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\lesssim\log(n)\bigg[\sqrt{C}\sqrt{\frac{\ell}{n}}+\sqrt{L_{0}}C^{1/4}\left(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg]\right]\geq 0.99.$
Up to poly-logarithmic factors in $n,d_{x},d_{y}$, the rate above is indeed optimal. Before proving this result, we will explain Algorithm 3 in more detail.
Bias-Reduced Gradient Oracle
We say a random variable $N$ is distributed as a truncated geometric with parameters $(p,M)$ if $\mathbb{P}[N=k]=p^{k}\mathbbm{1}_{(k\in\{0,1,...,M\})}/C_{M}$, where the normalizing constant is $p\leq C_{M}\leq 1$. We denote it by $\mbox{TG}(p,M)$. To construct a bias-reduced estimator of the gradient at $(x,y)$, we start by sampling $N\sim\mbox{TG}(0.5,M)$. Then, we construct a batch $B$ of $\max\{1,2^{N^{t}}/\alpha\}$ fresh samples, where $\alpha$ is a parameter to be choosen later. This batch $B$ combined with the gradient estimates (see Algorithm 2 and note similar algorithms in [Blanchet:2015]) reduces the bias while preserving privacy and accuracy (see Thm. 3.14). To do so, we control the bias and second moment of the gradient estimator.
Algorithm 2 BiasReducedGradient($x,y,N,B$)
 Sample $(\hat{x}^{i},\hat{y}^{i})_{i\in[2^{N+1}]}\overset{\text{iid}}{\sim}P_{x}\times P_{y}$
$(\bar{x}_{+},\bar{y}_{+})=\frac{1}{2^{N+1}}\sum_{i\in[2^{N+1}]}(\hat{x}^{i},\hat{y}^{i})$ ; $(\bar{x}_{-},\bar{y}_{-})=\frac{1}{2^{N}}\sum_{i\in[2^{N}]}(\hat{x}^{i},\hat{y}^{i})$
$g_{x}=C_{M}2^{N}(\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B))+\nabla_{x}f(\hat{x}^{1},\hat{y}^{1};B)$
$g_{y}=-C_{M}2^{N}(\nabla_{y}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{y}f(\bar{x}_{-},\bar{y}_{-};B))-\nabla_{y}f(\hat{x}^{1},\hat{y}^{1};B)$
return $g_{x},g_{y}$
Lemma 3.15 (Bias and Second Moment of Algorithm 2).
Suppose the objective of the SSP problem in the $\ell_{1}$ setting is second-order-smooth. Let $x\in\Delta_{x},y\in\Delta_{y}$, $N\sim\mbox{TG}(0.5,M)$ and $B\sim\mathcal{D}^{N}$. Let $(g_{x},g_{y})=\text{BiasReducedGradient}(x,y,N,B)$. Then,
$\|\mathbb{E}[\Phi_{\mathcal{D}}(x,y)-(g_{x},g_{y})]\|_{\infty}$ $\leq\frac{2L_{2}}{2^{M}},$
$\text{and}\quad\mathbb{E}[\|(g_{x},g_{y})\|_{\infty}^{2}]$ $\lesssim L_{0}^{2}+L_{2}^{2}+L_{1}^{2}M[\log(d_{x})+\log(d_{y})].$
Consequently, the bias scales inversely proportional to $2^{M}$, where $M$ is largest possible realization of $N$. Hence the expected bias is roughly same as if we were sampling $2^{M}$ times while we actually sample only $2^{N}$ times. This leads to a more favorable privacy-utility tradeoff. Another consequence is the order of the second moment, $M\ell$. Here $M$ will be choosen as a logarithmic quantity (i.e., $M\lesssim\log(\min\{n\varepsilon,n\})$). Hence, the method allows us to improve the bias at a only a logarithmic cost in the second moment bound. Note, however, a disadvantage of this gradient oracle is we can only control its second moment for second-order-smooth objective functions.
Proof 3.16 (Proof of Lemma 3.15).
Following the notation in Algorithm 2, we have
$\mathbb{E}[g_{x}]=\textstyle\mathbb{E}\left[\sum_{k=0}^{M}\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)+\frac{2^{-k}}{C_{M}}\nabla_{x}f(\hat{x}^{1},\hat{y}^{1};B)\mid N=k\right]$
$=\textstyle\mathbb{E}\left[\sum_{k=0}^{M}\big(\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})-\nabla_{x}F_{\mathcal{D}}(\bar{x}_{-},\bar{y}_{-})\big)+\frac{2^{-k}}{C_{M}}\nabla_{x}F_{\mathcal{D}}(\hat{x}^{1},\hat{y}^{1})\mid N=k\right]$
$=\textstyle\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})\mid N=M\right]-\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(x_{-},y_{-})\mid N=0\right]+\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\hat{x}^{1},\hat{y}^{1})\right]$
$=\textstyle\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})\mid N=M\right],$
where the first equality follows from the law of total expectation, the second from the fact that $(\bar{x}_{+},\bar{y}_{+})$$(\bar{x}_{-},\bar{y}_{-})$, $(\hat{x}^{1},\hat{y}^{1})$ and $B$ are conditionally independent given $N$, the third from $\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})\mid N=k\right]=\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{-},\bar{y}_{-})\mid N=k+1\right]$ and the last one from $\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(x_{-},y_{-})\mid N=0\right]=\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\hat{x}^{1},\hat{y}^{1})\right]$. Finally, we get $\|\mathbb{E}[g_{x}]-\nabla_{x}F_{\mathcal{D}}(x,y)\|_{\infty}=\max_{j\in[d]}|\mathbb{E}[g_{x,j}]-\nabla_{x,j}F_{\mathcal{D}}(x,y)|\leq\frac{2L_{2}}{2^{M}}$ by Corollary 2.3. Clearly, an analogous bound holds for the gradient with respect to $y$. Writing $\|\mathbb{E}[(g_{x},g_{y})]-\Phi_{\mathcal{D}}(x,y)\|_{\infty}=\max\{\|\mathbb{E}[g_{x}]-\nabla_{x}F_{\mathcal{D}}(x,y)\|_{\infty},\|\mathbb{E}[g_{y}]-\nabla_{y}F_{\mathcal{D}}(x,y)\|_{\infty}\}$ gives the bias bound.
We now proceed with the second moment bound. Then, $\mathbb{E}\|g_{x}\|_{\infty}^{2}$ is equal to
$\textstyle\sum_{k=0}^{M}\mathbb{E}\|C_{M}2^{k}[\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B_{k})-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)]+\nabla_{x}f(\hat{x}^{1},\hat{y}^{1};B)\mid N=k\|_{\infty}^{2}\frac{2^{-k}}{C_{M}}$
$\textstyle\leq 2\left[\sum_{k=0}^{M}\!C_{M}2^{k}\mathbb{E}\|\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)\!-\!\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)\mid N=k\|_{\infty}^{2}\!+\!\frac{2^{-k}}{C_{M}}L_{0}^{2}\right]$
$\textstyle\leq 2L_{0}^{2}+4C_{M}\sum_{k=0}^{M}2^{k}\mathbb{E}\|\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(x,y;B)\mid N=k\|_{\infty}^{2}$
$\textstyle\quad+4C_{M}\sum_{k=0}^{M}\mathbb{E}2^{k}\|\nabla_{x}f(x,y;B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)\mid N=k\|_{\infty}^{2}.$
By Corollary 2.12, $\mathbb{E}\|\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(x,y;B)\mid N=k\|_{\infty}^{2}\lesssim\frac{L_{2}^{2}}{2^{2(k+1)}}+\frac{\log(d_{x})L_{1}^{2}}{2^{k+1}}$, and similarly $\mathbb{E}\|\nabla_{x}f(x,y;B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)\mid N=k\|_{\infty}^{2}\lesssim\frac{L_{2}^{2}}{2^{2k}}+\frac{\log(d_{x})L_{1}^{2}}{2^{k}}$. Hence, $\mathbb{E}\|g_{x}\|_{\infty}^{2}\lesssim L_{0}^{2}+L_{2}^{2}+M\log(d_{x})L_{1}^{2}$. Analogously, $\mathbb{E}\|g_{y}\|_{\infty}^{2}\lesssim L_{0}^{2}+L_{2}^{2}+M\log(d_{y})L_{1}^{2}$.
Bias-reduced Algorithm
We propose Algorithm 3, which is essentially Algorithm 1 combined with the bias-reduced gradient oracle. The main challenges encountered are the oracles becoming heavy-tailed (moments are exponentially larger than their expectation), and, due to a randomization of batch sizes, it is difficult to control the privacy budget. To handle the heavy-tailed nature of our estimators, we resort to efficiency estimates from Markov’s inequality which only certify success with constant probability. For handling the privacy budget challenge, we use the Fully Adaptive Composition Theorem of DP [Whitehouse:2023], which allows us to set up the mechanisms and their privacy parameters in a random (but predictable) way.
Algorithm 3 Private stochastic mirror descent with bias-reduced vertex sampling
0:Dataset $S$, Step-size $\tau$, Stopping parameter $U\geq 1$, TG parameter $M$, Batch parameter $\alpha$.
$x^{1}=(1/d_{x},...,1/d_{x})$, $y^{1}=(1/d_{y},...,1/d_{y}),t=1,N^{1}\sim\mbox{TG}(0.5,M)$
while $\sum_{i\in[t-1]}2^{N^{i}}\leq U-2^{M}$ do
  Let $B^{t}$ be a fresh batch of $\max\{1,2^{N^{t}}/\alpha\}$ samples.
  $(g^{t}_{x},g^{t}_{y})=\operatorname{BiasReducedGradient}(x^{t},y^{t},N^{t},B^{t})$
  $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau g^{t}_{x,j}\right),\quad\forall j\in[d_{x}]$  and  $y^{t+1}_{i}\propto y^{t}_{i}\exp\left(\tau g^{t}_{y,i}\right),\quad\forall i\in[d_{y}]$
  $N^{t+1}\sim\mbox{TG}(0.5,M)$  and set $t=t+1$
end while
return $(\tilde{x},\tilde{y})=\frac{1}{t}\sum_{i=1}^{t}(\tilde{x}^{i},\tilde{y}^{i})$ where $(\tilde{x}^{i},\tilde{y}^{i})_{i\in[t]}\stackrel{{\scriptstyle iid}}{{\sim}}P_{x^{i}}\times P_{y^{i}}$
Next, we will prove Theorem 3.14. It will be useful to define the stopping time $\mathcal{T}=\sup\big\{T\in\mathbb{N}:\sum_{t=1}^{T}2^{N_{t}}\leq U-2^{M}\big\}.$ We note that Algorithm 3 performs $\mathcal{T}+1$ optimization steps (see condition in Line 2). The definition of $\mathcal{T}$ ensures that $\sum_{t=1}^{\mathcal{T}}2^{N_{t}}\leq U-2^{M}$, and, since $2^{N_{t}}\leq 2^{M}$, we have $\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq U$. We will first prove the privacy guarantees and then convergence.
3.2.1 Privacy Analysis of Algorithm 3
Since this algorithm runs with random batch sizes and for a random number of steps, we use Fully Adaptive Composition (Theorem 3.5) to provide a sufficient condition on $U$ that ensures privacy.
Proposition 3.17.
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. If $U\leq\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}}$, then Algorithm 3 is $(\varepsilon,\delta)$-DP.
Proof 3.18.
By unraveling the recursion that defines $x^{t+1}$ iterate, it is easy to see that its $j$-th coordinate, $x^{t+1}_{j}$, is proportional to
$\exp\left\{-\tau\sum_{r=1}^{t}\Big[2^{N_{r}}\Big(\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big)+\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{r})\Big]\right\}.$
Hence sampling a vertex from $\Delta_{x}$ where the $j$-th vertex is choosen with probability $x^{t}_{j}$ corresponds to an application of the exponential mechanism with score function given by $s^{t}_{x}=-\tau\sum_{r=1}^{t}\Big[2^{N_{r}}\Big(\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big)+\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{r})\Big].$ Next, note that in iteration $t$ of Algorithm 3, we sample a vertex $\tilde{x}^{t}\sim P_{x^{t}}$ in Line 8 and we also sample $2^{N_{t}+1}$ times from $P_{x^{t}}$ to construct the bias-reduced gradient (with respect to $x$) estimate in Line 1 of Algorithm 2. Analogously, sampling vertices from $\Delta_{y}$ corresponds to the exponential mechanism with score function $s^{t}_{y}$, which is defined similarly to $s^{t}_{x}$ but using gradients w.r.t  the $y$ variable. Denote by $\Delta(s_{x}^{t})$ and $\Delta(s_{y}^{t})$ the sensitivities of the score functions $s_{x}^{t}$ and $s_{y}^{t}$, respectively. Recall that exponential mechanism with an score function that has sensitivity $\Delta$ is $2\Delta$-DP. Hence, based on the fully adaptive composition of DP (Theorem 3.5), the algorithm is $(\varepsilon,\delta)$-DP if $\sqrt{8\log(1/\delta)S}+2S\leq\varepsilon\text{ where }S=\sum_{t=1}^{\mathcal{T}+1}(2^{N_{t}+1}+1)(\Delta(s_{x}^{t})^{2}+\Delta(s_{y}^{t})^{2})$. Next, we bound $\Delta(s_{x}^{t})$. Fix a pair of neighboring datasets $S$ and $S^{\prime}$. Denote by $s_{x}^{t}$ the score function with $S$ and by $s_{x}^{\prime t}$ with $S^{\prime}$. When conditioning on the randomness up to iteration $t$ (recall the batch-size is a random quantity), either $B^{r}=B^{\prime r}$ for all $r\in[t]$, in which case $s_{x}^{t}=s_{x}^{\prime t}$, or there is exactly one index $r$ such that $B^{r}$ differ from $B^{\prime r}$ in exactly one datapoint (say $z^{*}$ and $z^{\prime*}$). In this case,
$|s_{x}^{t}-s_{x}^{\prime t}|=\tau\Big|2^{N_{r}}\Big(\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{r})\Big)+\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{r})$
$\qquad-2^{N_{r}}\Big(\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};B^{\prime r})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};B^{\prime r})\Big)-\nabla_{x}f(x_{0}^{r},y_{0}^{r};B^{\prime r})\Big|$
$\leq\tau\frac{2^{N_{r}}\alpha}{2^{N_{r}}}\Big|\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};z^{*})-\nabla_{x}f(\bar{x}_{+}^{r},\bar{y}_{+}^{r};z^{\prime*})+\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};z^{*})-\nabla_{x}f(\bar{x}_{-}^{r},\bar{y}_{-}^{r};z^{\prime*})\Big|$
$\qquad+\frac{\tau\alpha}{2^{N_{r}}}\Big|\nabla_{x}f(x_{0}^{r},y_{0}^{r};z^{*})-\nabla_{x}f(x_{0}^{r},y_{0}^{r};z^{\prime*})\Big|\leq\tau\alpha L_{0}(4+1/2^{N_{r}})\leq 4.5\tau\alpha L_{0}.$
Analogously, $\Delta(s_{y}^{t})\leq 4.5\tau\alpha L_{0}$. Hence, every sampled vertex (either from $\Delta_{x}$ or $\Delta_{y}$), is $9\tau\alpha L_{0}$-DP. Plugging in the condition for privacy mentioned above, we need $\sqrt{2\log(1/\delta)\sum_{t=1}^{\mathcal{T}+1}(4\cdot 2^{N_{t}}+2)(9\tau\alpha L_{0})^{2}}+\frac{1}{2}\sum_{t=1}^{\mathcal{T}+1}(4\cdot 2^{N_{t}}+2)(9\tau\alpha L_{0})^{2}\leq\varepsilon$. We further restrict this stopping time by upper bounding each of the above terms by $\varepsilon/2$. Further noticing that such a bound for the first term implies the same bound for the second term so long as $\varepsilon\leq 8\log(1/\delta)$, we conclude that it suffices
$4\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}+2(\mathcal{T}+1)\leq 6\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq\frac{\varepsilon^{2}}{8\log(1/\delta)(9\tau\alpha L_{0})^{2}}.$
This condition is satisfied when $U\leq\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}}$, since
$4\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}+2(\mathcal{T}+1)$ $\leq 6\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq 6\sum_{t=1}^{\mathcal{T}}2^{N_{t}}+6\cdot 2^{M}\leq 6(U-2^{M})+6\cdot 2^{M}=6U.$
3.2.2 Convergence Analysis of Algorithm 3
Proof 3.19 (Proof of Theorem 3.14).
For privacy we need $U\leq\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}}$ (Proposition 3.17). At the same time, we need $\sum_{t\in[\mathcal{T}+1]}\max\{1,2^{N_{t}}/\alpha\}\leq n$, to ensure Algorithm 3 does not exceed the sample size. For this condition to hold, it suffices that $\mathcal{T}+1\leq n/2$ and $\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq n\alpha/2$. This way, we only need $\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq\min\{n,n\alpha/2\}$. Since $\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq U$, then for this last condition to hold, it is sufficient that $U\leq\min\{n,n\alpha/2\}$. It is easy to see that under our parameter setting it holds that $U=\min\left\{\frac{\varepsilon^{2}}{48\log(1/\delta)(9\tau\alpha L_{0})^{2}},n\alpha/2,n/2\right\}$, so we don’t exceed our privacy budget and we satisfy the sample size constraint.
[…]
(9) $\mathbb{P}\Big[\operatorname{Gap}(\tilde{x},\tilde{y})\geq\frac{A}{B}\Big]\leq\mathbb{P}[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\geq A]+\mathbb{P}\Big[\frac{1}{(\mathcal{T}+1)}\geq\frac{1}{B}\Big].$
Note the second probability in the right hand side can be bounded as follows:
$\mathbb{P}\Big[\frac{1}{\mathcal{T}+1}\geq\frac{1}{B}\Big]=\mathbb{P}[\mathcal{T}\leq B-1]\leq\mathbb{P}\Big[{\textstyle\sum_{t\in[B]}}2^{N_{t}}>U-2^{M}\Big]\leq\frac{B\mathbb{E}[2^{N_{1}}]}{U-2^{M}},$
where the first inequality holds since $\mathcal{T}\leq B-1$ implies that $\sum_{t\in[B]}2^{N_{i}}>U$, and the second follows from Markov’s inequality. To bound the first probability on the right hand side of (9), we use Markov’s inequality again
$\mathbb{P}\left[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\geq A\right]\leq\frac{\mathbb{E}\left[(\mathcal{T}+1)(\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y}))+(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})\right]}{A}.$
Let us bound $\mathbb{E}\left[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})\right]$. From Lemma 3.8, we have
$(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})\lesssim\frac{\ell}{\tau}+\frac{\tau}{2}\sum_{t=1}^{\mathcal{T}+1}(L_{0}^{2}+\|g^{t}_{x}\|_{\infty}^{2}+\|g^{t}_{y}\|_{\infty}^{2})+\sum_{t=1}^{\mathcal{T}+1}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle,$
where $\widetilde{\Delta}_{t}=\Phi_{\mathcal{D}}(x^{t},y^{t})-(g_{x}^{t},g_{y}^{t})$. Since $\mathcal{T}+1\leq\sum_{t\in[\mathcal{T}+1]}2^{N_{t}}\leq U$, then
$\textstyle\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\|g^{t}_{x}\|_{\infty}^{2}\right]$ $=\textstyle\mathbb{E}\left[\sum_{t=1}^{U}\mathbb{E}[\|g^{t}_{x}\|_{\infty}^{2}\mathbbm{1}_{(\mathcal{T}\geq t-1)}\mid\mathcal{F}_{t}]\right]=\mathbb{E}\left[\sum_{t=1}^{U}\mathbb{E}[\|g^{t}_{x}\|_{\infty}^{2}]\mathbbm{1}_{(\mathcal{T}\geq t-1)}\right]$
$\lesssim\textstyle\left(L_{0}^{2}+L_{2}^{2}+\log(d_{x})ML_{1}^{2}\right)\mathbb{E}\left[\sum_{t=1}^{U}\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]$
$\lesssim\left(L_{0}^{2}+L_{2}^{2}+\log(d_{x})ML_{1}^{2}\right)\mathbb{E}[\mathcal{T}+1],$
where $\mathcal{F}_{t}$ is the sigma algebra generated by all the randomness in the algorithm that generated $x^{t}$. In the second and third equalities, we used that $\{\mathcal{T}\geq t-1\}=\big\{\sum_{k\in[t-1]}2^{N_{k}}\leq U-2^{M}\big\}$ is $\mathcal{F}_{t}$-measurable (since $2^{N_{1}},...,2^{N_{t-1}}$ are used to generate $x^{t}$), and hence it can be taken out of the conditional expectation. Next, for the first inequality we used that $2^{N_{t}}$ is independent of $\mathcal{F}_{t}$, so we can use Lemma 3.15 to bound the conditional expectation of the second moment. In the last line we used the fact that the expectation can be written as the sum of the tail probabilities. Analogously, we can obtain $\mathbb{E}\big[\sum_{t=1}^{\mathcal{T}+1}\|g^{t}_{y}\|_{\infty}^{2}\big]\lesssim L_{0}^{2}+L_{2}^{2}+\log(d_{y})ML_{1}^{2}\mathbb{E}[\mathcal{T}+1]$. In a similar way, we can use the same technique of introducing the indicator of $\mathcal{T}+1\geq t$ and then Lemma 3.15 to bound the bias term: $\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle\right]\lesssim\frac{L_{2}\mathbb{E}[\mathcal{T}+1]}{2^{M}}$. Now, we take expectations on both sides of the expression given by Lemma 3.8 and then use the bounds we established above for the second moment and bias of the bias-reduced gradient estimators:
$\mathbb{E}[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})]\lesssim\frac{\ell}{\tau}+\tau[L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}]\mathbb{E}[\mathcal{T}+1]+\frac{L_{2}\mathbb{E}[\mathcal{T}+1]}{2^{M}}.$
To bound $\mathbb{E}[\mathcal{T}+1]$ we note that $\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\leq U$, so using Wald’s identity [Wald:1944] yields
$U$ $\geq\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}2^{N_{t}}\right]=\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\mathbb{E}[2^{N_{t}}\mathbbm{1}_{(\mathcal{T}+1\geq t)}|N_{t-1},...,N_{1}]\right]=\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\mathbb{E}[2^{N_{t}}]\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]$
$=\textstyle\mathbb{E}[2^{N_{1}}]\mathbb{E}\left[\sum_{t=1}^{\mathcal{T}+1}\mathbbm{1}_{(\mathcal{T}+1\geq t)}\right]=\mathbb{E}[2^{N_{1}}]\mathbb{E}\left[\mathcal{T}+1\right]\eqsim M\mathbb{E}\left[\mathcal{T}+1\right].$
Hence, $\mathbb{E}\left[\mathcal{T}+1\right]\lesssim U/\mathbb{E}[2^{N_{1}}]\eqsim U/M$. It follows that
$\mathbb{E}[(\mathcal{T}+1)\operatorname{Gap}(\bar{x},\bar{y})]\lesssim\frac{\ell}{\tau}+\tau[L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}]\frac{U}{M}+\frac{L_{2}U}{M2^{M}}\\ \lesssim\frac{\ell}{\tau}+\tau[L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}]U+\frac{L_{2}U}{2^{M}}\lesssim\sqrt{\ell[L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}]U},$
where the last step follows from $M=\log_{2}(\sqrt{U})$ and $\tau=\sqrt{\frac{\ell}{(L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2})U}}$.
Now, let us bound $\mathbb{E}\left[(\mathcal{T}+1)(\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y}))\right]$. We can assume that $\tilde{x}$’s and $\tilde{y}$’s are sampled after the While Loop is over in the Algorithm, thus by Lemma 3.11
$\mathbb{E}\left[(\mathcal{T}+1)(\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y}))\right]=\mathbb{E}\left[(\mathcal{T}+1)\mathbb{E}_{\tilde{x},\tilde{y}}\left[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})\mid\mathcal{T}\right]\right]\\ \textstyle\leq\mathbb{E}\left[(\mathcal{T}+1)\left(\frac{4L_{1}}{\mathcal{T}+1}+\frac{2L_{1}\sqrt{\ell}}{\sqrt{\mathcal{T}+1}}\right)\right]=4L_{1}+2L_{1}\sqrt{\ell}\mathbb{E}\left[\sqrt{\mathcal{T}+1}\right]\lesssim\sqrt{\ell L_{1}^{2}\frac{U}{M}},$
where the second step follows by Jensen’s inequality. Combining the bounds obtained, we conclude that $\mathbb{P}\left[(\mathcal{T}+1)\operatorname{Gap}(\tilde{x},\tilde{y})\geq A\right]\lesssim\frac{\sqrt{\ell[L_{0}^{2}+L_{2}^{2}+(\ell)L_{1}^{2}]U}}{A}$.
Plugging all the bounds in Equation 9, $\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\geq\frac{A}{B}\right]\lesssim\frac{\sqrt{\ell[L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}]U}}{A}+\frac{BM}{U-\sqrt{U}}$. Since $U\geq 4$, then $U-\sqrt{U}\geq U/2$. Choosing $A\eqsim\sqrt{\ell[L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}]U}$, $B\eqsim U/M$ and using $M\lesssim\log(n)$, gives the claimed bound.
3.3 A high probability duality gap bound via Private Stochastic Mirror Descent with Bias-Reduced Sampling from Vertices
[…]
Theorem 3.20.
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. Suppose we are solving the DPSSP problem in the $\ell_{1}$ setting where the objective $f(\cdot,\cdot;z):\Delta_{x}\times\Delta_{y}\to[-B,B]$ is convex-concave and second-order-smooth for all $z\in\mathcal{Z}$. Let $(x_{i^{*}},y_{i^{*}})$ be the output of Algorithm 4. Then, $(x_{i^{*}},y_{i^{*}})$ is $(\varepsilon,\delta)$-DP and for any $\beta>0$, with probability at least $1-\beta$, $\operatorname{Gap}(x_{i^{*}},y_{i^{*}})$ is bounded above (up to $\operatorname{polylog}(n,1/\beta)$ and constant factors) by $\sqrt{\frac{\ell}{n}}+\frac{\ell}{\sqrt{n\varepsilon}}+\frac{B}{\sqrt{n}}+\frac{B}{n\varepsilon}$.
Algorithm 4 Boosted private stochastic mirror descent with bias-reduced vertex sampling
1:Input: Dataset $S$, parameters $I,J$, , privacy parameters $\varepsilon,\delta$
2: Let $\mathcal{A}^{DPSSP}$ be Algorithm 3 and $\mathcal{A}^{DPSCO}$ be Algorithm 3 from [asi2021private].
3: Partition $S$ into 4 equal parts $S_{1},S_{2},S_{3},S_{4}$
4: Partition $S_{1}$ into II equal parts: $S_{1}^{1},S_{1}^{2},\dots,S_{1}^{I}$
5: For each $i\in[I]$, compute $(x_{i},y_{i})=\mathcal{A}^{DPSSP}(S_{1}^{i},\varepsilon,\delta)$
6: Partition $S_{2}$ and $S_{3}$ into IJIJ equal parts: $S_{2}^{11},S_{2}^{12},\dots,S_{2}^{IJ};S_{3}^{11},S_{3}^{12},\dots,S_{3}^{IJ}$
7: For each $i\in[I],j\in[J]$, compute $x_{ij}=\mathcal{A}^{DPSCO}(S_{2}^{ij},\varepsilon,\delta,F_{\mathcal{D}}(\cdot,y_{i}))$
8: For each $i\in[I],j\in[J]$, compute $y_{ij}=\mathcal{A}^{DPSCO}(S_{3}^{ij},\varepsilon,\delta,-F_{\mathcal{D}}(x_{i},\cdot)))$
9: For each $i\in[I]$, define $G_{i}(S_{4})=\max_{j\in[J]}F_{S_{4}}(x_{i},y_{ij})+\max_{j\in[J]}-F_{S_{4}}(x_{ij},y_{i})$
10: Sample $i^{*}$ such that $\forall i\in[I]$, $\mathbb{P}[i^{*}=i]\propto\exp\left(-n\varepsilon G_{i}(S_{4})/[32B]\right)$
11:Return: $(x_{i^{*}},y_{i^{*}})$
We first prove the privacy guarantee.
Proof 3.21 (Privacy of Algorithm 4).
$\mathcal{A}^{DPSSP}$ and $\mathcal{A}^{DPSCO}$ are $(\varepsilon,\delta)$-DP algorithms. Since we run them on disjoint parts of $S$, steps 5 and 6 are $(\varepsilon,\delta)$-DP by parallel composition of DP. The last step is an exponential mechanism with score function $G_{i}(S_{4})$. Conditioned on $(x_{i},y_{i}),x_{ij},y_{ij}$, $i\in[I],j\in[J]$, the sensitivity of $G_{i}(S_{4})$ is $4B/|S_{4}|$. Since $|S_{4}|=n/4$, we get that the sensitivity is $16B/n$. Theorem 3.4 gives that step 10 is $(\varepsilon,0)$-DP. We conclude using parallel composition one more time that the whole algorithm is $(\varepsilon,\delta)$-DP.
To prove the accuracy guarantee we will need two results.
Theorem 3.22 (Theorem 7 in [asi2021private]).
Consider DPSCO in the $\ell_{1}$ setting as in Section 4. Then, Private Variance Reduced Stochastic Frank-Wolfe (Algorithm 3 in [asi2021private]) returns $\tilde{x}$ such that $\mathbb{E}[F_{\mathcal{D}}(\tilde{x})-\min_{x\in\Delta_{x}}F_{\mathcal{D}}(x)]\lesssim C(d_{x},n)$ where
(10) $C(d,n):=(L_{0}+L_{1})\frac{\sqrt{\log(d)}\log(n)}{\sqrt{n}}+\left(\frac{\sqrt{L_{1}}L_{0}\log(1/\delta)\log(d)\log(n)^{2}}{n\varepsilon}\right)^{2/3}.$
Remark 3.23.
Combining the above with Markov’s inequality gives that if we run Private Variance Reduced Frank-Wolfe $k$ times on datasets $S_{1},...,S_{k}\overset{iid}{\sim}\mathcal{D}^{n}$, and get outputs $\tilde{x}_{1},...,\tilde{x}_{k}$, then with probability at least $1-2^{k}$, $\min_{i\in[k]}\big(F_{\mathcal{D}}(\tilde{x}_{i})-\min_{x\in\Delta_{x}}F_{\mathcal{D}}(x)\big)\lesssim 2C(d_{x},n)$.
Theorem 3.24 (Theorem 3.11 in [dwork2014DPfoundations]).
Consider the same setting as in Theorem 3.4. Let $j^{*}$ be the object sampled from $J$ according to $\mathbb{P}(j=e_{i})\propto\exp\left(\frac{\varepsilon s(S,i)}{2\Delta_{s}}\right)$. Then, for any $\alpha>0$: $\mathbb{P}\left[s(S,j^{*})\leq\max_{j\in[J]}s(S,j)-\frac{2\Delta_{s}(\log(J)+\log(1/\alpha))}{\varepsilon}\right]\leq\alpha$.
Proof 3.25 (Accuracy of Algorithm 4).
Given constants $C_{1},C_{2},C_{3},C_{4},C_{5}$, define the following events using the notation from the Algorithm 4
$A_{i}^{x}=\left\{\max_{x\in\Delta_{x}}-F_{\mathcal{D}}(x,y_{i})\lesssim\max_{j\in[J]}-F_{\mathcal{D}}(x_{ij},y_{i})+C_{1}\right\},$
$A_{i}^{y}=\left\{\max_{y\in\Delta_{y}}F_{\mathcal{D}}(x_{i},y)\lesssim\max_{j\in[J]}F_{\mathcal{D}}(x_{i},y_{ij})+C_{2}\right\},$
$A_{ij}^{x}=\left\{|F_{\mathcal{D}}(x_{ij},y_{i})-F_{S_{4}}(x_{ij},y_{i})|\lesssim C_{3}\right\},A_{ij}^{y}=\left\{|F_{\mathcal{D}}(x_{ij},y_{i})-F_{S_{4}}(x_{ij},y_{i})|\lesssim C_{3}\right\},$
$A_{exp}=\left\{G_{i^{*}}(S_{4})\lesssim\min_{i\in[I]}G_{i}(S_{4})+C_{4}\right\},A_{BR}=\left\{\min_{i\in[I]}\operatorname{Gap}(x_{i},y_{i})\lesssim C_{5}\right\}.$
By Theorem 3.22, picking $C_{1}=C(d_{x},n/[4IJ]),C_{2}=C(d_{y},n/[4IJ])$, with $C(d,n)$ as defined in Equation (10), gives $\mathbb{P}[A_{i}^{x}]\geq 1-2^{-J},\mathbb{P}[A_{i}^{y}]\geq 1-2^{-J}$. Furthermore, since $|f(x,y;z)|\leq B$, Hoeffding’s inequality gives $\mathbb{P}[A_{ij}^{x}]\geq 1-\alpha,\mathbb{P}[A_{ij}^{y}]\geq 1-\alpha$ for $C_{3}\eqsim B\sqrt{\frac{\log(1/\alpha)}{n}}$. By Theorem 3.24, if we pick $C_{4}\eqsim\frac{B(\log(I)+\log(1/\gamma))}{n\varepsilon}$, then $\mathbb{P}[A_{exp}]\geq 1-\gamma$. Finally, by Theorem 3.14, letting $C=L_{0}^{2}+L_{2}^{2}+\ell ML_{1}^{2}$ we have that for $C_{5}\eqsim\log(n)\bigg[\sqrt{C}\sqrt{\frac{I\ell}{n}}+\sqrt{L_{0}}C^{1/4}\left(\frac{I\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg]$, $\mathbb{P}(A_{BR})\geq 1-2^{-I}$. Let $E$ be the intersection of all the events above. Then, with the proposed choices for the constants $C_{1},...,C_{5}$ it holds
$\mathbb{P}[E]\geq 1-(2^{-I}+\gamma+2\alpha IJ+2I2^{-J}),$
and furthermore, choosing
$I=\log_{2}(4/\beta),\gamma=\beta/4,J=\log_{2}(8I/\beta),\alpha=\beta/(8IJ)$
gives that $\mathbb{P}[E]\geq 1-\beta$ for any $\beta\in(0,1)$. Finally, note that under $E$
$\operatorname{Gap}(x_{i^{*}},y_{i^{*}})$ $=\max_{y\in\Delta_{y}}F_{\mathcal{D}}(x_{i^{*}},y)+\max_{x\in\Delta_{x}}-F_{\mathcal{D}}(x,y_{i^{*}})$
$\leq\max_{j\in[J]}F_{\mathcal{D}}(x_{i^{*}},y_{i^{*}j})+\max_{j\in[J]}-F_{\mathcal{D}}(x_{i^{*}j},y_{i^{*}})+C_{1}+C_{2}$
$\leq\max_{j\in[J]}F_{S_{4}}(x_{i^{*}},y_{i^{*}j})+\max_{j\in[J]}-F_{S_{4}}(x_{i^{*}j},y_{i^{*}})+C_{1}+C_{2}+2C_{3}$
$=G_{i^{*}}(S_{4})+C_{1}+C_{2}+2C_{3}$
$\leq\min_{i\in[I]}G_{i}(S_{4})+C_{1}+C_{2}+2C_{3}+C_{4}$
$=\min_{i\in[I]}\left\{\max_{j\in[J]}F_{S_{4}}(x_{i},y_{ij})+\max_{j\in[J]}-F_{S_{4}}(x_{ij},y_{i})\right\}+C_{1}+C_{2}+2C_{3}+C_{4}$
$\leq\min_{i\in[I]}\left\{\max_{j\in[J]}F_{\mathcal{D}}(x_{i},y_{ij})+\max_{j\in[J]}-F_{\mathcal{D}}(x_{ij},y_{i})\right\}+C_{1}+C_{2}+4C_{3}+C_{4}$
$\leq\min_{i\in[I]}\left\{\max_{y\in\Delta_{y}}F_{\mathcal{D}}(x_{i},y)+\max_{x\in\Delta_{x}}-F_{\mathcal{D}}(x,y_{i})\right\}+C_{1}+C_{2}+4C_{3}+C_{4}$
$=\min_{i\in[I]}\operatorname{Gap}(x_{i},y_{i})+C_{1}+C_{2}+4C_{3}+C_{4}$
$\leq C_{1}+C_{2}+4C_{3}+C_{4}+C_{5}.$
Using the definition of $E$ and our choices for $C_{1},C_{2},C_{3},C_{4},C_{5}$, we conclude that with probability at least $1-\beta$,
$\operatorname{Gap}(x_{i^{*}},y_{i^{*}})\lesssim\frac{(L_{0}+L_{1})\sqrt{\ell\log_{2}(8\log_{2}(4/\beta)/\beta)}\log(n)}{\sqrt{n}}$
$+\left(\frac{\sqrt{L_{1}}L_{0}\log(1/\delta)\ell\log(n)^{2}\log_{2}(8\log_{2}(4/\beta)/\beta)}{n\varepsilon}\right)^{2/3}$
$+B\sqrt{\frac{\log([8\log_{2}(4/\beta)\log_{2}(8\log_{2}(4/\beta)/\beta))]/\beta)}{n}}+\frac{B(\log(\log_{2}(4/\beta))+\log(4/\beta))}{n\varepsilon}$
$+\log(n)\bigg[\sqrt{C}\sqrt{\frac{\log_{2}(4/\beta)\ell}{n}}+\sqrt{L_{0}}C^{1/4}\left(\frac{\log_{2}(4/\beta)\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg].$
This bound is of order $\sqrt{\frac{\ell}{n}}+\frac{\ell}{\sqrt{n\varepsilon}}+\frac{B}{\sqrt{n}}+\frac{B}{n\varepsilon}$ up to $\operatorname{polylog}(n,1/\beta)$ factors.
Algorithm 5 Private stochastic mirror descent with vertex-sampling for DPSCO
0: Dataset $S=\{z^{1},...,z^{n}\}$, Total number of steps $T$, Step-size $\tau$, Sampling parameter $K$, Round length $q$, Initial point $x^{1}=(1/d_{x},...,1/d_{x})$
1:for $t=1$ to $T$ do
2:  $w^{t}=\frac{(t-1)w^{t-1}+x^{t}}{t}$
3:  if $t\leq q$ or $t=0$ mod $q$ then
4:   $\hat{w}^{t}=\frac{1}{K}\sum_{k\in[K]}\hat{w}^{t,k}$, where $(\hat{w}^{t,i})_{i\in[K]}\overset{\text{iid}}{\sim}P_{w^{t}}$
5:  else
6:   $\hat{w}^{t}=\hat{w}^{t-1}$
7:  end if
8:  $g^{t}=\nabla f(\hat{w}^{t};B^{t})$, where $B^{t}$ is a batch of $B=n/T$ fresh samples.
9:  $x^{t+1}_{j}\propto x^{t}_{j}\exp\left(-\tau g^{t}_{j}\right),\quad\forall j\in[d_{x}]$
10:end for
11:return $\hat{w}^{T}=\frac{1}{K}\sum_{k\in[K]}\hat{w}^{T,k}$, where $(\hat{w}^{T,i})_{i\in[K]}\overset{\text{iid}}{\sim}P_{w^{T}}$
4 Differentially Private Stochastic Convex Optimization
We show how our techniques can be used for DP-SCO in the $\ell_{1}$ setting. The setup is similar to DP-SSP in the $\ell_{1}$ setting: consider an unkown distribution $\mathcal{D}$ over a sample space $\mathcal{Z}$. We have access to $\mathcal{D}$ only through a sample $S=\{z^{1},...,z^{n}\}\overset{\text{iid}}{\sim}\mathcal{D}$. Consider a loss function $f:\Delta_{x}\times\mathcal{Z}\to\mathbb{R}$, where $\Delta_{x}$ is the standard simplex in $\mathbb{R}^{d_{x}}$. We assume that for all $z\in\mathcal{Z}$, the function $f(\cdot,z)$ is $L_{1}$-smooth and $L_{0}$-Lipschitz w.r.t $\|\cdot\|_{1}$ over $\Delta_{x}$. Let $F_{\mathcal{D}}(x)=\mathbb{E}_{z\sim\mathcal{D}}[f(x;z)]$ be the population loss. We are interested in solving $\min_{x\in\Delta_{x}}F_{\mathcal{D}}(x)$ with $(\varepsilon,\delta)$-DP algorithms. Many ideas used in the previous section for the SSP problem will be useful for here as well. We observe that guarantees for (entropic) stochastic mirror descent can be obtained by applying online-to-batch conversions on exponentiated gradient descent [orabona_online_learning]. By contrast, we will next apply anytime online-to-batch conversions to exponentiated gradient descent [anytime-online-to-batch]. The version we provide below for this result is slightly different than [anytime-online-to-batch] (see [online_to_batch] for a proof).
Lemma 4.1 (Anytime online-to-batch [anytime-online-to-batch, online_to_batch]).
Let $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ be a convex set. For any sequence $\beta_{t}\in\mathbb{R}^{+},g^{t}\in\mathbb{R}^{d_{x}}$, suppose an online learner predicts $x^{t}$ and receives $t$-th loss $\ell_{t}(x)=\langle g^{t},x\rangle$. Define $w^{t}=\sum_{i\in[t]}\frac{\beta_{i}x^{i}}{\sum_{j\in[t]}\beta_{j}}$ and $\operatorname{Regret}_{T}(x)=\sum_{t\in[T]}\langle g_{t},x_{t}-x\rangle$. Then, for any convex, differentiable $F$ over $\mathcal{X}$ and any $x\in\mathcal{X},$
$F(w^{T})-F(x)\leq\frac{1}{\sum_{j\in[T]}\beta_{j}}\bigg(\operatorname{Regret}_{T}(x)+\sum_{t\in[T]}\langle\beta_{t}\nabla F(w^{t})-g_{t},x_{t}-x\rangle\bigg).$
An advantage of this conversion is that the weighted averages $w^{t}$ change slowly from one iteration to the next, which is very useful in the smooth case [online_to_batch]. Using Lemma 4.1 with weights $\beta_{j}=1$ to optimize over $\Delta_{x}$, the average iterates satisfy $\|w^{t}-w^{t-1}\|_{1}=\|x^{t}-w^{t-1}\|_{1}/t\leq 2/t$. Consequently, $P_{w^{t-1}}$ and $P_{w^{t}}$ are progressively closer distributions. We exploit this idea in Algorithm 5 by maintaining estimates of the private iterates for a large number of steps, aggressively reducing the number of sampled vertices and improving the privacy. Theorem 4.2 provides a faster rate compared to the one we would get with the plain online-to-batch conversion.
Theorem 4.2 (Guarantees for Algorithm 5).
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. Let $f(\cdot;z)$ be convex and first-order-smooth. Then, there exists a parameter setting for $T,K,q,\tau$ such that Algorithm 5 run with these as inputs returns $\hat{w}^{T}$ such that $\hat{w}^{T}$ is $(\varepsilon,\delta)$-DP and
$\mathbb{E}[F_{\cal D}(\hat{w}^{T})-F_{\cal D}(x)]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\log(d_{x})}{n}}+L_{0}\frac{\log(d_{x})^{7/10}\log(1/\delta)^{1/5}}{(n\varepsilon)^{2/5}}\right]\\ +L_{1}\log(\sqrt{n}\log(d_{x}))\left[\frac{1}{\log(d_{x})\sqrt{n}}+\frac{\log(1/\delta)^{1/5}}{\log(d_{x})^{4/5}(n\varepsilon)^{2/5}}\right].$
If $f(\cdot;z)$ is convex and second-order-smooth, Algorithm 5 with a different choice for $T,K,q,\tau$ outputs $\hat{w}^{T}$ that is $(\varepsilon,\delta)$-DP and satisfies
$\mathbb{E}[F_{\cal D}(\hat{w}^{T})-F_{\cal D}(x)]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\log(d_{x})}{n}}+\frac{\log(d_{x})\log(1/\delta)^{1/4}}{\sqrt{n\varepsilon}}\right]+\frac{L_{2}}{\sqrt{n}\log(d_{x})^{1/4}}\\ +\frac{L_{2}\log(d_{x})^{1/4}\log(1/\delta)^{1/4}}{\sqrt{n\varepsilon}}+L_{1}\log(\sqrt{n\log(d_{x})})\Big[\frac{1}{\sqrt{n}}+\frac{\log(1/\delta)^{1/4}}{\sqrt{n\varepsilon}}\Big].$
4.1 Privacy Analysis of Algorithm 5
Now we make explicit the advantage of using the anytime online-to-batch approach in terms of privacy. Namely, due to the substantial sampling reduction, our privacy budget degrades much more nicely.
Proposition 4.3.
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. Then, Algorithm 5 is $(\varepsilon,\delta)$-DP if $\tau\leq\frac{B\varepsilon}{8L_{0}\sqrt{2(TK/q+qK)\log(1/\delta)}}$.
Proof 4.4.
Note that $w^{t}=\frac{1}{t}\sum_{i\in[t]}x^{i}$. Hence, for any $j\in[d_{x}]$
$w^{t}_{j}=\frac{1}{t}\sum_{i\in[t]}x^{i}_{j}\propto\sum_{i\in[t]}\exp\left(-\tau\sum_{k\in[i-1]}g^{k}_{j}\right)=\exp\left(\log\left(\sum_{i\in[t]}\exp\left(-\tau\sum_{k\in[i]}g^{k}_{j}\right)\right)\right).$
Denote by $G_{j}(S)=\big(\sum_{k\in[i-1]}g^{k}_{j}\big)_{i\in[t]}$ the vector of the cumulative sum of the gradients that our algorithm produces when it has been fed with dataset $S$. Also, define the log-sum-exp function as $\operatorname{LSE}((x_{1},...,x_{k}))=\log(\sum_{i=1}^{k}\exp\{x_{i}\})$. With this new notation, $w^{t}_{j}\propto\exp(\operatorname{LSE}(-\tau G_{j}(S)))$. Two well-known properties of the $\operatorname{LSE}(x)$ function are that it is convex in its argument and its gradient is the softmax function of $x$. As a consequence of the second property mentioned, it is clear that $\|\nabla\operatorname{LSE}(x)\|_{1}=1$ for every $x$. By convexity of $\operatorname{LSE}$, we have
$\operatorname{LSE}(-\tau G_{j}(S))$ $-\operatorname{LSE}(-\tau G_{j}(S^{\prime}))\leq\langle\nabla\operatorname{LSE}(-\tau G_{j}(S)),\tau G_{j}(S^{\prime})-\tau G_{j}(S)\rangle$
$\leq\tau\|\nabla\operatorname{LSE}(-\tau G(S))\|_{1}\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}=\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}.$
Furthermore, $\max_{j\in[d_{x}]}\tau\|G_{j}(S)-G_{j}(S^{\prime})\|_{\infty}=\tau\max_{j\in[d_{x}],i\in[t]}\left|\sum_{k\in[i-1]}(g^{k}_{j}-g^{\prime k}_{j})\right|\leq\frac{2\tau L_{0}}{B}$ by $L_{0}$-Lipschitzness of $f$. Hence sampling vertices from $\Delta_{x}$ according to the probabilities induced by $w^{t}$ is equivalent to running the exponential mechanism with score function $s(S,j)=\operatorname{LSE}(-\tau G_{j}(S))$, whose sensitivity is bounded by $\frac{2\tau L_{0}}{B}$. Since we sample $K$ times during the first $q$ iterations, then every $q$ iterations and then $K$ times to sparsify the last iterate, the overall number of samplings will be $TK/q+qK+K$. Since each of them is $\frac{4\tau L_{0}}{B}$-DP, then by Advanced Composition of DP (Corollary 3.6) we only need that $\frac{4\tau L_{0}}{B}\leq\frac{\varepsilon}{2\sqrt{2(TK/q+(q+1)K)\log(1/\delta)}}$.
4.2 Convergence Analysis of Algorithm 5
Proof 4.5 (Proof of Theorem 4.2).
Let $x\in\Delta_{x}$ be a minimizer of $F_{\mathcal{D}}(\cdot)$. From Lemma 4.1, $F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)\leq\frac{1}{T}\big[\sum_{t\in[T]}\langle g_{t},x_{t}-x\rangle+\sum_{t\in[T]}\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\big]$. The standard analysis of mirror descent yields, $\sum_{t\in[T]}\langle g_{t},x_{t}-x\rangle\leq\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2}$ [JNLS:2009]. Using that $\|g^{t}\|_{\infty}\leq L_{0}$ and taking expectations, we obtain
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]\leq\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right].$
Let us bound $\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$. For $t\leq q$, we have $g^{t}\!=\!\nabla f(\hat{w}^{t};B^{t})$, and
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]=\mathbb{E}[\langle\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]$
$=\mathbb{E}_{x^{t}}[\langle\mathbb{E}_{\hat{w}^{t}}[\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t})\mid x^{t}],x^{t}-x\rangle]$
$\leq 2\|\mathbb{E}_{\hat{w}^{t}}[\nabla f(w^{t};B^{t})-\nabla f(\hat{w}^{t};B^{t})]\|_{\infty}\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{4L_{2}}{K}&\text{for second-order-smooth $f(\cdot;z)$.}\end{cases}$
where in the first equality we used that $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ and for the second inequality we used Corollaries 2.6 and 2.4. Now consider the case $t>q$. First, denote by $t_{0}=q\lfloor t/q\rfloor$. That is, $t_{0}$ was the last time that $\hat{w}$ was updated by the algorithm before the $t$-th iteration. Hence, $g^{t}=\nabla f(\hat{w}^{t_{0}};B^{t})$. Now, we decompose the error into three terms and bound each of them individually:
$\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle$ $=\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle$
$+\langle\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t}),x^{t}-x\rangle$
$+\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle.$
Clearly $\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle]=0$, since $\nabla f(w^{t};B^{t})$ is an unbiased estimator of $\nabla F_{\mathcal{D}}(w^{t})$ conditionally on the randomness up to iteration $t$. For the second term, note that
$\langle\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t}),x^{t}-x\rangle\leq\|\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t})\|_{\infty}\|x^{t}-x\|_{1}$
$\leq 2L_{1}\|w^{t}-w^{t_{0}}\|_{1}\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\|w^{k+1}-w^{k}\|_{1}\leq 2L_{1}\sum_{k=t_{0}}^{t-1}\frac{2}{k+1}\leq\frac{4L_{1}q}{q\lfloor t/q\rfloor+1},$
where the first inequality follows from Hölder, the second from the $L_{1}$-smoothness of $f$, the third from the triangle inequality for $\|\cdot\|_{1}$, and the fourth from the stability of the average iterates. Now, we decompose
$\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle=A_{1}+A_{2},$
with $A_{1}:=\langle\widetilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle$ and $A_{2}:=\langle\widetilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle$, where $\widetilde{\Delta}^{t_{0}}$ denotes $\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})$ and $\tilde{x}^{t}$ is a ficticious iterate obtained by starting at $x^{t_{0}}$ and performing updates $\tilde{x}_{j}^{t+1}\propto\tilde{x}_{j}^{t}\exp\big(-\tau\nabla_{j}f(w^{t_{0}};B^{k})\big)$. We start by bounding $A_{1}$:
$\mathbb{E}_{\hat{w}^{t_{0}}}[\langle\widetilde{\Delta}^{t_{0}},\tilde{x}^{t}-x\rangle]=\langle\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})],\tilde{x}^{t}-x\rangle$
$\leq 2\|\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})]\|_{\infty}\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{4L_{2}}{K}&\text{for second-order-smooth $f(\cdot;z)$,}\end{cases}$
where the equality holds since $\tilde{x}^{t}$, as opposed to $x^{t}$, does not depend on $\hat{w}^{t_{0}}$, and the last inequality follows from Corollaries 2.6 and 2.4. We bound now $A_{2}$. Note that $\langle\widetilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle\leq\|\widetilde{\Delta}^{t_{0}}\|_{\infty}\|x^{t}-\tilde{x}^{t}\|_{1}$. Denote $\hat{G}_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(\hat{w}^{t_{0}};B^{k}),G_{j}=-\tau\sum_{k=t_{0}}^{t-1}\nabla_{j}f(w^{t_{0}};B^{k})$. We bound $|(x^{t}-\tilde{x}^{t})_{j}|$ as follows
$(x^{t}-\tilde{x}^{t})_{j}$ $=\frac{x^{t_{0}}_{j}\exp(\hat{G}_{j})}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp(\hat{G}_{i})}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$=\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)\exp(\hat{G}_{j}-G_{j})}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)\exp(\hat{G}_{i}-G_{i})}-\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}$
$\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp(G_{i})}\Big[\exp\big(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\big)-1\Big],$
and similarly $(\tilde{x}^{t}-x^{t})_{j}\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\left[2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\right]$. Combining both inequalities, we obtain $|(x^{t}-\tilde{x}^{t})_{j}|\leq\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\psi$, where
$\psi:=\max\Big\{2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|,\exp\big(2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\big)-1\Big\}$. It follows that $\|x^{t}-\tilde{x}^{t}\|_{1}=\sum_{j\in[d_{x}]}|(x^{t}-\tilde{x}^{t})_{j}|\leq\sum_{j\in[d_{x}]}\frac{x^{t_{0}}_{j}\exp\left(G_{j}\right)}{\sum_{i\in[d_{x}]}x^{t_{0}}_{i}\exp\left(G_{i}\right)}\psi=\psi$. Next, we use that $e^{x}-1\leq 2x$ for $x\leq 1$. Assume for now $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$, so $\psi=4\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|$. Using the definition of $\hat{G}_{j},G_{j}$, it follows that
$\|x^{t}-\tilde{x}^{t}\|_{1}\leq\psi$ $=\textstyle 4\max_{j\in[d_{x}]}\tau\left|\sum_{k=t_{0}}^{t-1}[\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})]\right|$
$\textstyle\leq 4\tau\max_{j\in[d_{x}]}\sum_{k=t_{0}}^{t-1}\left|\nabla_{j}f(\hat{w}^{t_{0}};B^{k})-\nabla_{j}f(w^{t_{0}};B^{k})\right|$
$\textstyle\leq 4\tau\sum_{k=t_{0}}^{t-1}\left\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\right\|_{\infty}.$
This has two implications. The first is that $\psi\leq 8\tau qL_{0}$, by $L_{0}$-Lipschitzness of $f$. This further implies that in order to ensure that $2\max_{j\in[d_{x}]}|\hat{G}_{j}-G_{j}|\leq 1$ holds, it suffices that $8\tau qL_{0}\leq 2$, or equivalently $\tau\leq 1/(4L_{0}q)$. We impose this condition on $\tau$ later when optimizing the convergence rate. The second is the following bound on $A_{2}$
$A_{2}=\langle\widetilde{\Delta}^{t_{0}},x^{t}-\tilde{x}^{t}\rangle=\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-\tilde{x}^{t}\rangle$
$\leq\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}\|x^{t}-\tilde{x}^{t}\|_{1}$
$\textstyle\leq 4\tau\sum_{k=t_{0}}^{t-1}\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}$
$\textstyle\leq 2\tau\sum_{k=t_{0}}^{t-1}\Big(\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\Big)$
$\textstyle\leq 2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}$
$\lesssim\begin{cases}q\tau\big[\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big]&\text{for first-order-smooth $f(\cdot;z)$}\\ q\tau\big[\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big]&\text{for second-order-smooth $f(\cdot;z)$,}\end{cases}$
where we bounded the second moment of the estimation errors with the results in Corollaries 2.12 and 2.14. We conclude that for all $t>q$
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle]$
$=\mathbb{E}\big[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(w^{t};B^{t}),x^{t}-x\rangle\big]+\mathbb{E}\big[\langle\nabla f(w^{t};B^{t})-\nabla f(w^{t_{0}};B^{t}),x^{t}-x\rangle\big]$
$\quad+\mathbb{E}\big[\langle\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t}),x^{t}-x\rangle\big]$
$\leq 0+\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+2\|\mathbb{E}_{\hat{w}^{t_{0}}}[\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})]\|_{\infty}$
$\,+2\tau q\|\nabla f(w^{t_{0}};B^{t})-\nabla f(\hat{w}^{t_{0}};B^{t})\|_{\infty}^{2}+2\tau\sum_{k=t_{0}}^{t-1}\mathbb{E}\left[\|\nabla f(\hat{w}^{t_{0}};B^{k})-\nabla f(w^{t_{0}};B^{k})\|_{\infty}^{2}\right]$
$\lesssim\begin{cases}\frac{L_{1}}{\sqrt{K}}+q\tau\big[\frac{L_{1}^{2}}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}}{\sqrt{K}}\big]&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{L_{2}}{K}+q\tau\big[\frac{L_{2}^{2}}{K^{2}}+\frac{L_{1}^{2}\log(d_{x})}{K}\big]&\text{for second-order-smooth $f(\cdot;z)$.}\end{cases}$
And recall that for $t\leq q$,
$\mathbb{E}[\langle\nabla F_{\mathcal{D}}(w^{t})-\nabla f(\hat{w}^{t};B^{t}),x^{t}-x\rangle]\leq\begin{cases}\frac{8L_{1}}{\sqrt{K}}&\text{for first-order-smooth $f(\cdot;z)$}\\ \frac{4L_{2}}{K}&\text{for second-order-smooth $f(\cdot;z)$.}\end{cases}$
Finally, we plug this bounds into our excess risk bound. We start considering the case of second-order-smooth functions. We bound $\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$ by
$\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\textstyle\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\Big(\frac{L_{2}}{K}+\mathbbm{1}_{(t>q)}\Big[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}+q\frac{L_{2}^{2}\tau}{K^{2}}+q\frac{L_{1}^{2}\log(d_{x})\tau}{K}\Big]\Big)$
$\textstyle\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big]+\frac{L_{2}}{K}+\frac{1}{T}\sum_{t=q+1}^{T}\frac{L_{1}q}{q\lfloor t/q\rfloor+1}$
(11) $\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q\log(d_{x})}{K}+\frac{L_{2}^{2}q}{K^{2}}\big]+\frac{L_{2}}{K}+\frac{L_{1}q\log(T/q)}{T},$
where the last line follows from $\sum_{t=q+1}^{T}\left[\frac{1}{q\lfloor t/q\rfloor+1}\right]\leq q\left(\frac{1}{q}+\frac{1}{2q}+...+\frac{1}{(T/q)q}\right)\lesssim\log\big(\frac{T}{q}\big)$. Now, by Theorem 4.3, and recalling that the batch size is $B=n/T$, we have a constraint on $\tau>0$ in order to satisfy privacy. We also have the constraint of $\psi\leq 1$, which holds true only when $\tau\leq 1/(8qL_{0})$. Optimizing (11) under these constraints, we have that we should choose $\tau$ of order
$\min\bigg\{$ $\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\bigg\}.$
With this choice we obtain
$\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$ $\lesssim\sqrt{\frac{(L_{0}^{2}+L_{1}^{2}q\log(d_{x})/K+L_{2}^{2}q/K^{2})\log(d_{x})}{T}}+\frac{L_{0}q\log(d_{x})}{T}$
$\qquad+\frac{L_{0}\sqrt{(TK/q+qK)\log(1/\delta)}\log(d_{x})}{n\varepsilon}+\frac{L_{2}}{K}+\frac{L_{1}q\log(T/q)}{T}.$
Note that the sparsification error is, by Corollary 2.3, $\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(w^{T})]\lesssim\frac{L_{1}}{K}$. Combining this with the previous inequality and setting $q=\sqrt{T/\log(d_{x})},K=T/q=\sqrt{T\log(d_{x})}$, leads to
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}\\ \textstyle+\frac{L_{2}}{T^{3/4}\log(d_{x})^{1/4}}+\frac{L_{1}+L_{2}}{\sqrt{T\log(d_{x})}}+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}\\ \textstyle\lesssim(L_{0}+L_{1})\sqrt{\frac{\log(d_{x})}{T}}+\frac{L_{2}}{\sqrt{T}\log(d_{x})^{1/4}}+\frac{L_{1}\log(\sqrt{n\log(d_{x})})}{\sqrt{T\log(d_{x})}}+\frac{L_{0}\sqrt{T\log(1/\delta)}\log^{3/2}(d_{x})}{n\varepsilon}.$
Hence letting $T=\min\Big\{n,\frac{n\varepsilon}{\log(d_{x})\sqrt{\log(1/\delta)}}\Big\}$ gives the claimed rate for second-order-smooth functions. Now, we look at the case where the partial derivates are not necessarily smooth. Now, we can bound $\mathbb{E}[F_{\mathcal{D}}(w^{T})-F_{\mathcal{D}}(x)]$ by
$\frac{\log(d_{x})}{\tau T}+\frac{\tau L_{0}^{2}}{2}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\langle\nabla F_{\mathcal{D}}(w^{t})-g^{t},x^{t}-x\rangle\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau L_{0}^{2}+\frac{1}{T}\sum_{t\in[T]}\bigg(\frac{L_{1}}{\sqrt{K}}+\mathbbm{1}_{(t>q)}\bigg[\frac{4L_{1}q}{q\lfloor t/q\rfloor+1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{L_{1}^{2}q\tau}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q\tau}{\sqrt{K}}\bigg]\bigg)$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big]$
$\qquad\qquad+\frac{L_{1}}{\sqrt{K}}+\frac{1}{T}\sum_{t=q+1}^{T}\left[\frac{L_{1}q}{q\lfloor t/q\rfloor+1}\right]$
$\lesssim\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big]+\frac{L_{1}}{\sqrt{K}}+\frac{L_{1}q\log(T/q)}{T}.$
By Corollary 2.3, the sparsification error is $\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(w^{T})]\lesssim\frac{L_{1}}{K}$, which implies
$\mathbb{E}[F_{\mathcal{D}}(\hat{w}^{T})-F_{\mathcal{D}}(x)]\lesssim$ $\frac{\log(d_{x})}{\tau T}+\tau\Big[L_{0}^{2}+\frac{L_{1}^{2}q}{\sqrt{\log(d_{x})}K^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d_{x})}q}{\sqrt{K}}\Big]$
$\quad+\frac{L_{1}}{\sqrt{K}}+\frac{L_{1}q\log(T/q)}{T}.$
Setting $T=\min\left\{n,\frac{(n\varepsilon)^{4/5}}{(\log(d_{x})\log(1/\delta))^{2/5}}\right\}$, $q=\sqrt{T}/\log(d_{x})$, $K=T/\log(d_{x})$ and
$\tau\!\eqsim\!\!\min\!\left\{\sqrt{\frac{\log(d_{x})}{(L_{0}^{2}+(L_{0}^{2}+L_{1}^{2})q\sqrt{\log(d_{x})/K}+L_{1}^{2}q/[\sqrt{\log(d_{x})}K^{3/2}])T}},\frac{1}{L_{0}q},\frac{n\varepsilon}{TL_{0}\sqrt{(TK/q+qK)\log(1/\delta)}}\right\}$ leads to the convergence rate claimed for first-order-smooth functions.
5 Conclusion and Future Work
We presented mirror descent based algorithms for DP-SSP in the polyhedral setting. Our algorithms are the first achieving polylogarithmic-in-the-dimension error beyond bilinear objectives. In fact, for second-order-smooth objectives our rates are optimal up to logarithmic factors. An important question remains: Are the optimal rates achievable in expectation for first-order-smooth functions? We leave this question for future work. We also apply our ideas to DP-SCO in the polyhedral setting, where we also leave open the problem of achieving optimal rates with mirror descent based algorithms.
6 Acknowledgments
T. González’s research was partially done as a Research Intern at Google DeepMind. C. Guzmán’s research was partially done as a Visiting Researcher at Google. C. Guzmán’s research was partially supported by INRIA Associate Teams project, ANID FONDECYT 1210362 grant, ANID Anillo ACT210005 grant, and National Center for Artificial Intelligence CENIA FB210017, Basal ANID. C. Paquette is a CIFAR AI chair, (MILA) and C. Paquette was supported by a NSERC Discovery Grant, NSERC CREATE grant, an external grant from Google, and FRQNT New University Researcher’s Start-Up Program.